### v1

# Linear Discriminant Regularized Regression

Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. Moreover, our new formulation is generically easy to analyze compared to existing regression-based LDA procedures. In particular, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization that has not yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
Keywords: Multi-class classification, discriminant analysis, high-dimensional data, regularization, multivariate response regression, dimension reduction.
1 Introduction
This paper introduces a new regression-based approach to multi-class, high-dimensional linear discriminant analysis (LDA). The main premise that leads to LDA is that we have a sample of $n$ i.i.d. copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, with the label $Y\in\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ taking values in $L\geq 2$ classes and the conditional distributions of the feature vector $X\in\mathbb{R}^{p}$ satisfy
$X\mid Y={\bm{e}}_{\ell}~{}\sim~{}\mathcal{N}_{p}\left(\mu_{\ell},\Sigma_{w}\right),\qquad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$ (1.1)
Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite. For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$. Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem. Since we can always subtract the marginal mean of $X$, we assume
$\mathbb{E}(X)=0_{p}$ (1.2)
for simplicity, and furthermore that the class probabilities
$\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ (1.3)
are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) and (1.3) is to classify a new feature $x\in\mathbb{R}^{p}$ as
$g^{*}(x):=\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.4)
See, for instance, Izenman (2008). Estimation of this classifier based on $(X_{i},Y_{i})$ with $i\in[n]$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). Write
$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\qquad D_{\pi}={\rm diag}(\pi),\qquad{\mathbb{M}}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.5)
and denote the unconditional covariance matrix of $X$ by
$\Sigma={\rm Cov}(X)=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}.$
We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}$ the between-class covariance matrix. The latter due to (1.2) has rank less than $L$.
In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure to estimate this direction (see, also, Cai and Zhang (2019) for extension to $L>2$). It isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed later. Witten and Tibshirani (2011) use the fact that the Bayes rule (1.4) is equivalent with Fisher’s discriminant rule
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.6)
Here the columns of the matrix $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are implicitly defined via, for $1\leq k\leq K$,
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}&~{}\beta^{\top}\Sigma_{b}\beta\\ \text{such that}&\quad\beta^{\top}\Sigma_{w}\beta=1,\quad\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.7)
Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the implicit formulation (1.7) of the matrix $F$ makes it extremely difficult to obtain the global solution and to analyze the theoretical properties of the regularized estimator, hence the resulting classifier. More closely related to our approach, introduced in the next paragraph, is optimal scoring, which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein.
By writing the response matrix and feature matrix as
${\mathbb{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbb{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.8)
optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}\\ \text{such that }&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.9)
Here $\widehat{\theta}_{i}$ is the solution to (1.9) in the $i$th iteration. The minimization problem (1.9) can be done via solving a generalized eigenvalue problem, see Hastie et al. (1995). Moreover, Hastie et al. (1995) also show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.9) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.7) in the low-dimensional case ($p<n$). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$.
Since optimal scoring is a regression-based approach, as pointed out in Hastie et al. (1995, 1994), it is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the procedure of Cai and Zhang (2019).
In the high-dimensional setting, Hastie et al. (1995) show that the global solution to a penalized version of (1.9), that is, for $1\leq k\leq K<L$,
$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\\ \text{such that}&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.10)
can be computed for any $p\times p$ symmetric, positive definite matrix $\Omega$. A possible choice of $\Omega=\lambda\mathbb{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). In a more recent work (Gaynanova, 2020), the matrix formulation of (1.10)
$\begin{split}\min_{\Theta,~{}B}&~{}\|{\mathbb{Y}}\Theta-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B)\\ \text{such that }&~{}\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\Theta=n\mathbb{I}_{K},\quad\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}1_{n}=0_{K}.\end{split}$ (1.11)
is considered, and it is proved that the global solution $\widehat{B}\in\mathbb{R}^{p\times K}$ to (1.11) can be computed for any penalty $\text{pen}(B)$ that satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. This allows for the penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ used in the group-lasso, but excludes the familiar $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ used in the lasso regression as well as its extensions such as the elastic net penalty. That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.11), let alone any theoretical property of the resulting estimator.
Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1) and (1.3)
$B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}\in\mathbb{R}^{p\times L}$ (1.12)
and the regression matrix
$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)$ (1.13)
not only share the same column space, but have a closed-form connection, to wit,
$B^{*}$ $=BH^{-1}$
for some invertible $L\times L$ matrix $H$. This connection is novel and distinguishes our approach from the current literature. Although the fact that $B$ and $B^{*}$ share the same span is relatively well-known (see, for instance, Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022) and the references therein), their closed-form connection has not been discovered and a thorough theoretical investigation on the corresponding classifier has not been conducted, to the best of our knowledge. Specifically, we propose to estimate the discriminant matrix $B^{*}$ by estimating the regression matrix $B$ first
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbb{Y}}-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
and then estimating $H^{-1}$ based on $\widehat{B}$. The advantage of our approach is three-fold:
  1. (a)
    We can easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$.
  2. (b)
    After estimating $B$, we only need to estimate the inverse of an $L\times L$ matrix rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
  3. (c)
    We are able to quantify the estimation error of both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier.
While various penalties are feasible in our approach, more specifically in (1.14), we focus on a lasso-type estimator of $B$ in this paper. This also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. Our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ and unobserved noise $W$, that is independent of both $Z$ and the label $Y$ with $r<n<p$. The procedure is entirely different than the one proposed here. Future work will be devoted on how to adapt the procedure in this paper to this low-rank setting and its analysis.
Under sparsity of the regression matrix $B$, we carefully analyze the properties of $\widehat{B}$ in Section 3.1, the estimation of $B^{*}$ via estimating $H^{-1}$ in Section 3.2 as well as the rate of convergence for the excess misclassification risk of the corresponding classification rule in Section 3.3. Computationally, after computing $\widehat{B}$ from (1.14) we offer an alternative way in Section 2.3 to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which extracts fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization. Finally, we provide a thorough simulation study in Section 4 and we report our findings in several real data studies of Section 5.
2 Methodology
We discuss our classification procedure in this section. Our goal is to estimate the Bayes rule in (1.4) which, for a given $x\in\mathbb{R}^{p}$, requires estimation of the discriminant functions
$G_{\ell}(x):=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}),\qquad\text{for all }\ell\in[L].$ (2.1)
This, in turn, requires that we estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.13). Recall the definitions of $\pi$, $D_{\pi}$ and ${\mathbb{M}}$ from (1.5) as well as $\Sigma={\rm Cov}(X)$.
Lemma 1.
Under model (1.1) and (1.3), we have
$B^{*}=BH^{-1}$ (2.2)
for a matrix $H\in\mathbb{R}^{L\times L}$ that can be inverted and satisfies
$H=D_{\pi}-B^{\top}\Sigma B=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=D_{\pi}-B^{\top}{\mathbb{M}}D_{\pi}.$ (2.3)
Proof.
From the identity
$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={\mathbb{M}}D_{\pi}$ (2.4)
we see that
$B=\Sigma^{-1}{\mathbb{M}}D_{\pi}.$ (2.5)
Next, using
$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top},$ (2.6)
the Woodbury matrix identity yields
$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}.$
It follows that
$B^{*}$ $=\Sigma_{w}^{-1}{\mathbb{M}}$
$=\Sigma^{-1}{\mathbb{M}}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}$
$=\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}$
$=\Sigma^{-1}{\mathbb{M}}D_{\pi}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
$=B\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)^{-1}$
which completes the proof. ∎
Lemma 1 suggests that estimation of $B^{*}$ could be done by first estimating the matrix $B$ and then estimating $H^{-1}$ according to (2.3). We discuss these two steps in details in the next two sections.
2.1 Estimation of the regression matrix
In view of (1.13), it is natural to estimate $B$ via regularized regression methods in (1.14). Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand. For example, if we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions. If the subset of features, that are pertinent for predicting the label, varies across different levels, then the entry-wise lasso penalty is appropriate. To fix ideas and due to technical complications, we focus our analysis to the matrix estimator $\widehat{B}$ in (1.14) obtained with the following concrete lasso penalty
$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}$ (2.7)
for some tuning parameter $\lambda>0$. It is fairly straightforward to modify our analysis to other choices of the penalty term discussed below.
Remark 1 (Alternative penalty choices).
If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ is a more suitable choice. In this case, one can retain a much smaller subset of important features and achieve feature selection. Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty if the features tend to (highly) correlate with each other. Another example that is commonly used in the literature is when entries of $B$ exhibit certain smoothness structure, such as some imaging or audio applications in Hastie et al. (1995). In this case, one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$. On the other hand, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate. For example, if the matrix of conditional means ${\mathbb{M}}\in\mathbb{R}^{p\times L}$ has a reduced rank $r\ll\min\{p,L\}$ such as under the model in Bing and Wegkamp (2023), then Lemma 1 implies that ${\rm rank}(B)\leq r$. Estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using either the reduced rank penalty $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ or the nuclear norm penalty $\text{pen}(B)=\lambda\|B\|_{*}$. We will investigate this approach in another paper.
2.2 Estimation of the discriminant matrix and the Bayes rule
The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.4) necessitate that we estimate the matrix of the conditional means ${\mathbb{M}}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$. We adopt the following standard estimators
$\widehat{\mathbb{M}}={\mathbb{X}}^{\top}{\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbb{Y}}^{\top}{\mathbb{Y}}$ (2.8)
to estimate ${\mathbb{M}}$ and $D_{\pi}$, respectively. Recall from Lemma 1 that $B^{*}=BH^{-1}$. Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}={1\over n}\left({\mathbb{Y}}^{\top}{\mathbb{Y}}-\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}\right).$ (2.9)
Note that ${\mathbb{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14). Computation of ${\mathbb{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbb{X}}$ consist of basis expansions. Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}.$ (2.10)
Here $M^{+}$ denotes the moore-penrose inverse of any matrix $M$. Finally, in view of (2.1), we estimate the discriminant functions by
$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell})\quad\text{for all $\ell\in[L]$. }$ (2.11)
The corresponding classifier is
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad\text{ for any $x\in\mathbb{R}^{p}$}.$ (2.12)
Theoretically, we will show that $\widehat{H}$ is invertible with overwhelming probability for the lasso-penalty (see, Remark 3 as well as the proof of Theorem 3 in Section A.3). Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.14).
2.3 Classification based on reduced dimensions
Since in practice it is often desirable to extract fewer discriminant directions than those in $B^{*}$, we discuss in this section an alternative procedure for this purpose, which estimates the Bayes rule by also using $\widehat{B}$ from (1.14) but in a different way.
We start with the fact (its proof is stated in Section A.1 for completeness) that the Bayes rule in (1.4) is equivalent to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$ (2.13)
which, by Lemma 1, further equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B(B^{\top}\Sigma_{w}B)^{+}B^{\top}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (2.14)
The formulation in (2.14) has the following two-step interpretation:
  1. (1)
    Transform the original feature $x\in\mathbb{R}^{p}$ to $z:=B^{\top}x$, with the latter belonging to a subspace of $\mathbb{R}^{L}$;
  2. (2)
    Perform linear discriminant analysis on the space of $z$.
The second point follows by noticing that
$C_{w}:=B^{\top}\Sigma_{w}B\in\mathbb{R}^{L\times L}$
is indeed the within-class covariance matrix of $Z=B^{\top}X$ under model (1.1).
Due to the fact that $C_{w}$ is rank deficient with ${\rm rank}(C_{w})<L$ and for the purpose of extracting fewer discriminant directions, we adopt the equivalent formulation of (2.14) from the Fisher’s perspective in the sequel. Write the between-class covariance matrix of $Z=B^{\top}X$ as
$C_{b}:=B^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the minimization problem
$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\\ \text{such that }&~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\text{ for all }~{}i<k.\end{split}$
The resulting classification rule based on these $K$ discriminant directions is
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B\sum_{k=1}^{K}\alpha_{k}\alpha_{k}^{\top}B^{\top}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (2.15)
In practice, after computing $\widehat{B}$ from (1.14), we propose to estimate $C_{b}$ and $C_{w}$ by
$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{\mathbb{M}}D_{\widehat{\pi}}\widehat{\mathbb{M}}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}P_{{\mathbb{Y}}}{\mathbb{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}(\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B},$
with $P_{{\mathbb{Y}}}={\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1}{\mathbb{Y}}^{\top}$. Subsequently, the estimated discriminant directions $\widehat{\alpha}_{k}$, for $1\leq k\leq K$, are computed by solving
$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\\ \text{such that}&~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,\quad\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.16)
By similar reasoning for solving the Fisher’s discriminant analysis in (1.7), we propose to solve the optimization problem (2.16) via eigen-decomposing
$[\widehat{C}_{w}^{+}]^{1\over 2}\widehat{C}_{b}[\widehat{C}_{w}^{+}]^{1\over 2}.$
The eigenvalues of the above matrix can also be used to determine the upper bound for $K$. For instance, $K$ should be chosen no greater than the rank of the above matrix.
Remark 2 (Numerical stability of computing (2.16)).
The procedure in (2.16) computes Fisher’s discriminant directions $(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K})$ in
$\text{span}({\mathbb{X}}\widehat{B})\subset\text{span}({\mathbb{X}})$
of dimension less than $L$. Consequently, as long as
${\rm rank}((\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B})={\rm rank}({\mathbb{X}}\widehat{B}),$ (2.17)
we do not suffer numerically in solving (2.16), in stark contrast with the algorithmic issue of solving (1.7) (or a penalized version thereof) in the original $p$-dimensional space. As pointed out by Wu et al. (2015), the latter leads to meaningless solutions with the objective function equal to infinity. Finally, we remark that the requirement in (2.17) is easy to satisfy as ${\rm rank}({\mathbb{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbb{I}_{n}-P_{{\mathbb{Y}}})=n-L$ when we observe at least one data point per class.
3 Theoretical guarantees
In this section we provide theoretical guarantees for our procedure. We start by listing the main assumptions used in our analysis. The following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$ is assumed.
Assumption 1.
There exists some absolute constants $0<c\leq C<\infty$ such that
${c\over L}\leq\min_{k\in[L]}\pi_{k}\leq\max_{k\in[L]}\pi_{k}\leq{C\over L}.$ (3.1)
This is a common regularity assumption even in binary classification problems. We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 15 of Appendix C) at the cost of more technical proofs and a less transparent presentation.
As we mentioned earlier, we decided to focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns. Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
$\Theta_{s}:=\left\{B\in\mathbb{R}^{p\times L}:\max_{\ell\in[L]}\|B_{\ell}\|_{0}\leq s\right\}.$
For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$. However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$. Moreover, the analysis of $\widehat{B}$ relies on the standard Restricted Eigenvalue condition (RE) on the matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Since ${\mathbb{X}}$ is a matrix with random entries, we resort to the following quantity on the within-class covariance matrix $\Sigma_{w}$.
Definition 1.
For any integer $1\leq s\leq p$, define
$\kappa_{s}:=\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{v^{\top}\Sigma_{w}v\over v^{\top}v}$
where $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
Our analysis needs the following regularity conditions on $\Sigma_{w}$ and the matrix ${\mathbb{M}}$ of conditional means. For any matrix $M$, write $\|M\|_{\infty}=\max_{i,j}|M_{ij}|$.
Assumption 2.
For some $1\leq s\leq p$, there exists some absolute constants $0<c\leq C<\infty$ such that $c<\kappa_{s}\leq\|\Sigma_{w}\|_{\infty}\leq C$.
Assumption 3.
There exists an absolute constant $C<\infty$ such that $\|{\mathbb{M}}\|_{\infty}\leq C$.
A condition akin to 2 is commonly used in sparse linear regression models, see, for instance, Bühlmann and Van de Geer (2011); Giraud (2021). 3 is a mild regularity condition on the largest entry of ${\mathbb{M}}$. We remark that the requirements that $\kappa_{s}$, $\|\Sigma_{w}\|_{\infty}$ and $\|{\mathbb{M}}\|_{\infty}$ are absolutely bounded are not essential, and can be relaxed to allow dependence on $n$.
3.1 Finite sample bounds of the proposed regression matrix estimation
The following theorem states finite sample upper bounds of both the (in-sample and out-of-sample) prediction error and the estimation error of the regression matrix estimator $\widehat{B}$ in (1.14) with the lasso-penalty in (2.7). For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$. For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
Theorem 2.
Under model (1.1) and (1.3) with Assumptions 13, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with
$(s\vee L)\log(n\vee p)\leq c~{}n$
Then, with probability at least $1-n^{-1}$, we have
$\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbb{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
$\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
$\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}$
for the estimator $\widehat{B}$ obtained from (1.14) with any $\lambda$ satisfying
$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.2)
and some finite, large enough constant $C>0$.
Proof.
See Section A.2. ∎
The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbb{Y}}$ is not a linear model in ${\mathbb{X}}$. We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$ in Lemma 6. For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbb{X}}$ are generated from a mixture of Gaussians. In Lemma 7 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b). Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}^{\prime}-{\mathbb{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbb{Y}}^{\prime}$ and different target matrix $B^{\prime}$. Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
In the rest of the paper, we assume that $\lambda$ is chosen as the order specified in (3.2). In practice, one can simply choose $\lambda$ via cross-validation, for instance, via the built-in function of the R-package glmnet.
3.2 Estimation of the discriminant directional matrix
We analyze the estimation of the discriminant matrix $B^{*}$ in this section. Recall that
$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$
with $\widehat{H}$ obtained from (2.9). The following theorem states upper bounds of the transformed $(\widehat{B}^{*}-B^{*})$. Since $H^{-1}$ depends on ${\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$ (see Lemma 9 of Appendix B for its proof), our results depend in turn on the quantity
$\Delta_{\infty}=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under (1.2))
${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
Theorem 3.
Under the conditions in Theorem 2, assume that $\Delta_{\infty}\geq 1$ and assume there exists some sufficiently small constant $c>0$ such that
$\Delta_{\infty}^{2}\sqrt{sL\log(n\vee p)\over n}\leq c.$ (3.3)
Then, with probability at least $1-n^{-1}$, we have
$\max_{\ell\in[L]}~{}\bigl{\|}\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\bigr{\|}_{2}\!\!~{}\lesssim~{}\Delta_{\infty}^{5/2}\sqrt{sL\log(n\vee p)\over n}.$
Proof.
See Section A.3. ∎
Remark 3.
Condition (3.3) ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds, and hence $\widehat{H}$ is invertible. In Theorem 4 below we will see that condition (3.3) is also needed for consistent estimation of the discriminant function. The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish. It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
3.3 Finite sample bounds of the misclassification rates
We first analyze the estimation of the discriminant functions in (2.1). For any $X$ that is distributed according to (1.1), we have
$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right).$
As a consequence of Theorem 3, the Gaussian distribution assumption of $X\mid Y$ and the independence between $X$ and $({\mathbb{X}},{\mathbb{Y}})$, we have the following control of the estimated discriminant function $\widehat{G}_{\ell}(X)$ uniformly over $\ell\in[L]$.
Theorem 4.
Under the conditions of Theorem 3, we have
$\mathbb{P}\left\{\max_{\ell\in[L]}|\widehat{G}_{\ell}(X)-G_{\ell}(X)|~{}\gtrsim~{}\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}\right\}\leq 2n^{-\Delta_{\infty}}+n^{-1}.$
Proof.
See Section A.4. ∎
Armed with Theorem 4, we are ready to analyze the proposed classifier $\widehat{g}$ in (2.12) in terms of its excess risk relative to error rate of the Bayes rule in (1.4). Specifically, for any pair $(X,Y)$ from (1.1) that is independent of $({\mathbb{X}},{\mathbb{Y}})$, the excess risk of $\widehat{g}$ is defined as
$\mathcal{R}(\widehat{g}):=\mathbb{P}\left\{Y\neq{\widehat{g}(X)}\right\}-\mathbb{P}\left\{Y\neq{g^{*}(X)}\right\}.$
Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. In the proof of the following theorem, we show that, for any $t>0$,
$\mathcal{R}(\widehat{g})$ $\leq(L-1)~{}t~{}\max_{k\in[L]}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}$
$\quad+(L-1)~{}\mathbb{P}\left\{\max_{\ell\in[L]}\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|>t\right\}$
This fact, in combination with Theorem 4, yields the following result on the rates of convergence of $\mathcal{R}(\widehat{g})$.
Theorem 5.
Under the same conditions of Theorem 3 and $\Delta_{\infty}\asymp 1$, we have
$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}.$
Proof.
See Section A.5. ∎
Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$. The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$.
Theorem 5 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$. Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)). Indeed, provided
$\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$
using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}e^{-c\Delta_{\infty}}+n^{-a}$
for some arbitrary large constant $a>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
4 Simulation study
In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR). We also include the procedure in Section 2.3 (called LDRR-F) that is based on Fisher’s discriminant rule after estimating $B$. For estimating $B$ from (1.14), in addition to the lasso penalty (2.7), we also consider the elastic net penalty, $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{2}+(1-\alpha)\|B_{\ell}\|_{1}]$, for some $\alpha\in(0,1)$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet. We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al. (2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al. (2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011). These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters. Finally, the oracle procedure that uses the true values of ${\mathbb{M}}$, $\Sigma_{w}$, and $\pi$ in the computation of the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
We generate the data from the following mean-shift scenario. A similar setting is also used in Witten and Tibshirani (2011). Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $\mu_{\ell,j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0. Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$. The quantity $\sigma>0$ controls the overall noise level while the coefficient $\rho\in[0,1]$ controls the within-class correlation among the features. We generate the class probabilities $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ as follows. For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$. Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$. The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes. At the other extreme, $\alpha=0$ corresponds to the balanced case $\pi=(1/L)1_{L}$. In the following, we examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{100,300,500,700,900\}$; the feature dimension $p\in\{100,300,500,700,900\}$; the number of classes $L\in\{2,5,8,11,14\}$; the correlation coefficient $\rho\in\{0,0.2,0.4,0.6,0.8\}$; the noise parameter $\sigma\in\{0.6,0.9,1.2,1.5,1.8\}$; the imbalance of the class probabilities $\alpha\in\{0,0.4,0.8,1.2,1.6\}$.
For all settings, we fix $n=300$, $p=500$, $L=5$, $\rho=0.6$, $\sigma=1$ and $\alpha=0$ when they are not varied. We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report the averaged misclassification errors in Figure 1.
From Figure 1, it is clear that our procedures consistently outperform other methods across all settings. We observe that the performance of our algorithms gets improved as $n$ increases, $L$ decreases and $\sigma$ decreases. Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms. These findings are in line with our theory in Section 3. It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$. Nevertheless, both our algorithms and RDA adapt to this situation. Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios. This is as expected as the data is simulated according to model (1.1) and (1.3), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
Figure 1: The averaged misclassification errors of each algorithm for varying each parameter one at a time.
5 Real data analysis
In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets that are commonly used in the literature for discriminant analysis.
The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001). The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT). These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma. Each sample has $p=2,308$ gene expressions as features. The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes. These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia. Each sample has $p=4026$ gene expressions as its features. We used the preprocessed Lymphoma data in Dettling (2004).
For all three data sets, the features are centered to zero mean. Each dataset is randomly divided with 75% of samples as the training data and 25% as the test data. Such process was repeated 10 times and the averaged misclassification errors of each algorithm were reported in Table 1.
Table 1: The averaged misclassification errors (in percentage).
LDRR PenalizedLDA PAMR RDA
Ramaswamy 14.4 37.6 28.8 15.4
SRBCT 1.0 5.2 1.4 1.0
Lymphoma 0.0 0.6 0.6 0.6
Given the similar performance between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification. From Table 1, we see that our procedure LDRR-F has the smallest misclassification errors across all three datasets. Furthermore, in Figure 2, we show that LDRR-F with $K=2$ in (2.15) can be used for visualization in the 2D space of the first two discriminant directions.
Figure 2: Visualization of Ramaswamy, SRBCT and Lymphoma data points in the space of the first two discriminant vectors.
Acknowledgements
Wegkamp is supported in part by the National Science Foundation grant NSF DMS-2210557.
Appendix A Proofs
Throughout the proofs, we will use the following notation. Define
$\Delta={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$
and write
$\Delta_{\infty}=\max_{\ell\in[L]}\Delta_{\ell\ell},\qquad\Delta_{\rm op}=\|\Delta\|_{\rm op}.$
Further write
$\pi_{\min}=\min_{\ell\in[L]}\pi_{\ell},\qquad\pi_{\max}=\max_{\ell\in[L]}\pi_{\ell}.$
The following event will be frequently used in our proofs. Define the event
$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (A.1)
for some absolute constant $C>0$. Lemma 15 in Appendix C ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1. Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (A.2)
as well as $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
A.1 Proof of the equivalence between (1.4) and (2.13)
Proof.
Recall that $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. We find that
$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}$
so that, for any $\ell\in[L]$,
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
Write the singular value decomposition of $\Sigma_{w}^{-1/2}{\mathbb{M}}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{\mathbb{M}})$. It follows that the above equals to
$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
which, in view of (1.4) and (2.1), completes the proof. ∎
A.2 Proof of Theorem 2
Proof.
We work on the intersection of the event
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (A.3)
with the event
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbb{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (A.4)
Recall $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$. The main difficulty in proving Theorem 2 is to establish the order of $\lambda$ as well as to show that $\mathcal{E}_{\lambda}\cap\mathcal{E}_{RE}$ holds with overwhelming probability.
Lemma 6 ensures that $\mathcal{E}_{\lambda}$ holds with probability at least $1-n^{-1}$ for any $\lambda$ satisfying (3.2).
We proceed to verify that $\mathbb{P}(\mathcal{E}_{RE})\geq 1-(n\vee p)^{-2}$. To this end, for any set $S\subseteq[p]$, we write $E_{S}:=\text{span}\{{\bm{e}}_{j}:j\in S\}$. Under 2, we define the space
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (A.5)
for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$. Let $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Lemma 7 and the condition $(s\vee L)\log(n\vee p)\leq cn$ imply that
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (A.6)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (A.6) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (A.7)
holds for any $u\in\mathcal{C}(S,3)$. Eq. A.7 in conjunction with 2 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof. Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
${1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
from which one can deduce that
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (A.8)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (A.7), we find on the event $\mathcal{E}_{RE}$
${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$
$\leq{3\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}$
$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (A.8)
$\leq 6\lambda\sqrt{s}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}.$
The proof is complete. ∎
A.2.1 Two key lemmas used in the proof of Theorem 2
Lemma 6.
Under model (1.1) and (1.3) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}({\mathbb{Y}}_{\ell}-{\mathbb{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{\mathbb{M}}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
Proof.
Note that
${1\over n}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (A.9)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 15 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 13 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (B.5), (B.6) and (B.7) in Lemma 9 to derive that the inequalities
$\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over n}$
$\quad+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\|{\mathbb{M}}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
hold with probability at least $1-(n\vee p)^{-3}$. After we take the union bound over $j\in[p]$ and $\ell\in[L]$, we have that
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (A.10)
holds with probability $1-n^{-1}$.
To bound the first term in (A.9), by recalling that $\Sigma_{XY}={\mathbb{M}}D_{\pi}$, we find
${1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{\mathbb{M}}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{\mathbb{M}}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){\mathbb{M}}_{j\ell}$
$={n_{\ell}\over n}(\widehat{\mathbb{M}}_{j\ell}-{\mathbb{M}}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){\mathbb{M}}_{j\ell}.$
Invoking 1, Lemma 15 and Lemma 16 with $t=\sqrt{\log(p\vee n)}$ and taking a union bound over $j\in[p]$ and $\ell\in[L]$, we obtain
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{\mathbb{M}}\|_{\infty}\sqrt{\log(n)\over nL}$ (A.11)
with probability $1-n^{-1}$. Combining (A.10) and (A.11) completes the proof. ∎
Recall the set
$\Gamma_{s}=\bigcup_{S\subseteq[p]:|S|=\lfloor Cs\rfloor}E_{S}$
with $E_{S}=\text{span}\{{\bm{e}}_{j}:j\in S\}$. The following lemma gives an upper bound for $v^{\top}(\widehat{\Sigma}-\Sigma)v$, uniformly over $v\in\Gamma_{s}$.
Lemma 7.
Under model and (1.3) with Assumptions 1 & 2, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$. Then, on the event $\mathcal{E}_{\pi}$, with probability at least $1-(n\vee p)^{-2}$, the following holds uniformly over $v\in\Gamma_{s}$,
$v^{\top}(\widehat{\Sigma}-\Sigma)v~{}\lesssim~{}v^{\top}\Sigma v\left(\sqrt{s\log(n\vee p)\over n}+\sqrt{L\log(n)\over n}\right).$
Proof.
Without loss of generality, we prove the result for
$v\in\Gamma_{s}^{\prime}:=\Gamma_{s}\cap\mathcal{S}^{p}$
via a discretization argument. For any subset $S\subseteq[p]$ with $|S|=\lfloor Cs\rfloor$, let $\mathcal{N}_{S}$ be the $(1/3)$-net of $E_{S}\cap\mathcal{S}^{p}$ satisfying
$\mathcal{N}_{S}\subset E_{S}\cap\mathcal{S}^{p},\qquad|\mathcal{N}_{S}|\leq 7^{|S|}$
The existence of such net is ensured by Rudelson and Zhou (2012, Lemma 23). Hence, for any fixed $v\in E_{S}\cap\mathcal{S}^{p}$, there exists $u\in\mathcal{N}_{S}$ such that $\|u-v\|_{2}\leq 1/3$, and we find
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\leq(v-u)^{\top}(\widehat{\Sigma}-\Sigma)v+u^{\top}(\widehat{\Sigma}-\Sigma)(v-u)+u^{\top}(\widehat{\Sigma}-\Sigma)u$
$\leq 2\|u-v\|_{2}\sup_{v\in E_{S}\cap\mathcal{S}^{p}}v^{\top}(\widehat{\Sigma}-\Sigma)v+u^{\top}(\widehat{\Sigma}-\Sigma)u.$
The second inequality uses the fact that $(u-v)/\|u-v\|_{2}\in E_{S}\cap\mathcal{S}^{p}$ whenever $u\neq v$. Taking the supremum over $v\in E_{S}\cap\mathcal{S}^{p}$ and the maximum over $u\in\mathcal{N}_{S}$ gives
$\sup_{v\in E_{S}\cap\mathcal{S}^{p}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}_{S}}u^{\top}(\widehat{\Sigma}-\Sigma)u$
and
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (A.12)
Here
$\mathcal{N}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}\mathcal{N}_{S}$
has cardinality
$|\mathcal{N}|\leq\sum_{|S|=\lfloor Cs\rfloor}|\mathcal{N}_{S}|\lesssim{p}^{Cs}.$
We will bound the right hand side (A.12) via the same arguments used to prove Lemma 13, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 13 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{\mathbb{M}}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
holds, with probability at least $1-6e^{-t}$. Note that
${\|{\mathbb{M}}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v$
under 1. By using the decomposition of $\Sigma$ in (2.6), choosing $t=Cs\log(n\vee p)$ for some large constant $C>0$, we obtain that
$v^{\top}(\widehat{\Sigma}-\Sigma)v\lesssim v^{\top}\Sigma v\left(\sqrt{s\log(n\vee p)\over n}+{s\log(n\vee p)\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$
holds uniformly over $v\in\mathcal{N}$, with probability at least $1-(n\vee p)^{-C^{\prime\prime}s}$. On the event $\mathcal{E}_{\pi}$, invoking 1, we conclude that
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\sqrt{L\log(n)\over n}$
holds uniformly over $\mathcal{N}$. The proof is complete in view of (A.12). ∎
A.3 Proof of Theorem 3
Proof.
Define
$\delta\asymp\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$ (A.13)
We work on the event
$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\infty,1}\leq\delta\right\}$ (A.14)
which, according to Lemma 8, holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{1,\infty}$ $\text{by writing }\Omega:=H^{-1}$
[…]
Hence, $\widehat{H}$ is invertible on the event $\mathcal{E}_{H}$. For simplicity, we write $\widehat{\Omega}=\widehat{H}^{-1}$. By definition,
$\widehat{B}^{*}-B^{*}$ $=(\widehat{B}-B)\widehat{\Omega}+B(\widehat{\Omega}-\Omega)$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}H(\widehat{\Omega}-\Omega)$ by Lemma 1
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})$ by $\text{by }H\Omega=\mathbb{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbb{I}_{L}.$
Recall that (A.17) and pick any $\ell\in[L]$. We it then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}_{\ell}\|_{2}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\max_{k\in[L]}\sqrt{B_{k}^{*\top}\Sigma_{w}B^{*}_{k}}~{}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\sqrt{\Delta_{\infty}}\|\widehat{H}-H\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}$ by $\text{by }B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}.$
To bound from above $\|\widehat{\Omega}_{\ell}\|_{1}$, we have that
$\|\widehat{\Omega}_{\ell}\|_{1}$ $\leq\|\Omega_{\ell}\|_{1}+\|(\widehat{\Omega}-\Omega){\bm{e}}_{\ell}\|_{1}$
$=\|\Omega_{\ell}\|_{1}+\|\Omega(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
$\leq\|\Omega_{\ell}\|_{1}+\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}.$
Since the event $\mathcal{E}_{H}$ and (3.3) ensure that
$\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\leq\left(\min_{k\in[K]}{1\over\pi_{k}}+\|\Delta\|_{1,\infty}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
we conclude, for any $\ell\in[L]$,
$\|\widehat{\Omega}_{\ell}\|_{1}\leq 2\|\Omega_{\ell}\|_{1}\lesssim L\Delta_{\infty}.$ (A.15)
It then follows that
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
$\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}+\delta\sqrt{\Delta_{\infty}}\right)L\Delta_{\infty}$
$\lesssim\Delta_{\infty}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}+\sqrt{\Delta_{\infty}^{5}}\sqrt{L\log(n)\over n}+\Delta_{\infty}^{2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}$
$\lesssim\Delta_{\infty}^{5/2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n},$ (A.16)
completing the proof. ∎
The following lemma provide upper bounds for $\|\widehat{H}-H\|_{\infty,1}$ under condition of Theorem 3.
Lemma 8.
Under conditions of Theorem 3, with probability at least $1-n^{-1}$,
$\|\widehat{H}-H\|_{\infty,1}\lesssim\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$
Proof.
Recall that
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (A.17)
We first bound the $\|\cdot\|_{\infty,1}$ norm. By triangle inequality, we have
$\|\widehat{H}-H\|_{\infty,1}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\infty,1}$
$\quad+\|B^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty,1}.$
We proceed to bound each term separately. On the event $\mathcal{E}_{\pi}$, we find that
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (A.18)
Next, pick any $k,\ell\in[L]$. We bound from above
$\textrm{I}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}(\widehat{B}_{k}-B_{k})\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2},$ (A.19)
$\textrm{II}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}B_{k}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\|\widehat{\Sigma}^{1/2}B_{k}\|_{2},$
$\textrm{III}=B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}.$
Note that
$\|\widehat{\Sigma}^{1/2}B_{k}\|_{2}\leq\left(B_{k}^{\top}\Sigma B_{k}+B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}\right)^{1/2}.$
Meanwhile Lemma 14 ensures that, on the event $\mathcal{E}_{\pi}$ in (A.1),
$\textrm{III}\leq\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}$ (A.20)
with probability at least $1-n^{-3}$. The above two bounds together with (B.5) in Lemma 9 give that
$\textrm{II}\lesssim{\sqrt{\Delta_{\infty}}\over L}\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}.$ (A.21)
By combining (A.18), (A.19), (A.20) and (A.21) and taking the union bounds over $k,\ell\in[L]$, we conclude that
$\|\widehat{H}-H\|_{\infty,1}$ $\lesssim~{}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{L\log(n)\over n}$
$\qquad+\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}\sqrt{\Delta_{\infty}}+L\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2}$
with probability at least $1-n^{-1}$. Invoking Theorem 2 and using (3.3) to simplify the expression complete the proof. ∎
A.4 Proof of Theorem 4
Proof.
Pick any $\ell\in[L]$. Start with
$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$
We work on $\mathcal{E}_{\pi}$ in (A.1) throughout the rest of the proof. For the first term in $\widehat{G}_{\ell}(X)-G_{\ell}(X)$, invoking Lemma 16 with $k=\ell$ and $v=B^{*}_{\ell}$, using the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and taking a union bounds over $\ell\in[L]$ gives that
$\mathbb{P}\left\{\left\{\max_{\ell\in[L]}|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|\leq\sqrt{\Delta_{\infty}}\sqrt{aL\log(n)\over n}\right\}\cap\mathcal{E}_{\pi}\right\}\geq 1-2n^{-a}.$ (A.22)
To bound from above
$|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|,$
using Cauchy-Schwarz inequality gives
$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\mu_{\ell}^{\top}\Sigma_{w}^{-1/2}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}.$ (A.23)
On the other hand, by using the same arguments of proving (A.3), we have that on the event $\mathcal{E}_{H}$
$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B(\widehat{\Omega}_{\ell}-\Omega_{\ell})|$
$\leq\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})_{\ell}|$
$\leq\left(\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}+\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{\infty}\delta\right)2\|\Omega_{\ell}\|_{1}.$
Here $\delta$ is defined in (A.13). Invoking (A.15), (A.22), (A.23) and Lemma 16 with $v={\bm{e}}_{j}$ and union bounds over $j\in[p]$ yields that, on $\mathcal{E}_{H}$ defined in (A.14),
$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $~{}\lesssim L\Delta_{\infty}\sqrt{\|\Sigma_{w}\|_{\infty}}\sqrt{L\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}$
$\qquad+L\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L\log(n)\over n}$
with probability at least $1-4(n\vee p)^{-2}$. Therefore, with probability at least $1-5(n\vee p)^{-2}$,
$\begin{split}|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|&~{}\lesssim~{}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}+\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L^{3}\log(n)\over n}\\ &\qquad+\Delta_{\infty}\sqrt{L^{3}\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}.\end{split}$ (A.24)
Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\mathbb{X}},{\mathbb{Y}}$ gives that, for all $t\geq 0$,
$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\mathbb{X}},{\mathbb{Y}}\right\}\leq 2e^{-t^{2}/2}.$
Choosing $t=\sqrt{4a\log(n)}$, taking the union bounds over $\ell\in[L]$ and unconditioning yield
$\max_{\ell\in[L]}|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\geq\max_{k,\ell\in[L]}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{4a\log(n)}$
$\geq 3\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{2,\infty}\sqrt{(a\vee\Delta_{\infty})\log(n)}$ (A.25)
with probability less than $2n^{-a}$. The last step uses (A.23). Collecting (A.2), (A.22), (A.24) and (A.4) and invoking Theorem 3 complete the proof. ∎
A.5 Proof of Theorem 5
For any $x\in\mathbb{R}^{p}$, write
$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
For any $t\geq 0$, define the event
$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t~{}\Big{|}~{}{\mathbb{X}},{\mathbb{Y}}\right\}.$ (A.26)
Proof.
By definition, we start with
$\quad\mathcal{R}(\widehat{g})$
$=\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]\Bigr{\}}$
$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y={\bm{e}}_{k}\right]$
$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid L=\ell\right]\Bigr{\}}.$
Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
[…]
where, for all $k,\ell\in[L]$,
$G^{(\ell|k)}(x):=\left(x-{\mu_{\ell}+\mu_{k}\over 2}\right)^{\top}\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{k})+\log{\pi_{\ell}\over\pi_{k}}={1\over 2}\left(G_{k}(x)-G_{\ell}(x)\right).$
Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
$00>G_{z}^{(\ell|k)}(z)\overset{\mathcal{E}_{t}}{\geq}{1\over 2}\left(\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)\right)-t\geq-t,\quad\forall~{}t>$ 0 .
By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that, for any $t>0$,
$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\left\{t\pi_{k}\mathbb{E}\left[\mathbbm{1}\{-t\leq G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y={\bm{e}}_{k}\}\right]+\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
$\leq t\sum_{k\in[L]}\pi_{k}\left\{\mathbb{P}\left(-t\leq\max_{\ell\in[L]\setminus\{k\}}G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right)+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
$\leq t(m-1)\sum_{k\in[L]}\pi_{k}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c})$
Note that
$G^{(\ell|k)}(X)\mid Y={\bm{e}}_{k}~{}\sim~{}N\left(-{1\over 2}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}+\log\left(\pi_{\ell}\over\pi_{k}\right),~{}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}\right).$
By the mean-value theorem, we conclude
$\mathcal{R}(\widehat{g})$ $\leq(m-1)t^{2}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$
Choosing
$t=\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}$
and invoking Theorem 4 complete the proof. ∎
Appendix B Technical lemmas
The following lemma provides a few useful facts on quantities related with $B$ in (1.13). Write
$\Omega:=D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}.$ (B.1)
Lemma 9.
With $B$ defined in (1.13), we have
${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1},$ (B.2)
$H=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\Omega^{-1},$ (B.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1}.$ (B.4)
Consequently, for any $k,\ell\in[L]$, we have
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (B.5)
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (B.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (B.7)
Proof.
From Lemma 1 and its proof, we know that
$B=B^{*}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)=\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)$
so that
${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}-{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
Rearranging terms gives
${\mathbb{M}}^{\top}B$ $=\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}$
$=\mathbb{I}_{L}-\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
$=\mathbb{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$ (B.8)
$={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$
proving (B.2). By using the above identity (B.8), we have
$D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1},$
proving (B.3). Since Lemma 1 and (B.3) imply
$B=B^{*}\Omega^{-1},$
(B.4) follows from
$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\Omega^{-1}.$
To prove (B.5) – (B.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (B.9)
yielding the second bound in (B.5). The other bound in (B.5) follows by observing that
[…]
For (B.6), on the one hand, similar arguments yield
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
On the other hand, using (B.2) and (B.3) proves
$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
Finally, the last statement follows by noting that
$|\mu_{k}^{\top}B_{\ell}|^{2}\leq B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq 1$
and
$|\mu_{k}^{\top}B_{\ell}|=\pi_{\ell}|\mu_{k}^{\top}\Sigma^{-1}\mu_{\ell}|\leq\pi_{\ell}\sqrt{\mu_{k}^{\top}\Sigma^{-1}\mu_{k}}\sqrt{\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}}.$
The following lemmas contain deviation inequalities related with $\widehat{\mathbb{M}}-{\mathbb{M}}$ and $\widehat{\pi}-\pi$.
Lemma 10.
Let $v\in\mathbb{R}^{p}$ and $u\in\mathbb{R}^{L}$ be any fixed vectors. Under (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t^{2}/2}$, we have
$\left|\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}\right|\leq t\|u\|_{2}\sqrt{v^{\top}\Sigma_{w}v\over n}\sqrt{\max_{k}n_{k}\over n}.$
Proof.
Recall that, for any $k\in[L]$,
$\widehat{\mu}_{k}={1\over n_{k}}\sum_{i=1}^{n}X_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}.$
We obtain
$\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}$ $=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-\mu_{k}\right)u_{k}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}.$
For $i\in[n]$, by conditioning on ${\mathbb{Y}}$, we know that $W_{i}:=X_{i}-{\mathbb{M}}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbb{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}u\over n}\right).$
The claim follows by invoking the standard Gaussian tail probability bounds and after unconditioning. ∎
Lemma 11.
Let $u,v\in\mathbb{R}^{L}$ be any fixed vectors. Under (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t/2}$,
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v_{k}\right|\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}u_{k}^{2}v_{k}^{2}\over n}+{t\|u\|_{\infty}\|v\|_{\infty}\over n}.$
Proof.
Observe that we can write
$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v_{k}$ $=\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}Y_{ik}u_{k}v_{k}-\sum_{k=1}^{L}\pi_{k}u_{k}v_{k}$
$={1\over n}\sum_{i=1}^{n}Z_{i}-\pi_{k}u_{k}v_{k}$
Here
$Z_{i}=\sum_{k=1}^{L}Y_{ik}u_{k}v_{k}$
are independent, bounded random variables (as $|Z_{i}|\leq\|u\|_{\infty}\|v\|_{\infty}$) with mean
$\mathbb{E}[Z_{i}]=\sum_{k=1}^{L}\pi_{k}u_{k}v_{k}$
and variance
${\rm Var}(Z_{i})={\rm Var}(y_{i}^{\top}{\rm diag}(u)v)\leq\sum_{k=1}^{L}\pi_{k}v_{k}^{2}u_{k}^{2}$
for $i\in[n]$. The proof follows after a straightforward application of Bernstein’s inequality for bounded random variables. ∎
The following lemma controls the difference
$\widehat{\Sigma}_{w}-\Sigma_{w}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}.$
Lemma 12.
Let $u,v\in\mathbb{R}^{p}$ be any fixed vectors. Under model (1.1), for any $t\geq 0$, with probability at least $1-2e^{-t}$, one has
$\left|u^{\top}\left(\widehat{\Sigma}_{w}-\Sigma_{w}\right)v\right|\lesssim\sqrt{u^{\top}\Sigma_{w}u}\sqrt{v^{\top}\Sigma_{w}v}\left(\sqrt{t\over n}+{t\over n}\right)$
Proof.
Start with
$\widehat{\Sigma}_{w}-\Sigma_{w}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}$
$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{\mathbb{M}}Y_{i})(X_{i}-{\mathbb{M}}Y_{i})^{\top}-\Sigma_{w}\right].$
Recall that, conditioning ${\mathbb{Y}}$, $W_{i}=X_{i}-{\mathbb{M}}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$. The result follows from standard concentration inequalities for the quadratic term of Gaussian random vectors. ∎
B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (A.1).
Lemma 13.
Under model (1.1) and 1,assume $L\log(n)\leq n$. On the event $\mathcal{E}_{\pi}$, for any fixed $v_{1},v_{2}\in\mathbb{R}^{p}$ and any $t\geq 0$, with probability at least $1-8e^{-t}$, one has
$\left|v_{1}^{\top}(\widehat{\Sigma}-\Sigma)v_{2}\right|$ $\lesssim~{}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}\left(\sqrt{t\over n}+{t\over n}\right)$
$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$
$\quad+\left({\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
Proof.
By definition, we have
$\widehat{\Sigma}$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
$={1\over n}\sum_{k=1}^{L}\left[\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}+n_{k}(\widehat{\mu}_{k}\mu_{k}^{\top}+\mu_{k}\widehat{\mu}_{k}^{\top})-n_{k}\mu_{k}\mu_{k}^{\top}\right]$
$=\widehat{\Sigma}_{w}+\sum_{k=1}^{L}{n_{k}\over n}\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}{n_{k}\over n}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right]$
where we write
$\widehat{\Sigma}_{w}:={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}.$
By further using the decomposition of $\Sigma$ in (2.6) and $\widehat{\pi}_{k}=n_{k}/n$ for $k\in[L]$, we find that
$\widehat{\Sigma}-\Sigma$ $=\widehat{\Sigma}_{w}-\Sigma_{w}+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}\widehat{\pi}_{k}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right].$
The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 12. We bound the second term by invoking 1 and Lemma 11 with $u={\mathbb{M}}^{\top}v_{1}$ and $v={\mathbb{M}}^{\top}v_{2}$. This yields
$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}\over n}$
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$ (B.11)
with probability at least $1-2e^{-t}$. Furthermore, on the event $\mathcal{E}_{\pi}$, 1 and $L\log(n)\leq n$ imply
$\widehat{\pi}_{k}\lesssim\pi_{k}+\sqrt{\log(n)\over nL}\lesssim{1\over L},\quad\forall\ k\in[L],$
and after we invoke Lemma 10 twice with $v=v_{1},u={\mathbb{M}}^{\top}v_{2}$ and $v=v_{2},u={\mathbb{M}}^{\top}v_{1}$, respectively, we get that
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (B.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (B.13)
hold with probability at least $1-4e^{-t}$. Collecting all three bounds completes the proof. ∎
As an immediate application of Lemma 13, we have the following bounds on the sup-norm and operator norm of the matrix $B^{\top}(\widehat{\Sigma}-\Sigma)B$.
Lemma 14.
Under conditions of Lemma 13, assume $\Delta_{\infty}\gtrsim 1$. On the event $\mathcal{E}_{\pi}$, the following holds with probability at least $1-n^{-2}$,
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL},$
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$
Proof.
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 13 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (B.5), (B.6) and (B.7) of Lemma 9 gives that, with probability at least $1-n^{-2}$,
$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\quad+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
$\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)\sqrt{\log(n)\over nL^{2}}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{2}{\log(n)\over n}$
$\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}.$
The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim. Regarding the operator norm bound, a standard discretization argument gives
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\leq 3\max_{u\in\mathcal{N}_{L}(1/2)}u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$
where $\mathcal{N}_{L}(1/3)$ denotes the $(1/3)$-epsilon net of $\{u\in\mathbb{R}^{L}:\|u\|_{2}=1\}$. Note that $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$. Then invoking Lemma 13 with $t=CL\log(n)$ and using
$\|B^{\top}\Sigma_{w}B\|_{\rm op}\leq{\Delta_{\rm op}\over L(L+\Delta_{\rm op})},$
$\|B^{\top}{\mathbb{M}}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
$\|{\mathbb{M}}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
deduced from Lemma 9 give that, with probability at least $1-8n^{-CL+L\log(7)}\geq 1-n^{-L}$,
$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{\mathbb{M}}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
$\quad+\|{\mathbb{M}}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
$\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}$
holds uniformly over $u\in\mathcal{N}_{L}(1/3)$. We also use $\Delta_{\infty}\gtrsim 1$ to simplify expressions in the last step above. This completes the proof. ∎
Appendix C Auxiliary lemmas
The following lemmas are proved in Bing and Wegkamp (2023).
Lemma 15.
Assume $\pi_{\min}\geq 2\log n/n$ for some sufficiently large constant $C$. Then, for any $k\in[L]$
$\mathbb{P}\left\{|\widehat{\pi}_{k}-\pi_{k}|<\sqrt{16\pi_{k}(1-\pi_{k})\log n\over n}\right\}\geq 1-n^{-2}.$
Furthermore, if $\pi_{\min}\geq C\log n/n$ for some sufficiently large constant $C$, then
$\mathbb{P}\left\{c\pi_{k}\leq\widehat{\pi}_{k}\leq c^{\prime}\pi_{k}\right\}\geq 1-n^{-2}.$
Lemma 16.
Suppose that model (1.1) holds. For any deterministic vector $v\in\mathbb{R}^{p}$ and $k\in[L]$, for all $t>0$,
$\mathbb{P}\left\{\left|v^{\top}(\widehat{\mu}_{k}-\mu_{k})\right|\geq t\sqrt{v^{\top}\Sigma_{w}v\over n_{k}}\right\}\leq 2e^{-{t^{2}/2}}.$

### v2

@@
 This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods.
-Moreover, our new formulation is generically easy to analyze compared to existing regression-based LDA procedures.
-In particular, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization that has not yet been fully analyzed in the LDA context.
+Moreover, our new formulation is amenable to analysis: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques.
+As applications, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization as well as for using the reduced-rank regression, neither of which has yet been fully analyzed in the LDA context.
 Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
@@
 Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem.
-Since we can always subtract the marginal mean of $X$, we assume
-$\mathbb{E}(X)=0_{p}$ (1.2)
-for simplicity, and furthermore that the class probabilities
-$\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ (1.3)
-are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) and (1.3) is to classify a new feature $x\in\mathbb{R}^{p}$ as
+Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$ throughout this paper mainly for simplicity, and furthermore that the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ are strictly positive for all $\ell\in[L]$.
+The Bayes rule under model (1.1) is to classify a new feature $x\in\mathbb{R}^{p}$ as
 $g^{*}(x):=\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.2)
@@
 Estimation of this classifier based on $(X_{i},Y_{i})$ with $i\in[n]$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
-Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). Write
-$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\qquad D_{\pi}={\rm diag}(\pi),\qquad{\mathbb{M}}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.5)
+Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022).
+We recall various approaches to deal with (i).
+We will use the following notation.
+Write
+$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\quad D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L},\quad{M}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.3)
 and denote the unconditional covariance matrix of $X$ by
-$\Sigma={\rm Cov}(X)=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}.$
-We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}$ the between-class covariance matrix.
-The latter due to (1.2) has rank less than $L$.
-In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure to estimate this direction (see, also, Cai and Zhang (2019) for extension to $L>2$).
-It isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed later.
-Witten and Tibshirani (2011) use the fact that the Bayes rule (1.4) is equivalent with Fisher’s discriminant rule
+$\Sigma={\rm Cov}(X)=\Sigma_{w}+{M}D_{\pi}{M}^{\top}.$ (1.4)
+We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={M}D_{\pi}{M}^{\top}$ the between-class covariance matrix.
+The latter has rank less than $L$ due to $\mathbb{E}(X)=0_{p}$.
+In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$ to estimate this direction.
+Other procedures based on the same or similar equations include, for instance, Qiao et al.
+(2009); Shao et al.
+(2011); Fan et al.
+(2012) and Tibshirani et al.
+(2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022) for $L>2$.
+However, it isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed below shortly.
+Witten and Tibshirani (2011) alternatively use the fact that the Bayes rule (1.2) is equivalent with Fisher’s discriminant rule
 $\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.5)
 Here the columns of the matrix $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are implicitly defined via, for $1\leq k\leq K$,
-$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}&~{}\beta^{\top}\Sigma_{b}\beta\\ \text{such that}&\quad\beta^{\top}\Sigma_{w}\beta=1,\quad\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.7)
+$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~{}\beta^{\top}\Sigma_{b}\beta\quad\text{subject to}&~{}\beta^{\top}\Sigma_{w}\beta=1,~{}\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.6)
 Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$.
-However, the implicit formulation (1.7) of the matrix $F$ makes it extremely difficult to obtain the global solution and to analyze the theoretical properties of the regularized estimator, hence the resulting classifier.
-More closely related to our approach, introduced in the next paragraph, is optimal scoring, which dates back to De Leeuw et al.
+However, the implicit formulation (1.6) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
+More closely related to our approach is optimal scoring which dates back to De Leeuw et al.
 (1976); Young et al.
@@
 By writing the response matrix and feature matrix as
-${\mathbb{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbb{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.8)
+${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.7)
 optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
-$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}\\ \text{such that }&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.9)
+$\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k.\end{split}$ (1.8)
 Here $\widehat{\theta}_{i}$ is the solution to (1.8) in the $i$th iteration.
-The minimization problem (1.9) can be done via solving a generalized eigenvalue problem, see Hastie et al.
-(1995).
-Moreover, Hastie et al.
-(1995) also show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.9) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.7) in the low-dimensional case ($p<n$).
+In the low-dimensional case ($p<n$), Hastie et al.
+(1995) show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.8) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.6).
 As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$.
-Since optimal scoring is a regression-based approach, as pointed out in Hastie et al.
-(1995, 1994), it is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the procedure of Cai and Zhang (2019).
-In the high-dimensional setting, Hastie et al. (1995) show that the global solution to a penalized version of (1.9), that is, for $1\leq k\leq K<L$,
-$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbb{Y}}\theta-{\mathbb{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\\ \text{such that}&~{}\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\theta=n,\quad\theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.10)
-can be computed for any $p\times p$ symmetric, positive definite matrix $\Omega$.
-A possible choice of $\Omega=\lambda\mathbb{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989).
-In a more recent work (Gaynanova, 2020), the matrix formulation of (1.10)
-$\begin{split}\min_{\Theta,~{}B}&~{}\|{\mathbb{Y}}\Theta-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B)\\ \text{such that }&~{}\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}\Theta=n\mathbb{I}_{K},\quad\Theta^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}1_{n}=0_{K}.\end{split}$ (1.11)
-is considered, and it is proved that the global solution $\widehat{B}\in\mathbb{R}^{p\times K}$ to (1.11) can be computed for any penalty $\text{pen}(B)$ that satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$.
-This allows for the penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ used in the group-lasso, but excludes the familiar $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ used in the lasso regression as well as its extensions such as the elastic net penalty.
+In the high-dimensional setting, regularization in (1.8) is needed.
+For instance, Hastie et al.
+(1995) consider
+$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.9)
+for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution to (1.9) can be computed.
+A possible choice of $\Omega=\lambda\mathbf{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989).
+To accommodate more general regularization, Gaynanova (2020) studies the matrix formulation of (1.9),
+$\begin{split}\min_{\Theta,\ B}&~{}\|{\mathbf{Y}}\Theta-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B)\quad\text{s.t.}~{}\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\Theta=n\mathbf{I}_{K},\quad\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}1_{n}=0_{K},\end{split}$ (1.10)
+and proves that the global solution to (1.10) can be computed, provided pen$(B)$ satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$.
+Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ and analyzes the resulting classifier.
+However, it is not clear how the analysis can be extended to other penalized estimation.
+Moreover, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ as well as its extensions such as the elastic net penalty.
 That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.10), let alone any theoretical property of the resulting estimator.
-Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1) and (1.3)
-$B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}\in\mathbb{R}^{p\times L}$ (1.12)
+Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1)
+$B^{*}=\Sigma_{w}^{-1}{M}\in\mathbb{R}^{p\times L}$ (1.11)
 and the regression matrix
-$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)$ (1.13)
+$B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)\in\mathbb{R}^{p\times L}$ (1.12)
 not only share the same column space, but have a closed-form connection, to wit,
-$B^{*}$ $=BH^{-1}$
-for some invertible $L\times L$ matrix $H$.
+$B^{*}$ $=BH^{-1}$ (1.13)
+for some invertible $L\times L$ matrix $H$ with explicit expression.
 This connection is novel and distinguishes our approach from the current literature.
-Although the fact that $B$ and $B^{*}$ share the same span is relatively well-known (see, for instance, Hastie et al.
+Although the fact that columns of $B$ and $B^{*}$ share the same span is known (see, for instance, Hastie et al.
 (1995); Ye (2007); Lee and Kim (2015); Nie et al.
-(2022) and the references therein), their closed-form connection has not been discovered and a thorough theoretical investigation on the corresponding classifier has not been conducted, to the best of our knowledge.
-Specifically, we propose to estimate the discriminant matrix $B^{*}$ by estimating the regression matrix $B$ first
-$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbb{Y}}-{\mathbb{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
-and then estimating $H^{-1}$ based on $\widehat{B}$. The advantage of our approach is three-fold:
+(2022) and the references therein), their closed-form connection given by (1.13) has not been discovered, and a thorough theoretical investigation on the corresponding classifier has not been conducted.
+To the best of our knowledge, only in the binary case ($L=2$), Mai et al.
+(2012) leverages sparse linear regression between an appropriately encoded label vector and the feature matrix ${\mathbf{X}}$ for classification and proves consistency.
+Extension of such procedure to handle multi-class responses with high-dimensional features has been a longstanding open problem.
+In view of (1.13), we propose to estimate the discriminant direction matrix $B^{*}$ by first estimating the regression matrix $B$ as
+$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
+and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are three-fold:
   1. (a)
-    We can easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$.
+    Akin to optimal scoring, our approach in (1.14) is regression-based, which, as pointed out in Hastie et al.
+(1995, 1994), is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the aforementioned procedures based on approximately solving the equation $\Sigma_{w}B^{*}=M$ for $B^{*}$.
+We can also easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$ in (1.14).
   2. (b)
-    After estimating $B$, we only need to estimate the inverse of an $L\times L$ matrix rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
+    After computing $\widehat{B}$ from (1.14), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
+Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys computational advantage as well as numerical stability, compared to the existing approaches.
+In Section 2.3, we further offer an alternative way to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$.
+This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization.
   3. (c)
-    We are able to quantify the estimation error of both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier.
-While various penalties are feasible in our approach, more specifically in (1.14), we focus on a lasso-type estimator of $B$ in this paper.
-This also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively.
-Our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ and unobserved noise $W$, that is independent of both $Z$ and the label $Y$ with $r<n<p$.
-The procedure is entirely different than the one proposed here.
-Future work will be devoted on how to adapt the procedure in this paper to this low-rank setting and its analysis.
-Under sparsity of the regression matrix $B$, we carefully analyze the properties of $\widehat{B}$ in Section 3.1, the estimation of $B^{*}$ via estimating $H^{-1}$ in Section 3.2 as well as the rate of convergence for the excess misclassification risk of the corresponding classification rule in Section 3.3.
-Computationally, after computing $\widehat{B}$ from (1.14) we offer an alternative way in Section 2.3 to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$.
-This leads to a practically useful method, which extracts fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization.
-Finally, we provide a thorough simulation study in Section 4 and we report our findings in several real data studies of Section 5.
+    We are able to quantify the errors of estimating both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier, thanks to (1.13).
+Concretely, in Section 3.1 and Section 3.2 we provide a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for a generic estimator $\widehat{B}$ of $B$.
+While various choices of $\widehat{B}$ are feasible in our approach, more specifically in (1.14), we apply the general result to two popular estimators: the lasso estimator in Section 3.3 and the reduced rank estimator in Section 3.4.
+The former also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively.
+The latter is particularly suitable when the rank of the discriminant direction matrix $B^{*}$ is low and the number of classes $L$ is moderate / large.
+For instance, our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ with $r<p$ and random noise $W$, that is independent of both $Z$ and the label $Y$.
+In this case, the discriminant direction matrix $B^{*}$ has rank no greater than $r$ when $r<L$.
+The procedure in Bing and Wegkamp (2023), however, is different than the one proposed here.
+We refer to Remark 1 in Section 2.2 for more discussion.
+Finally, we provide a thorough simulation study in Section 4 to corroborate our theoretical findings.
+In Section 5, we report our findings in several real data studies.
+The Appendix contains most of the proofs and additional simulation study.
+Notation.
+For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$.
+For any vector $v$ and $1\leq q\leq\infty$, we use $\|v\|_{q}$ to denote the standard $\ell_{q}$-norm.
+For any positive integer $d$, we write $[d]:=\{1,\ldots,d\}$.
+For any matrix $Q\in\mathbb{R}^{d_{1}\times d_{2}}$, any $i\in[d_{1}]$ and $j\in[d_{2}]$, we write $Q_{i\cdot}$ for its $i$th row and $Q_{j}$ for its $j$th column.
+For norms, we write $\|Q\|_{\infty}=\max_{i,j}|Q_{ij}|$, $\|Q\|_{F}^{2}=\sum_{i,j}Q_{ij}^{2}$, $\|Q\|_{1}=\sum_{i,j}|Q_{ij}|$, $\|Q\|_{1,2}=\sum_{i}\|Q_{i\cdot}\|_{2}$ and $\|Q\|_{\rm op}=\sup_{v:\|v\|_{2}=1}\|Qv\|_{2}$.
+For any symmetric, semi-positive definite matrix $Q\in\mathbb{R}^{d\times d}$, we use $\lambda_{1}(Q),\lambda_{2}(Q),\ldots,\lambda_{d}(Q)$ to denote its eigenvalues in non-increasing order.
+For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
 2 Methodology
@@
 $G_{\ell}(x):=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}),\qquad\text{for all }\ell\in[L].$ (2.1)
-This, in turn, requires that we estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$.
+This, in turn, requires to estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{M}$.
 The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.12).
-Recall the definitions of $\pi$, $D_{\pi}$ and ${\mathbb{M}}$ from (1.5) as well as $\Sigma={\rm Cov}(X)$.
+Recall the definitions of $\pi$, $D_{\pi}$ and ${M}$ from (1.3) as well as $\Sigma$ from (1.4).
 Lemma 1.
-Under model (1.1) and (1.3), we have
+Under model (1.1), we have
 $B^{*}=BH^{-1}$ (2.2)
 for a matrix $H\in\mathbb{R}^{L\times L}$ that can be inverted and satisfies
-$H=D_{\pi}-B^{\top}\Sigma B=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=D_{\pi}-B^{\top}{\mathbb{M}}D_{\pi}.$ (2.3)
+$H=D_{\pi}-B^{\top}\Sigma B=D_{\pi}-D_{\pi}{M}^{\top}B=D_{\pi}-B^{\top}{M}D_{\pi}.$ (2.3)
 Proof.
 From the identity
-$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={\mathbb{M}}D_{\pi}$ (2.4)
+$\Sigma_{XY}=\mathbb{E}(XY^{\top})-\mathbb{E}(X)(\mathbb{E}(Y))^{\top}={M}D_{\pi}$ (2.4)
 we see that
-$B=\Sigma^{-1}{\mathbb{M}}D_{\pi}.$ (2.5)
+$B=\Sigma^{-1}{M}D_{\pi}.$ (2.5)
 Next, using
-$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top},$ (2.6)
+$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
 the Woodbury matrix identity yields
-$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}.$
+$\Sigma_{w}^{-1}=\Sigma^{-1}+\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}D_{\pi}{M}^{\top}\Sigma^{-1}.$
 It follows that
-$B^{*}$ $=\Sigma_{w}^{-1}{\mathbb{M}}$
-$=\Sigma^{-1}{\mathbb{M}}+\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}$
-$=\Sigma^{-1}{\mathbb{M}}\left(\mathbb{I}_{L}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}\right)^{-1}$
-$=\Sigma^{-1}{\mathbb{M}}D_{\pi}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
-$=B\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)^{-1}$
+$B^{*}$ $=\Sigma_{w}^{-1}{M}$
+$=\Sigma^{-1}{M}+\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}D_{\pi}{M}^{\top}\Sigma^{-1}{M}$
+$=\Sigma^{-1}{M}\left(\mathbf{I}_{L}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}\right)^{-1}$
+$=\Sigma^{-1}{M}D_{\pi}\left(D_{\pi}-D_{\pi}{M}^{\top}\Sigma^{-1}{M}D_{\pi}\right)^{-1}$
+$=B\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)^{-1}$
 which completes the proof. ∎
@@
 2.1 Estimation of the regression matrix
-In view of (1.13), it is natural to estimate $B$ via regularized regression methods in (1.14).
+To accommodate high-dimensional data, we estimate $B$ via regularized regression methods in (1.14).
 Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand.
-For example, if we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions.
+For instance, when entries of $B$ exhibit certain smoothness structure such as in some imaging or audio applications (Hastie et al., 1995), one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$.
+In this paper we mainly focus on the following two types of structural estimation of $B$.
+Sparsity-based structural estimation.
+If we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions.
 If the subset of features, that are pertinent for predicting the label, varies across different levels, then the entry-wise lasso penalty is appropriate.
-To fix ideas and due to technical complications, we focus our analysis to the matrix estimator $\widehat{B}$ in (1.14) obtained with the following concrete lasso penalty
+The corresponding matrix estimator $\widehat{B}$ is defined in (1.14) with the following lasso penalty (Tibshirani, 1996)
 $\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}$ (2.7)
-for some tuning parameter $\lambda>0$. It is fairly straightforward to modify our analysis to other choices of the penalty term discussed below.
-Remark 1 (Alternative penalty choices).
-If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ is a more suitable choice.
+for some tuning parameter $\lambda>0$.
+In Section 3.3 we provide a complete analysis of this estimator.
+If there exist many features that are not predictive for all levels of the response label, then the group-lasso penalty $\text{pen}(B)=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (Yuan and Lin, 2006) is a more suitable choice.
 In this case, one can retain a much smaller subset of important features and achieve feature selection.
-Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty if the features tend to (highly) correlate with each other.
-Another example that is commonly used in the literature is when entries of $B$ exhibit certain smoothness structure, such as some imaging or audio applications in Hastie et al.
-(1995).
-In this case, one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$.
-On the other hand, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate.
-For example, if the matrix of conditional means ${\mathbb{M}}\in\mathbb{R}^{p\times L}$ has a reduced rank $r\ll\min\{p,L\}$ such as under the model in Bing and Wegkamp (2023), then Lemma 1 implies that ${\rm rank}(B)\leq r$.
-Estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using either the reduced rank penalty $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ or the nuclear norm penalty $\text{pen}(B)=\lambda\|B\|_{*}$.
-We will investigate this approach in another paper.
+Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty (Zou and Hastie, 2005) if the features tend to (highly) correlate with each other.
+Our numerical results in Sections 4 and 5 include both the lasso penalty and the elastic net penalty.
+Low-rank structural estimation.
+Different from sparsity, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.14) could be more appropriate.
+For example, if the matrix of conditional means ${M}\in\mathbb{R}^{p\times L}$ has a reduced-rank $r\ll\min\{p,L\}$, then Lemma 1 implies that ${\rm rank}(B)\leq r$.
+In this case estimation of $B\in\mathbb{R}^{p\times L}$ in (1.14) could be done by using the following reduced-rank penalty (see, for instance, Izenman (1975))
+$\text{pen}(B)=\lambda~{}{\rm rank}(B)$ (2.8)
+for some tuning parameter $\lambda>0$.
+A full theoretical analysis of using (2.8) is presented in Section 3.4.
+Some alternative penalties include, but are not limited to, the reduced-rank penalty plus the matrix ridge penalty in Mukherjee and Zhu (2011), the nuclear norm penalty in Koltchinskii et al.
+(2011), adaptive nuclear norm penalty in Chen et al.
+(2013) as well as the generalized cross-validation rank penalty in Bing and Wegkamp (2019).
+Remark 1.
+We conclude this section by giving another concrete example under which $B$ has low rank.
+In Bing and Wegkamp (2023) the authors consider the model $X=AZ+W$ with latent factors $Z\in\mathbb{R}^{r}$ for some $r\ll p$ and some general additive noise $W$, that is independent of both $Z$ and the label $Y$.
+Conditioning on $Y={\bm{e}}_{\ell}$ for all $\ell\in[L]$, the low-dimensional factors $Z$ are assumed to follow Gaussian distributions with different conditional means $\alpha_{\ell}\in\mathbb{R}^{r}$ but the same within-class covariance matrix $\Sigma_{z}\in\mathbb{R}^{r\times r}$.
+When $W$ is also Gaussian, this model is a sub-model of (1.1) with $\mu_{\ell}=A\alpha_{\ell}$ and $\Sigma_{w}=A\Sigma_{z}A^{\top}+{\rm Cov}(W)$.
+When $r\leq\min\{p,L\}$, it is easy to verify that ${\rm rank}(B)\leq r$.
+The two-step approach in Bing and Wegkamp (2023) uses the unsupervised Principal Component Analysis (PCA) to reduce the feature dimension and then base classification on the reduced dimension.
+It is entirely different from our approach in Section 2.2 but connected to the one in Section 2.3 in the way that our approach in Section 2.3 is also a two-step procedure but reduces the feature space in the first step in a supervised way via the regression step in (1.14).
+Numerical comparison of both procedures are also included in our simulation study in Appendix A.
 2.2 Estimation of the discriminant matrix and the Bayes rule
-The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.4) necessitate that we estimate the matrix of the conditional means ${\mathbb{M}}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$.
+The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.2) necessitate that we estimate the matrix of the conditional means ${M}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$.
 We adopt the following standard estimators
-$\widehat{\mathbb{M}}={\mathbb{X}}^{\top}{\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbb{Y}}^{\top}{\mathbb{Y}}$ (2.8)
-to estimate ${\mathbb{M}}$ and $D_{\pi}$, respectively.
+$\widehat{M}={\mathbf{X}}^{\top}{\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbf{Y}}^{\top}{\mathbf{Y}}$ (2.9)
+to estimate ${M}$ and $D_{\pi}$, respectively.
 Recall from Lemma 1 that $B^{*}=BH^{-1}$.
 Since the estimator $\widehat{B}$ of $B$ is computed in (1.14), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
-$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}={1\over n}\left({\mathbb{Y}}^{\top}{\mathbb{Y}}-\widehat{B}^{\top}{\mathbb{X}}^{\top}{\mathbb{X}}\widehat{B}\right).$ (2.9)
-Note that ${\mathbb{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14).
-Computation of ${\mathbb{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbb{X}}$ consist of basis expansions.
+$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}={1\over n}\left({\mathbf{Y}}^{\top}{\mathbf{Y}}-\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}\right).$ (2.10)
+Note that ${\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14).
+Computation of ${\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbf{X}}$ consist of basis expansions.
 Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
 $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}.$ (2.11)
-Here $M^{+}$ denotes the moore-penrose inverse of any matrix $M$. Finally, in view of (2.1), we estimate the discriminant functions by
-$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell})\quad\text{for all $\ell\in[L]$. }$ (2.11)
+Here for any matrix $M$, $M^{+}$ denotes its moore-penrose inverse. Finally, in view of (2.1), we estimate the discriminant functions by
+$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$. }$ (2.12)
 The corresponding classifier is
 $\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad\text{ for any $x\in\mathbb{R}^{p}$}.$ (2.13)
-Theoretically, we will show that $\widehat{H}$ is invertible with overwhelming probability for the lasso-penalty (see, Remark 3 as well as the proof of Theorem 3 in Section A.3).
+Theoretically, we show in Theorem 5 of Section 3.2 that $\widehat{H}$ is invertible with overwhelming probability under the same mild condition that is required for consistent classification.
+In Theorem 6 of Section 3.3 and Theorem 8 of Section 3.4, the condition is verified for the lasso estimator, and the reduced rank estimator, of $B$, respectively.
 Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.14).
@@
 Since in practice it is often desirable to extract fewer discriminant directions than those in $B^{*}$, we discuss in this section an alternative procedure for this purpose, which estimates the Bayes rule by also using $\widehat{B}$ from (1.14) but in a different way.
-We start with the fact (its proof is stated in Section A.1 for completeness) that the Bayes rule in (1.4) is equivalent to
+We start with the fact (its proof is stated in Section B.1 for completeness) that the Bayes rule in (1.2) is equivalent to
 $\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$ (2.14)
@@
 Write the between-class covariance matrix of $Z=B^{\top}X$ as
-$C_{b}:=B^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
-For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the minimization problem
-$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\\ \text{such that }&~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\text{ for all }~{}i<k.\end{split}$
+$C_{b}:=B^{\top}{M}D_{\pi}{M}^{\top}B.$
+For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the optimization problem
+$\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}C_{b}~{}\alpha\quad\text{ s.t.}~{}\alpha^{\top}C_{w}~{}\alpha=1,~{}\alpha^{\top}C_{w}~{}\alpha_{i}=0,~{}~{}\forall~{}i<k.\end{split}$
 The resulting classification rule based on these $K$ discriminant directions is
@@
 In practice, after computing $\widehat{B}$ from (1.14), we propose to estimate $C_{b}$ and $C_{w}$ by
-$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{\mathbb{M}}D_{\widehat{\pi}}\widehat{\mathbb{M}}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}P_{{\mathbb{Y}}}{\mathbb{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbb{X}}^{\top}(\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B},$
-with $P_{{\mathbb{Y}}}={\mathbb{Y}}({\mathbb{Y}}^{\top}{\mathbb{Y}})^{-1}{\mathbb{Y}}^{\top}$.
+$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{M}D_{\widehat{\pi}}\widehat{M}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}P_{{\mathbf{Y}}}{\mathbf{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}(\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B},$
+with $P_{{\mathbf{Y}}}={\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1}{\mathbf{Y}}^{\top}$.
 Subsequently, the estimated discriminant directions $\widehat{\alpha}_{k}$, for $1\leq k\leq K$, are computed by solving
-$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\\ \text{such that}&~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,\quad\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.16)
+$\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~{}\alpha^{\top}\widehat{C}_{b}~{}\alpha\quad\text{ s.t.}~{}\alpha^{\top}\widehat{C}_{w}~{}\alpha=1,~{}\alpha^{\top}\widehat{C}_{w}~{}\widehat{\alpha}_{i}=0,~{}~{}\forall~{}i<k,\end{split}$ (2.17)
 By similar reasoning for solving the Fisher’s discriminant analysis in (1.6), we propose to solve the optimization problem (2.17) via eigen-decomposing
@@
 The procedure in (2.17) computes Fisher’s discriminant directions $(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K})$ in
-$\text{span}({\mathbb{X}}\widehat{B})\subset\text{span}({\mathbb{X}})$
+$\text{span}({\mathbf{X}}\widehat{B})\subset\text{span}({\mathbf{X}})$
 of dimension less than $L$. Consequently, as long as
-${\rm rank}((\mathbb{I}_{n}-P_{{\mathbb{Y}}}){\mathbb{X}}\widehat{B})={\rm rank}({\mathbb{X}}\widehat{B}),$ (2.17)
+${\rm rank}((\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B})={\rm rank}({\mathbf{X}}\widehat{B}),$ (2.18)
 we do not suffer numerically in solving (2.17), in stark contrast with the algorithmic issue of solving (1.6) (or a penalized version thereof) in the original $p$-dimensional space.
@@
 (2015), the latter leads to meaningless solutions with the objective function equal to infinity.
-Finally, we remark that the requirement in (2.17) is easy to satisfy as ${\rm rank}({\mathbb{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbb{I}_{n}-P_{{\mathbb{Y}}})=n-L$ when we observe at least one data point per class.
+Finally, we remark that the requirement in (2.18) is easy to satisfy as ${\rm rank}({\mathbf{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbf{I}_{n}-P_{{\mathbf{Y}}})=n-L$ when we observe at least one data point per class.
 3 Theoretical guarantees
-In this section we provide theoretical guarantees for our procedure.
-We start by listing the main assumptions used in our analysis.
-The following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$ is assumed.
+In this section we provide a unified theory for analyzing the classifier $\widehat{g}$ in (2.13) based on estimating the regression coefficient matrix.
+In Section 3.1 we start with full generality and bound the excess risk of $\widehat{g}$ by a quantity that is related with the error of estimating $B^{*}$.
+The results there hold for any estimator $\widehat{B}^{*}$ of $B^{*}$.
+In Section 3.2 we focus on the proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ and provide a unified analysis of its estimation error that is valid for any generic estimator $\widehat{B}$.
+Finally we apply the general result to two particular choices of $\widehat{B}$, the lasso estimator in Section 3.3 and the reduced-rank estimator in Section 3.4.
+3.1 A reduction scheme for bounding the excess risk
+Let $(X,Y)$ be a new pair from (1.1) that is independent of the training data ${\bm{D}}:=\{{\mathbf{X}},{\mathbf{Y}}\}$.
+For any classifier $\widehat{g}:\mathbb{R}^{p}\to[L]$ in (2.13), its excess risk relative to the error rate of the Bayes rule in (1.2) is defined as
+$\mathcal{R}(\widehat{g}):=\mathbb{P}\{Y\neq\widehat{g}(X)\mid{\bm{D}}\}-\mathbb{P}\{Y\neq g^{*}(X)\}.$
+Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition.
+Our analysis of $\mathcal{R}(\widehat{g})$ requires the following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$.
 Assumption 1.
@@
 This is a common regularity assumption even in binary classification problems.
-We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 15 of Appendix C) at the cost of more technical proofs and a less transparent presentation.
-As we mentioned earlier, we decided to focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns.
+We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 21 of Appendix D) at the cost of more technical proofs and a less transparent presentation.
+The following proposition states that for the purpose of bounding $\mathcal{R}(\widehat{g})$, it suffices to study the estimation error of each discriminant function, that is, $\widehat{G}_{\ell}(X)-G_{\ell}(X)$ for $\ell\in[L]$.
+The proof essentially follows the argument of proving Theorem 12 in Bing and Wegkamp (2023) but with modifications to capture the sum of squared errors of estimating all discriminant functions.
+Proposition 2.
+Under model (1.1) with 1, for any sequences $t_{1},\ldots,t_{L}$ (that possibly depend on ${\bm{D}}$), with probability equal to one, one has
+$\mathcal{R}(\widehat{g})~{}\leq~{}C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$
+Proof.
+See Section B.2. ∎
+2 is valid for any estimator $\widehat{G}_{\ell}$ of the discriminant function, in particular, for our proposed $\widehat{G}_{\ell}$ in (2.12).
+The choice of $(t_{1},\ldots,t_{L})$ depends on the estimation error of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
+We therefore proceed to analyze
+$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)$ (3.2)
+for each $\ell\in[L]$.
+It is worth mentioning that in this section we allow $\widehat{B}^{*}$ in the above display to be any estimator of $B^{*}$ whereas $\widehat{\mu}_{\ell}$ and $\widehat{\pi}_{\ell}$ are given in (2.9).
+Since the variance of the first term in (3.2) depends on $\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$, we define
+$\Delta_{\infty}:=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
+This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under $\mathbb{E}(X)=0_{p}$)
+${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
+1, the Gaussian tail of $X\mid Y$ and the independence between $X$ and $({\mathbf{X}},{\mathbf{Y}})$ in (3.2) entail the following upper bound of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
+Theorem 3.
+Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$ and $L\log(n)\leq n$. With probability at least $1-3n^{-1}$, we have, for all $\ell\in[L]$,
+$\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|~{}\lesssim~{}\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}$ $+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}$
+$+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
+Proof.
+See Section B.3. ∎
+Condition $L\log(n)\leq n$ together with 1 ensures that $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
+The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish.
+It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
+Combining 2 with Theorem 3 immediately yields the following corollary.
+Denote by $\mathbb{E}_{{\bm{D}}}$ the expectation taken with respect to the training data ${\bm{D}}$ only.
+For any estimator $\widehat{B}^{*}$ of $B^{*}$, define the following (random) quantity
+$\mathcal{Q}(\widehat{B}^{*}):=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}(\Delta_{\infty}+\log n)+\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}.$ (3.3)
+Corollary 4.
+Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$. For any estimator $\widehat{B}^{*}$ of $B^{*}$, the corresponding classifier $\widehat{g}$ in (2.13) satisfies
+$\mathbb{E}_{\bm{D}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\mathbb{E}_{{\bm{D}}}\left[\min\left\{{\Delta_{\infty}}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right].$
+Proof.
+See Section B.4. ∎
+4 ensures that bounding $\mathcal{R}(\widehat{g})$ can be reduced to controlling $\mathcal{Q}(\widehat{B}^{*})$ for any estimator $\widehat{B}^{*}$ of $B^{*}$.
+In particular, 4 is valid for any existing approach based on estimating $B^{*}$.
+In the next section we provide a unified analysis of $\mathcal{Q}(\widehat{B}^{*})$ for our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ based on a generic $\widehat{B}$.
+As shown later, the first term in (3.3) is typically the dominating term.
+3.2 A unified analysis of the estimation of $B^{*}$ based on generic regression coefficient estimation
+In this section we provide more explicit bounds of $\mathcal{Q}(\widehat{B}^{*})$ in (3.3) for the estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ with $\widehat{H}$ obtained from (2.10).
+Our results in this section are valid for any estimator $\widehat{B}$ of $B$ that satisfies the following assumption.
+Assumption 2.
+There exists some deterministic sequences $\omega_{1}$ and $\omega_{2}$ such that the following holds with probability at least $1-n^{-1}$,
+$\max\left\{n^{-1/2}\|{\mathbf{X}}(\widehat{B}-B)\|_{F},~{}\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\right\}\leq\omega_{2},\quad\|\widehat{B}-B\|_{1,2}\leq\omega_{1}.$
+We need the following regularity conditions on $\Sigma_{w}$. It assumes the largest entry of $\Sigma_{w}$ is bounded from above and the smallest eigenvalue of $\Sigma_{w}$ is bounded away from zero.
+Assumption 3.
+There exists some positive constants $0<c\leq C<\infty$ such that $c<\lambda_{p}(\Sigma_{w})\leq\|\Sigma_{w}\|_{\infty}\leq C$.
+Theorem 5.
+Under model (1.1) with Assumptions 1 & 3, assume $\Delta\asymp 1$ and $L\log(n)\leq cn$ for some small constant $c>0$.
+For any $\widehat{B}$ satisfying 2 with $\omega_{2}\sqrt{L}\leq c,$ with probability at least $1-2n^{-1}$,
+the matrix $\widehat{H}$ in (2.10) is invertible,
+and
+$\mathcal{Q}(\widehat{B}^{*})\lesssim~{}L^{2}{\log^{2}(n)\over n}+\omega_{2}^{2}L^{2}\log(n)+\omega_{1}^{2}L^{3}{\log(n\vee p)\over n}.$ (3.4)
+Proof.
+See Section B.5. ∎
+Theorem 5 relates $\mathcal{Q}(\widehat{B}^{*})$ to the estimation error of $\widehat{B}$ in both the Frobenius norm and the $\ell_{1}/\ell_{2}$-norm.
+In the bound of (3.4), the second term on the right-hand-side is usually the leading term and its rate depends on specific choice of $\widehat{B}$.
+We provide concrete examples in Section 3.3 for the lasso estimator and in Section 3.4 for the reduced-rank estimator.
+Condition $\omega_{2}\sqrt{L}\leq c$ ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds so that $\widehat{H}$ is invertible.
+Furthermore, the bound in (3.4) suggests that the condition $\omega_{2}\sqrt{L}\leq c$ is also needed for consistent classification.
+3.3 Application to estimation of $B$ with entry-wise $\ell_{1}$-regularization
+In this section we focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns.
 Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
@@
 However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$.
-Moreover, the analysis of $\widehat{B}$ relies on the standard Restricted Eigenvalue condition (RE) on the matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$.
-Since ${\mathbb{X}}$ is a matrix with random entries, we resort to the following quantity on the within-class covariance matrix $\Sigma_{w}$.
-Definition 1.
-For any integer $1\leq s\leq p$, define
-$\kappa_{s}:=\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{v^{\top}\Sigma_{w}v\over v^{\top}v}$
-where $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
-Our analysis needs the following regularity conditions on $\Sigma_{w}$ and the matrix ${\mathbb{M}}$ of conditional means. For any matrix $M$, write $\|M\|_{\infty}=\max_{i,j}|M_{ij}|$.
+Our analysis also needs the following mild regularity condition on the matrix ${M}$ of conditional means.
 Assumption 4.
-For some $1\leq s\leq p$, there exists some absolute constants $0<c\leq C<\infty$ such that $c<\kappa_{s}\leq\|\Sigma_{w}\|_{\infty}\leq C$.
-Assumption 3.
-There exists an absolute constant $C<\infty$ such that $\|{\mathbb{M}}\|_{\infty}\leq C$.
-A condition akin to 2 is commonly used in sparse linear regression models, see, for instance, Bühlmann and Van de Geer (2011); Giraud (2021).
-3 is a mild regularity condition on the largest entry of ${\mathbb{M}}$.
-We remark that the requirements that $\kappa_{s}$, $\|\Sigma_{w}\|_{\infty}$ and $\|{\mathbb{M}}\|_{\infty}$ are absolutely bounded are not essential, and can be relaxed to allow dependence on $n$.
-3.1 Finite sample bounds of the proposed regression matrix estimation
-The following theorem states finite sample upper bounds of both the (in-sample and out-of-sample) prediction error and the estimation error of the regression matrix estimator $\widehat{B}$ in (1.14) with the lasso-penalty in (2.7).
-For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$.
-For any sequences $a_{n}$ and $b_{n}$, we write $a_{n}\lesssim b_{n}$ if there exists some constant $C>0$ such that $a_{n}\leq Cb_{n}$, and $a_{n}\asymp b_{n}$ if $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$.
+There exists an absolute constant $C<\infty$ such that $\|{M}\|_{\infty}\leq C$.
+The following theorem establishes explicit rates of $\omega_{2}$ and $\omega_{1}$ in 2 for the regression matrix estimator $\widehat{B}$ in (1.14) using the lasso-penalty in (2.7).
 Theorem 6.
-Under model (1.1) and (1.3) with Assumptions 13, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with
-$(s\vee L)\log(n\vee p)\leq c~{}n$
-Then, with probability at least $1-n^{-1}$, we have
-$\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbb{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
-$\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}$
-$\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}$
-for the estimator $\widehat{B}$ obtained from (1.14) with any $\lambda$ satisfying
-$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.2)
-and some finite, large enough constant $C>0$.
+Under model (1.1) with Assumptions 1, 3 and 4, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with $(s\vee L)\log(n\vee p)\leq c~{}n.$ Let the estimator $\widehat{B}$ be obtained from (1.14) with any $\lambda$ satisfying
+$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.5)
+and some finite, large enough constant $C>0$. Then, with probability at least $1-3n^{-1}$, 2 holds for
+$\omega_{2}\asymp\lambda\sqrt{sL},\qquad\omega_{1}\asymp\lambda sL.$
 Proof.
-See Section A.2. ∎
-The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$.
-For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbb{Y}}$ is not a linear model in ${\mathbb{X}}$.
-We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}$ in Lemma 6.
-For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbb{X}}$ are generated from a mixture of Gaussians.
+See Section B.6. ∎
+Our proof is in fact based on a weaker condition than 3 by replacing $\lambda_{p}(\Sigma_{w})$ with the Restricted Eigenvalue condition (RE) on $\Sigma_{w}$.
+The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$.
+For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbf{Y}}$ is not a linear model in ${\mathbf{X}}$.
+We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$ in Lemma 11.
+For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbf{X}}$ are generated from a mixture of Gaussians.
 In Lemma 12 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b).
 Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
-Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbb{X}}^{\top}({\mathbb{Y}}^{\prime}-{\mathbb{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbb{Y}}^{\prime}$ and different target matrix $B^{\prime}$.
+Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}^{\prime}-{\mathbf{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbf{Y}}^{\prime}$ and different target matrix $B^{\prime}$.
 Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
@@
 In practice, one can simply choose $\lambda$ via cross-validation, for instance, via the built-in function of the R-package glmnet.
-3.2 Estimation of the discriminant directional matrix
-We analyze the estimation of the discriminant matrix $B^{*}$ in this section. Recall that
-$\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$
-with $\widehat{H}$ obtained from (2.9).
-The following theorem states upper bounds of the transformed $(\widehat{B}^{*}-B^{*})$.
-Since $H^{-1}$ depends on ${\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$ (see Lemma 9 of Appendix B for its proof), our results depend in turn on the quantity
-$\Delta_{\infty}=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
-This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under (1.2))
-${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
-Theorem 3.
-Under the conditions in Theorem 2, assume that $\Delta_{\infty}\geq 1$ and assume there exists some sufficiently small constant $c>0$ such that
-$\Delta_{\infty}^{2}\sqrt{sL\log(n\vee p)\over n}\leq c.$ (3.3)
-Then, with probability at least $1-n^{-1}$, we have
-$\max_{\ell\in[L]}~{}\bigl{\|}\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\bigr{\|}_{2}\!\!~{}\lesssim~{}\Delta_{\infty}^{5/2}\sqrt{sL\log(n\vee p)\over n}.$
-Proof.
-See Section A.3. ∎
-Remark 3.
-Condition (3.3) ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds, and hence $\widehat{H}$ is invertible.
-In Theorem 4 below we will see that condition (3.3) is also needed for consistent estimation of the discriminant function.
-The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish.
-It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
-3.3 Finite sample bounds of the misclassification rates
-We first analyze the estimation of the discriminant functions in (2.1). For any $X$ that is distributed according to (1.1), we have
-$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right).$
-As a consequence of Theorem 3, the Gaussian distribution assumption of $X\mid Y$ and the independence between $X$ and $({\mathbb{X}},{\mathbb{Y}})$, we have the following control of the estimated discriminant function $\widehat{G}_{\ell}(X)$ uniformly over $\ell\in[L]$.
-Theorem 4.
-Under the conditions of Theorem 3, we have
-$\mathbb{P}\left\{\max_{\ell\in[L]}|\widehat{G}_{\ell}(X)-G_{\ell}(X)|~{}\gtrsim~{}\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}\right\}\leq 2n^{-\Delta_{\infty}}+n^{-1}.$
-Proof.
-See Section A.4. ∎
-Armed with Theorem 4, we are ready to analyze the proposed classifier $\widehat{g}$ in (2.12) in terms of its excess risk relative to error rate of the Bayes rule in (1.4).
-Specifically, for any pair $(X,Y)$ from (1.1) that is independent of $({\mathbb{X}},{\mathbb{Y}})$, the excess risk of $\widehat{g}$ is defined as
-$\mathcal{R}(\widehat{g}):=\mathbb{P}\left\{Y\neq{\widehat{g}(X)}\right\}-\mathbb{P}\left\{Y\neq{g^{*}(X)}\right\}.$
-Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition. In the proof of the following theorem, we show that, for any $t>0$,
-$\mathcal{R}(\widehat{g})$ $\leq(L-1)~{}t~{}\max_{k\in[L]}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}$
-$\quad+(L-1)~{}\mathbb{P}\left\{\max_{\ell\in[L]}\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|>t\right\}$
-This fact, in combination with Theorem 4, yields the following result on the rates of convergence of $\mathcal{R}(\widehat{g})$.
-Theorem 5.
-Under the same conditions of Theorem 3 and $\Delta_{\infty}\asymp 1$, we have
-$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}.$
-Proof.
-See Section A.5. ∎
+As an immediate corollary of 4, Theorem 5 and Theorem 6, we have the following upper bound of the misclassification rate of $\widehat{g}$ based on the entry-wise lasso estimator with $\lambda$ chosen as the rate in (3.5).
+Corollary 7.
+Under the conditions in Theorem 6, assume that $\Delta_{\infty}\asymp 1$. Then
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{sL^{2}\over n}\log^{2}(n\vee p),~{}1\right\}.$
 Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$.
 The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$.
-Theorem 5 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$.
+Guarantees on the excess risk are stronger since they together with the Bayes error, which is oftentimes easy to calculate, imply guarantees of the risk.
+7 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$.
 Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)).
-Indeed, provided
-$\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$
-using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
-$\mathcal{R}(\widehat{g})~{}\lesssim~{}{sL^{2}\log(n\vee p)\log(n)\over n}e^{-c\Delta_{\infty}}+n^{-a}$
-for some arbitrary large constant $a>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
+Indeed, provided $\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$ using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}{sL^{2}\over n}\log^{2}(n\vee p)e^{-c\Delta_{\infty}}+n^{-C_{0}}$
+for some arbitrary large constant $C_{0}>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
+3.4 Application to estimation of $B$ via reduced-rank regression
+We provide analysis for the classifier that estimates the regression coefficient matrix $B$ by reduced-rank regression.
+Specifically, let $\widehat{B}$ be obtained from (1.14) with $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ for some tuning parameter $\lambda>0$.
+The following theorem establishes the rates of $\omega_{2}$ and $\omega_{1}$ for which 2 holds.
+Write $r={\rm rank}(B)$.
+Theorem 8.
+Under model (1.1) with Assumptions 1 and 3, assume $(p+L)\log(n)\leq c~{}n$ for some sufficiently small constant $c>0$. For the estimator $\widehat{B}$ in (1.14) with any
+$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$ (3.6)
+and some finite, large enough constant $C>0$, with probability at least $1-2n^{-1}$, 2 holds for
+$\omega_{2}\asymp\sqrt{\lambda r},\qquad\omega_{1}\asymp\sqrt{\lambda rp}.$
+Proof.
+The proof is deferred to Section B.7. ∎
+For the reduced-rank estimator, establishing its in-sample prediction risk can be done without any condition on the design matrix (see, for instance, Bunea et al.
+(2011); Giraud (2011)).
+However, since 2 is also related with the out-of-sample prediction risk, we need the smallest eigenvalue of the Gram matrix $\widehat{\Sigma}$ to be bounded away from zero.
+The inequality $(p+L)\log(n)\leq cn$ is assumed for this purpose.
+In case $p$ is large, it is recommended that an additional ridge penalty be added in $\text{pen}(B)$ to alleviate the singularity issue of $\widehat{\Sigma}$ (Mukherjee and Zhu, 2011).
+Another main ingredient of our proof of Theorem 8 is to control $\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}$ in Lemma 13.
+It requires non-standard analysis for the same reason mentioned before that ${\mathbf{Y}}$ is not linearly related with ${\mathbf{X}}$.
+Combining 4, Theorem 5 and Theorem 8 immediately yields the following guarantee on the excess risk of the classifier using the reduced-rank estimator of $B$.
+Corollary 9.
+Under the conditions in Theorem 8, assume $\Delta_{\infty}\asymp 1$ and $\|\Sigma_{w}\|_{\rm op}\lesssim 1$. Then
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{(p+L)rL\over n}\log^{2}(n),~{}1\right\}.$
+Note that for the $p\times L$ matrix $B$ with rank $r$, its effective number of parameters is exactly $(p+L)r$ (Izenman, 1975) which could be much smaller than $pL$ when $r\ll(p\wedge L)$.
+9 suggests that the classifier using the reduced-rank estimator has benefit when $r$ is relatively small comparing to $p$ and $L$.
+The same remark as the one at the end of Section 3.3 applies, and our analysis can be extended to $\Delta_{\infty}\to\infty$, where the rate in 9 becomes exponentially fast in $\Delta_{\infty}$.
 4 Simulation study
 In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR).
-We also include the procedure in Section 2.3 (called LDRR-F) that is based on Fisher’s discriminant rule after estimating $B$.
-For estimating $B$ from (1.14), in addition to the lasso penalty (2.7), we also consider the elastic net penalty, $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{2}+(1-\alpha)\|B_{\ell}\|_{1}]$, for some $\alpha\in(0,1)$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet.
+We also include the procedure in Section 2.3 that is based on Fisher’s discriminant rule after estimating $B$ (called LDRR-F).
+For estimating $B$ from (1.14), in the sparse scenarios of Section 4.1, we consider both the lasso penalty (2.7) (L1) and the elastic net penalty (L1+L2), $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{1}+(1-\alpha)/2\|B_{\ell}\|_{2}^{2}]$, for some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet.
+In the low-rank scenarios of Section 4.2, we consider both the reduced-rank penalty (2.8) (RR) and the reduced-rank penalty plus the ridge penalty (RR+L2), $\text{pen}(B)=\lambda[\alpha~{}{\rm rank}(B)+(1-\alpha)/2\|B\|_{F}^{2}]$, with tuning parameters $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation.
 We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al.
@@
 These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters.
-Finally, the oracle procedure that uses the true values of ${\mathbb{M}}$, $\Sigma_{w}$, and $\pi$ in the computation of the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
+Finally, the oracle procedure that uses the true values of ${M}$, $\Sigma_{w}$, and $\pi$ in computing the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
+4.1 Sparse scenarios
 We generate the data from the following mean-shift scenario.
 A similar setting is also used in Witten and Tibshirani (2011).
-Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $\mu_{\ell,j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0.
+Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $[\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0.
 Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$.
@@
 For all settings, we fix $n=300$, $p=500$, $L=5$, $\rho=0.6$, $\sigma=1$ and $\alpha=0$ when they are not varied.
-We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report the averaged misclassification errors in Figure 1.
+We use $500$ test data points to compute the misclassification errors of each procedure, and for each setting, we consider $50$ repetitions and report their averaged misclassification errors in Figure 1.
 From Figure 1, it is clear that our procedures consistently outperform other methods across all settings.
-We observe that the performance of our algorithms gets improved as $n$ increases, $L$ decreases and $\sigma$ decreases.
+We observe that the performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases.
 Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms.
-These findings are in line with our theory in Section 3.
+These findings are in line with our theory in Section 3.3.
 It is a bit surprising to see that the classification error of our benchmark decreases as the correlation coefficient $\rho$ keeps increasing after $0.4$.
@@
 Finally, we find that both LDRR and LDRR-F as well as their lasso and elastic net variants have very comparable performance in all scenarios.
-This is as expected as the data is simulated according to model (1.1) and (1.3), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
+This is as expected as the data is simulated according to model (1.1), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
 Figure 1: The averaged misclassification errors of each algorithm for varying each parameter one at a time.
+4.2 Low-rank scenarios
+In this section we evaluate our proposed procedure when the rank of the conditional mean matrix $M$ is low and hence $B$ is also low-rank.
+We consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Appendix A for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank.
+Specifically, let the within-class covariance matrix $[\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$.
+To generate $M$, for some scalar $\eta>0$ and some randomly generated orthogonal matrix $A\in\mathbb{R}^{p\times r}$ with $A^{\top}A=\mathbf{I}_{r}$, we set $M=\eta A\alpha$ with entries of $\alpha\in\mathbb{R}^{r\times L}$ generated i.i.d.
+from $N(0,32/r)$.
+The quantity $\eta$ controls the magnitude of separation between the conditional means.
+We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$.
+For each setting, we fix $n=1000,L=10,r=3,p=100$ and $\eta=1$ when they are not varied.
+We set $n=500$ when we vary $\eta$.
+Figure 2 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
+From Figure 2 we see that the proposed approaches of using the reduced-rank penalty generally outperform the other methods.
+As $L$ gets larger with $r$ fixed, the benefit of using reduced-rank penalty becomes more visible, in line with 9.
+When the dimension ($p$) increases and gets closer to the sample size, the performance of LDRR(RR) deteriorates quickly while LDRR(RR+L2) still performs the best.
+For relative large $p$, incorporating the ridge penalty is beneficial as it improves the estimation of $B$ when the sample covariance matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ becomes ill-conditioned.
+Figure 2: The averaged misclassification errors of each algorithm.
 5 Real data analysis
-In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets that are commonly used in the literature for discriminant analysis.
+In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets and one low-dimensional football dataset.
 The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001).
@@
 We used the preprocessed Lymphoma data in Dettling (2004).
-For all three data sets, the features are centered to zero mean.
-Each dataset is randomly divided with 75% of samples as the training data and 25% as the test data.
-Such process was repeated 10 times and the averaged misclassification errors of each algorithm were reported in Table 1.
-Table 1: The averaged misclassification errors (in percentage).
-LDRR PenalizedLDA PAMR RDA
-Ramaswamy 14.4 37.6 28.8 15.4
-SRBCT 1.0 5.2 1.4 1.0
-Lymphoma 0.0 0.6 0.6 0.6
-Given the similar performance between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification.
-From Table 1, we see that our procedure LDRR-F has the smallest misclassification errors across all three datasets.
+The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes.
+We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction.
+We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$.
+For all four data sets, the features are centered to have zero mean.
+Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data.
+This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
+Table 1: The averaged misclassification errors (in percentage) and standard errors
+LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA
+Ramaswamy 30.7(0.051) 14.4(0.054) 22.6(0.048) 40.0(0.058) 30.4(0.058) 16.7(0.05)
+SRBCT 0.7(0.022) 0.3(0.011) 0.1(0.007) 4.4(0.052) 1.0(0.019) 0.0(0.0)
+Lymphoma 0.3(0.012) 0.5(0.017) 0.4(0.015) 1.9(0.034) 1.3(0.025) 1.3(0.025)
+Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014)
+Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification.
+We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison.
+From Table 1, we see that our procedure LDRR-F is close to have the best performance across all datasets.
+Choice of the best penalty depends on the data structure.
+For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent.
+For the Football dataset with $p=118$ and $L=10$, the selected rank is found to be $\widehat{r}=3$.
 Furthermore, in Figure 3, we show that LDRR-F with $K=2$ in (2.16) can be used for visualization in the 2D space of the first two discriminant directions.
-Figure 2: Visualization of Ramaswamy, SRBCT and Lymphoma data points in the space of the first two discriminant vectors.
+The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
+Figure 3: Visualization of Ramaswamy, SRBCT, Lymphoma and Football data points in the space of the first two discriminant vectors.
 Acknowledgements
 Wegkamp is supported in part by the National Science Foundation grant NSF DMS-2210557.
-Appendix A Proofs
+Appendix A Additional simulation study when the conditional mean matrix is low-rank and $\Sigma_{w}$ is approximately low-rank
+In this setting, we consider the setting in Remark 1 and also compare with the method (PCLDA) in Bing and Wegkamp (2023).
+As mentioned in Remark 1, PCLDA is a two-step procedure similar to LDRR-F but with the first step done in a unsupervised way by using PCA.
+We generate the data according to model (1.1) with $M=\eta A\alpha$ as in Section 4.2 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $[\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.1 with $\rho=0.2$ and its diagonal elements from Uniform(0,1).
+The quantity $\eta$ represents the signal-to-noise ratio (SNR) between the low-dimensional signal $\eta AZ$ and the noise $W$.
+As pointed out in Bing and Wegkamp (2023), classification becomes easier as $\eta$ increases.
+We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$.
+For all settings, we fix $n=1000,L=10,r=3,p=100$ and $\eta=2$ when they are not varied.
+We set $n=500$ when we vary $\eta$.
+Figure 4 depicts the misclassification errors of each algorithm averaged over 50 repetitions.
+From Figure 4 we have similar findings as in the low-rank setting in Section 4.2 that the proposed approaches of using the reduced rank penalty outperform other methods when $n$ is much larger than $p$.
+In the last setting when we vary the SNR $\eta$, it is interesting to notice that PCLDA, the procedure that reduces the dimension by PCA in its first step, has comparable performance to the LDRR-F(RR+L2) when $\eta$ is sufficiently large ($\eta\geq 2$ in these settings).
+This aligns with the minimax optimality of PCLDA derived in Bing and Wegkamp (2023) in the regime of moderate / large $\eta$.
+But when $\eta$ is small such as $\eta<2$, LDRR-F has benefit by reducing the dimension in a supervised way.
+Figure 4: The averaged misclassification errors of each algorithm.
+Appendix B Proofs
 Throughout the proofs, we will use the following notation. Define
-$\Delta={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}$
+$\Delta={M}^{\top}\Sigma_{w}^{-1}{M}$
 and write
@@
 The following event will be frequently used in our proofs. Define the event
-$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (A.1)
+$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (B.1)
 for some absolute constant $C>0$.
-Lemma 15 in Appendix C ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1.
+Lemma 21 in Appendix D ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1.
 Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
-$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (A.2)
+$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (B.2)
 as well as $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
-A.1 Proof of the equivalence between (1.4) and (2.13)
+B.1 Proof of the equivalence between (1.2) and (2.14)
 Proof.
-Recall that $B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}$. We find that
-$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}$
+Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
+$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}$
 so that, for any $\ell\in[L]$,
 $\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$
-$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
-$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{+}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
-Write the singular value decomposition of $\Sigma_{w}^{-1/2}{\mathbb{M}}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{\mathbb{M}})$.
+$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
+$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
+Write the singular value decomposition of $\Sigma_{w}^{-1/2}{M}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{M})$.
 It follows that the above equals to
-$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
-$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
+$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
+$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
 which, in view of (1.2) and (2.1), completes the proof. ∎
-A.2 Proof of Theorem 2
+B.2 Proof of 2
+For any $x\in\mathbb{R}^{p}$, write
+$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
+For any $t_{1},\ldots,t_{L}\geq 0$, define the event
+$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (B.3)
+In the proof of 2, we use the notation $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$.
 Proof.
-We work on the intersection of the event
-$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (A.3)
+We condition on ${\bm{D}}$ throughout the proof.
+The law of $\mathbb{E}$ and $\mathbb{P}$ is with respect to the randomness of $(X,Y)$, which is independent of ${\bm{D}}$.
+By definition, the excess risk $\mathcal{R}(\widehat{g})$ equals to
+$\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y=k\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y=k\right]\Bigr{\}}$
+$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y=k\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y=k\right]$
+$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y=k\right]$
+$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=\ell\right]\Bigr{\}}.$
+Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
+[…]
+[…]
+where, for all $k,\ell\in[L]$,
+$G^{(\ell|k)}(x):=\left(x-{\mu_{\ell}+\mu_{k}\over 2}\right)^{\top}\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{k})+\log{\pi_{\ell}\over\pi_{k}}={1\over 2}\left(G_{k}(x)-G_{\ell}(x)\right).$
+Fix any $t_{1},\ldots,t_{L}\geq 0$. Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
+$0>2G^{(\ell|k)}(X)\overset{\mathcal{E}_{t}}{\geq}\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)-(t_{k}+t_{\ell})\geq-(t_{k}+t_{\ell}).$
+By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that,
+$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$
+$+{1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{E}\left[\mathbbm{1}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y=k\right\}\right].$
+For the first term, we have
+$\sum_{k\neq\ell}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$ $\leq(L-1)\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=(L-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$ (B.4)
+Regarding the second term, it is bounded from above by
+${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (B.5)
+Note that $G^{(\ell|k)}(X)\mid Y=k$ is normally distributed
+$N\left(-{1\over 2}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})+\log\left(\pi_{\ell}/\pi_{k}\right),~{}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\right).$
+By the mean-value theorem, the quantity in (B.5) is no greater than
+${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})^{2}$ $\leq{1\over 2}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
+Together with (B.4) and $\pi_{k}\leq C/L$, the proof is complete. ∎
+B.3 Proof of Theorem 3
+Proof.
+Pick any $\ell\in[L]$. Recall that
+$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$ (B.6)
+Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (B.1) which implies (B.2), an upper bound for the last term on the right of (B.6).
+Recall that $\mathbb{P}\{\mathcal{E}_{\pi}\}\geq 1-2n^{-2}$ and note that we consider both the randomness of $(X,Y)$ and that of ${\bm{D}}$ in this proof.
+For the first term in (B.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and take a union bound over $\ell\in[L]$ to conclude that the event
+$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~{}\text{for all }\ell\in[L]\right\}$ (B.7)
+has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (B.6)
+$|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|,$
+the Cauchy-Schwarz inequality gives
+$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\ell\ell}}$ (B.8)
+Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\bm{D}}$ gives that, for all $t\geq 0$,
+$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\bm{D}}\right\}\leq 2e^{-t^{2}/2}.$
+Choosing $t=2\sqrt{\log(n)}$, taking the union bound over $k\in[L]$ and unconditioning yield that for all $\ell\in[L]$,
+$|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq\sum_{k=1}^{L}\pi_{k}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\log(n)}$
+$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (B.9)
+holds with probability at least $1-2Ln^{-2}$. The last step uses (B.8). Collecting (B.2), (B.7), (B.8) and (B.3) completes the proof. ∎
+B.4 Proof of 4
+Proof.
+For any $\ell\in[L]$, let
+${t_{\ell}\over C}=\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
+Recall from the proof of Theorem 3 that $1-\mathbb{P}(\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi})\leq 2n^{-2}+n^{-2}=3n^{-2}$ and
+$\sum_{\ell=1}^{L}\mathbb{P}\left\{\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\geq t_{\ell}\right\}\cap\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\right\}\leq{1\over n}.$
+By the fact $\mathcal{R}(\widehat{g})\leq 1$ almost surely and invoking 2, we find that
+$\mathbb{E}_{{\bm{D}}}[\mathcal{R}(\widehat{g})]$ $\leq\mathbb{E}_{{\bm{D}}}\left[\min\left\{C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]$
+$\quad+1-\mathbb{P}_{{\bm{D}}}\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}$
+$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
+$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right],$
+completing the proof. ∎
+B.5 Proof of Theorem 5
+The proof of Theorem 5 uses the following lemma on $\|\widehat{H}-H\|_{\rm op}$. The proof of Lemma 10 appears at the end of this section.
+Lemma 10.
+Under the conditions of Theorem 5, with probability at least $1-n^{-1}$,
+$\|\widehat{H}-H\|_{\rm op}\lesssim\sqrt{\log(n)\over nL}+{\omega_{2}\over\sqrt{L}}.$
+Proof of Theorem 5.
+Define
+$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (B.10)
+for some constant $C>0$. We work on the event
+$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\rm op}\leq\delta\right\}.$
+Lemma 10 ensures that $\mathcal{E}_{H}$ holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
+$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
+$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ by $\text{by }\Omega=H^{-1}$
+[…]
+[…]
+Hence, on the event $\mathcal{E}_{H}$, $\widehat{H}$ is invertible. We proceed to prove
+$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~{}\lesssim~{}\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (B.11)
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (B.12)
+which, in conjunction with (B.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
+To prove (B.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
+$\widehat{B}^{*}-B^{*}$ $=(\widehat{B}-B)\widehat{\Omega}+B(\widehat{\Omega}-\Omega)$
+$=(\widehat{B}-B)\widehat{\Omega}+B^{*}H(\widehat{\Omega}-\Omega)$ by Lemma 1
+$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ by $\text{by }H\Omega=\mathbf{I}_{L}$
+$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (B.13)
+It then follows that
+$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$
+$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}\|_{F}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}\|_{F}$
+$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\|\widehat{\Omega}\|_{\rm op}+\|\Sigma_{w}^{1/2}B^{*}\|_{F}~{}\|(H-\widehat{H})\widehat{\Omega}\|_{\rm op}$
+$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\|\widehat{\Omega}\|_{\rm op}+\sqrt{L\Delta_{\infty}}\|\widehat{H}-H\|_{{\rm op}}\|\widehat{\Omega}\|_{\rm op}.$
+In the last step, we use $\|\Sigma_{w}^{1/2}B^{*}\|_{F}^{2}={\rm tr}(B^{*\top}\Sigma_{w}B^{*})={\rm tr}({M}^{\top}\Sigma_{w}^{-1}{M})\leq L\Delta_{\infty}$.
+We bound $\|\widehat{\Omega}\|_{\rm op}$ as follows:
+$\|\widehat{\Omega}\|_{\rm op}$ $\leq\|\Omega\|_{\rm op}+\|\widehat{\Omega}-\Omega\|_{\rm op}$
+$=\|\Omega\|_{\rm op}+\|\Omega(H-\widehat{H})\widehat{\Omega}\|_{\rm op}$
+$\leq\|\Omega\|_{\rm op}+\|\Omega\|_{\rm op}\|H-\widehat{H}\|_{\rm op}\|\widehat{\Omega}\|_{\rm op}.$
+Since the event $\mathcal{E}_{H}$ and $\omega_{2}\sqrt{L}\leq c$ ensure that
+$\|\Omega\|_{\rm op}\|H-\widehat{H}\|_{\rm op}\leq\left(\max_{k\in[K]}{1\over\pi_{k}}+\Delta_{\rm op}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
+we conclude
+$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (B.14)
+It now follows from 2 that
+$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$ $\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty}$
+$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (B.15)
+proving (B.11).
+We now prove (B.12). Using the identity (B.5), we have, on the event $\mathcal{E}_{H}$,
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}$
+$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (B.16)
+For the first term on the right, we find
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}$ $=\sum_{\ell=1}^{L}(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}(\widehat{\mu}_{\ell}-\mu_{\ell})$
+$\leq\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\sum_{\ell=1}^{L}\|(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}\|_{1}.$
+On the one hand, on the event $\mathcal{E}_{\pi}$, Lemma 22 with $k=\ell$ and $v={\bm{e}}_{j}$, in conjunction with union bounds over $j\in[p]$ and $\ell\in[L]$, yields that
+$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (B.17)
+with probability at least $1-n^{-2}$. On the other hand, we notice that
+$\sum_{\ell=1}^{L}\|(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}\|_{1}$ $=\sum_{\ell=1}^{L}\sum_{i,j\in[p]}\left|{\bm{e}}_{i}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\widehat{\Omega}_{\ell}^{\top}(\widehat{B}-B)^{\top}{\bm{e}}_{j}\right|$
+$\leq\sum_{i,j\in[p]}\left(\sum_{\ell=1}^{L}({\bm{e}}_{i}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell})^{2}\right)^{1/2}\left(\sum_{\ell=1}^{L}({\bm{e}}_{j}^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell})^{2}\right)^{1/2}$
+$=\sum_{i,j\in[p]}\|\widehat{\Omega}(\widehat{B}-B)^{\top}{\bm{e}}_{i}\|_{2}\|\widehat{\Omega}(\widehat{B}-B)^{\top}{\bm{e}}_{j}\|_{2}$
+$\leq\|\widehat{\Omega}\|_{\rm op}^{2}\sum_{i,j\in[p]}\|(\widehat{B}-B)^{\top}{\bm{e}}_{i}\|_{2}\|(\widehat{B}-B)^{\top}{\bm{e}}_{j}\|_{2}$
+$\leq\|\widehat{\Omega}\|_{\rm op}^{2}\|\widehat{B}-B\|_{1,2}^{2}.$
+By invoking 2 and applying inequalities (B.14) and (B.17), we obtain
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (B.18)
+with probability at least $1-n^{-2}$. Regarding the second term on the right in (B.5), we have
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $\leq\sum_{\ell=1}^{L}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}^{2}$
+$\leq\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|H-\widehat{H}\|_{\rm op}^{2}\sum_{\ell=1}^{L}\|\widehat{\Omega}_{\ell}\|_{2}^{2}$
+$\leq\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|\widehat{\Omega}\|_{F}^{2}$
+$\lesssim\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}~{}L\|\widehat{\Omega}\|_{\rm op}^{2}$
+on the event $\mathcal{E}_{H}$. Further invoking (B.7) and (B.14) yields that
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~{}\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (B.19)
+with probability at least $1-n^{-2}$. Combining (B.18) with (B.19) proves (B.12), and hence completes the proof. ∎
+Proof of Lemma 10.
+Recall that
+$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (B.20)
+By triangle inequality, we have
+$\|\widehat{H}-H\|_{\rm op}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}+2\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\rm op}$
+$\quad+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}.$
+We proceed to bound each term on the right-hand-side separately. On the event $\mathcal{E}_{\pi}$, we first find that
+$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (B.21)
+Regarding the second term, 2 ensures that
+$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (B.22)
+For the last two terms, invoking Lemma 19 ensures with probability at least $1-n^{-2}$,
+$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (B.23)
+It also follows that
+$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\rm op}$ $\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{\rm op}\|\widehat{\Sigma}^{1/2}B\|_{\rm op}$
+$\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{\rm op}\left(\|B^{\top}\Sigma B\|_{\rm op}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\right)^{1/2}$
+$\lesssim\omega_{2}\sqrt{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
+where the last step uses
+[…]
+deduced from Lemma 14. Combining the above bounds with (B.21), (B.22) and (B.23) gives
+$\|\widehat{H}-H\|_{\rm op}$ $\lesssim\sqrt{\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}$
+$\qquad+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}+\omega_{2}\sqrt{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}+\omega_{2}^{2}$
+with probability at least $1-n^{-1}$. The result follows by using $\Delta_{\infty}\asymp 1$ and $\omega_{2}\sqrt{L}\leq c$ to collect terms. ∎
+B.6 Proof of Theorem 6: the lasso estimator of $B$
+The proof of Theorem 6 uses the following Restricted Eigenvalue condition (RE) on the within-class covariance matrix $\Sigma_{w}$.
+Definition 1.
+For any integer $1\leq s\leq p$, define
+$\kappa_{s}:=\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{v^{\top}\Sigma_{w}v\over v^{\top}v}$
+where $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
+We prove Theorem 6 under the following weaker condition than 3.
+Assumption 5.
+For some $1\leq s\leq p$, there exists some absolute constants $0<c\leq C<\infty$ such that $c<\kappa_{s}\leq\|\Sigma_{w}\|_{\infty}\leq C$.
+Proof of Theorem 6.
+We first prove that
+$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (B.24)
+hold with probability at least $1-2n^{-1}$. We work on the intersection of the event
+$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (B.25)
 with the event
-$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbb{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (A.4)
+$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (B.26)
 Recall $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
@@
 To this end, for any set $S\subseteq[p]$, we write $E_{S}:=\text{span}\{{\bm{e}}_{j}:j\in S\}$.
-Under 2, we define the space
-$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (A.5)
-for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$. Let $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$. Lemma 7 and the condition $(s\vee L)\log(n\vee p)\leq cn$ imply that
-${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (A.6)
-holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (A.6) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
-$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (A.7)
-holds for any $u\in\mathcal{C}(S,3)$. Eq. A.7 in conjunction with 2 ensures that $\mathcal{E}_{RE}$ holds.
-It is now relatively straightforward to finish the proof. Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
-${1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
+Under 5, we define the space
+$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (B.27)
+for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$.
+Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$.
+Lemma 12 and the condition $(s\vee L)\log(n\vee p)\leq c~{}n$ imply that
+${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (B.28)
+holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (B.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
+$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (B.29)
+holds for any $u\in\mathcal{C}(S,3)$. Eq. B.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
+It is now relatively straightforward to finish the proof of (B.24).
+Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
+${1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$ $\leq{\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}+\lambda\|B_{\ell}\|_{1}-\lambda\|\widehat{B}_{\ell}\|_{1}$
 from which one can deduce that
-$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (A.8)
-with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (A.7), we find on the event $\mathcal{E}_{RE}$
-${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbb{X}}\widehat{B}_{\ell}-{\mathbb{X}}B_{\ell}\|_{2}^{2}$
+$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (B.30)
+with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (B.29), we find on the event $\mathcal{E}_{RE}$
+${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$
 $\leq{3\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}$
-$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (A.8)
+$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (B.30)
 $\leq 6\lambda\sqrt{s}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}.$
+We have proved (B.24).
+Finally, the claim follows by noting that
+$\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}\leq\sqrt{L}\max_{\ell\in[L]}\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2},$
+$\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\leq\sqrt{L}\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\leq\sqrt{L}\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2},$
+$\|\widehat{B}-B\|_{1,2}=\sum_{j=1}^{p}\|\widehat{B}_{j\cdot}-B_{j\cdot}\|_{2}\leq\sum_{j=1}^{p}\|\widehat{B}_{j\cdot}-B_{j\cdot}\|_{1}=\sum_{\ell=1}^{L}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}.$
 The proof is complete. ∎
-A.2.1 Two key lemmas used in the proof of Theorem 2
+B.6.1 Two key lemmas used in the proof of Theorem 6
 Lemma 11.
-Under model (1.1) and (1.3) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
-$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}({\mathbb{Y}}_{\ell}-{\mathbb{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{\mathbb{M}}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
+Under model (1.1) with 1, assume $L\log(n\vee p)\leq n$. With probability at least $1-n^{-1}$, one has
+$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}({\mathbf{Y}}_{\ell}-{\mathbf{X}}B_{\ell})\right|\lesssim\left(\sqrt{\|\Sigma_{w}\|_{\infty}}+\|{M}\|_{\infty}\right)\sqrt{\log(n\vee p)\over nL}.$
 Proof.
 Note that
-${1\over n}{\mathbb{X}}^{\top}({\mathbb{Y}}-{\mathbb{X}}B)$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (A.9)
+${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (B.31)
 We first analyze the term $(\widehat{\Sigma}-\Sigma)B$.
@@
 We observe from Lemma 21 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$.
-Next, we invoke Lemma 13 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (B.5), (B.6) and (B.7) in Lemma 9 to derive that the inequalities
+Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (C.5), (C.6) and (C.7) in Lemma 14 to derive that the inequalities
 $\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over n}$
-$\quad+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}\|{\mathbb{M}}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
-$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
-$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{\mathbb{M}}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
-$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
+$\quad+\|{M}_{j\cdot}\|_{\infty}\sqrt{B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{\infty}\|{M}^{\top}B_{\ell}\|_{\infty}{\log(n\vee p)\over n}$
+$\quad+\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{2}\sqrt{B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over nL}$
+$\lesssim\sqrt{[\Sigma_{w}]_{jj}+\|{M}_{j\cdot}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}+\|{M}_{j\cdot}\|_{\infty}{\log(n\vee p)\over n}$
+$\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$
 hold with probability at least $1-(n\vee p)^{-3}$. After we take the union bound over $j\in[p]$ and $\ell\in[L]$, we have that
-$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{\mathbb{M}}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (A.10)
+$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (B.32)
 holds with probability $1-n^{-1}$.
-To bound the first term in (A.9), by recalling that $\Sigma_{XY}={\mathbb{M}}D_{\pi}$, we find
-${1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{\mathbb{M}}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
-$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{\mathbb{M}}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){\mathbb{M}}_{j\ell}$
-$={n_{\ell}\over n}(\widehat{\mathbb{M}}_{j\ell}-{\mathbb{M}}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){\mathbb{M}}_{j\ell}.$
+To bound the first term in (B.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
+${1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{M}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
+$={1\over n}\sum_{i=1}^{L}\left(X_{ij}-{M}_{j\ell}\right)\mathbbm{1}\{Y_{i}=\ell\}+\left({n_{\ell}\over n}-\pi_{\ell}\right){M}_{j\ell}$
+$={n_{\ell}\over n}(\widehat{M}_{j\ell}-{M}_{j\ell})+(\widehat{\pi}_{\ell}-\pi_{\ell}){M}_{j\ell}.$
 Invoking 1, Lemma 21 and Lemma 22 with $t=\sqrt{\log(p\vee n)}$ and taking a union bound over $j\in[p]$ and $\ell\in[L]$, we obtain
-$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbb{X}}_{j}^{\top}{\mathbb{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{\mathbb{M}}\|_{\infty}\sqrt{\log(n)\over nL}$ (A.11)
-with probability $1-n^{-1}$. Combining (A.10) and (A.11) completes the proof. ∎
+$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (B.33)
+with probability $1-n^{-1}$. Combining (B.32) and (B.33) completes the proof. ∎
 Recall the set
@@
 Lemma 12.
-Under model and (1.3) with Assumptions 1 & 2, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$.
+Under model (1.1) with Assumptions 1 & 5, assume $(s\vee L)\log(n\vee p)\leq n$ for some integer $1\leq s<p$.
 Then, on the event $\mathcal{E}_{\pi}$, with probability at least $1-(n\vee p)^{-2}$, the following holds uniformly over $v\in\Gamma_{s}$,
@@
 and
-$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (A.12)
+$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (B.34)
 Here
@@
 $|\mathcal{N}|\leq\sum_{|S|=\lfloor Cs\rfloor}|\mathcal{N}_{S}|\lesssim{p}^{Cs}.$
-We will bound the right hand side (A.12) via the same arguments used to prove Lemma 13, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$.
+We will bound the right hand side (B.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$.
 Specifically, repeating the proof of Lemma 18 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
-$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{\mathbb{M}}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
+$v^{\top}(\widehat{\Sigma}-\Sigma)v$ $\lesssim v^{\top}\Sigma_{w}v\left(\sqrt{t\over n}+{t\over n}\right)+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v+2{\|{M}^{\top}v\|_{2}\over\sqrt{L}}\sqrt{v^{\top}\Sigma_{w}v}\sqrt{t\over n}.$
 holds, with probability at least $1-6e^{-t}$. Note that
-${\|{\mathbb{M}}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v$
+${\|{M}^{\top}v\|_{2}^{2}\over L}\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v$
 under 1. By using the decomposition of $\Sigma$ in (2.6), choosing $t=Cs\log(n\vee p)$ for some large constant $C>0$, we obtain that
@@
 holds uniformly over $v\in\mathcal{N}$, with probability at least $1-(n\vee p)^{-C^{\prime\prime}s}$. On the event $\mathcal{E}_{\pi}$, invoking 1, we conclude that
-$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\lesssim v^{\top}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}v\sqrt{L\log(n)\over n}$
-holds uniformly over $\mathcal{N}$. The proof is complete in view of (A.12). ∎
-A.3 Proof of Theorem 3
+$\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{M}D_{\pi}{M}^{\top}v\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v\sqrt{L\log(n)\over n}$
+holds uniformly over $\mathcal{N}$. The proof is complete in view of (B.34). ∎
+B.7 Proof of Theorem 8: the reduced-rank estimator of $B$
 Proof.
-Define
-$\delta\asymp\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$ (A.13)
-We work on the event
-$\mathcal{E}_{H}:=\left\{\|\widehat{H}-H\|_{\infty,1}\leq\delta\right\}$ (A.14)
-which, according to Lemma 8, holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
-$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
-$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{1,\infty}$ $\text{by writing }\Omega:=H^{-1}$
-[…]
-[…]
-Hence, $\widehat{H}$ is invertible on the event $\mathcal{E}_{H}$. For simplicity, we write $\widehat{\Omega}=\widehat{H}^{-1}$. By definition,
-$\widehat{B}^{*}-B^{*}$ $=(\widehat{B}-B)\widehat{\Omega}+B(\widehat{\Omega}-\Omega)$
-$=(\widehat{B}-B)\widehat{\Omega}+B^{*}H(\widehat{\Omega}-\Omega)$ by Lemma 1
-$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})$ by $\text{by }H\Omega=\mathbb{I}_{L}$
-$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbb{I}_{L}.$
-Recall that (A.17) and pick any $\ell\in[L]$. We it then follows that
-$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
-$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\widehat{\Omega}_{\ell}\|_{2}+\|\Sigma_{w}^{1/2}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}$
-$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\max_{k\in[L]}\sqrt{B_{k}^{*\top}\Sigma_{w}B^{*}_{k}}~{}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
-$\leq\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+\sqrt{\Delta_{\infty}}\|\widehat{H}-H\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}$ by $\text{by }B^{*}=\Sigma_{w}^{-1}{\mathbb{M}}.$
-To bound from above $\|\widehat{\Omega}_{\ell}\|_{1}$, we have that
-$\|\widehat{\Omega}_{\ell}\|_{1}$ $\leq\|\Omega_{\ell}\|_{1}+\|(\widehat{\Omega}-\Omega){\bm{e}}_{\ell}\|_{1}$
-$=\|\Omega_{\ell}\|_{1}+\|\Omega(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{1}$
-$\leq\|\Omega_{\ell}\|_{1}+\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\|\widehat{\Omega}_{\ell}\|_{1}.$
-Since the event $\mathcal{E}_{H}$ and (3.3) ensure that
-$\|\Omega\|_{\infty,1}\|H-\widehat{H}\|_{\infty,1}\leq\left(\min_{k\in[K]}{1\over\pi_{k}}+\|\Delta\|_{1,\infty}\right)\delta\lesssim\delta L\Delta_{\infty}\leq{1\over 2},$
-we conclude, for any $\ell\in[L]$,
-$\|\widehat{\Omega}_{\ell}\|_{1}\leq 2\|\Omega_{\ell}\|_{1}\lesssim L\Delta_{\infty}.$ (A.15)
-It then follows that
-$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}$
-$\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{2,\infty}+\delta\sqrt{\Delta_{\infty}}\right)L\Delta_{\infty}$
-$\lesssim\Delta_{\infty}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}+\sqrt{\Delta_{\infty}^{5}}\sqrt{L\log(n)\over n}+\Delta_{\infty}^{2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n}$
-$\lesssim\Delta_{\infty}^{5/2}\|{\mathbb{M}}\|_{\infty}\sqrt{sL\log(n\vee p)\over n},$ (A.16)
-completing the proof. ∎
-The following lemma provide upper bounds for $\|\widehat{H}-H\|_{\infty,1}$ under condition of Theorem 3.
+Let $C>0$ be some constant to be specified later. We work on the event
+$\mathcal{E}_{\lambda}^{\prime}=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}\leq\sqrt{\lambda/C}\right\}$
+intersecting with
+$\mathcal{E}_{X}=\left\{\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma)\geq{c\over 2}\right\}.$
+Here $c$ is defined in 3.
+Lemma 20, Lemma 13 and 3 together with (3.6) ensure that $\mathcal{E}_{X}\cap\mathcal{E}_{\lambda}^{\prime}$ hold with probability at least $1-n^{-1}$.
+Write $\widehat{r}={\rm rank}(\widehat{B})$ and $r={\rm rank}(B)$.
+By the optimality of $\widehat{B}$, we have
+${1\over n}\|{\mathbf{Y}}-{\mathbf{X}}\widehat{B}\|_{F}^{2}+\lambda\widehat{r}$ $~{}\leq~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\lambda r.$
+Working out the squares gives
+${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq{2\over n}|\langle{\mathbf{Y}}-{\mathbf{X}}B,{\mathbf{X}}(\widehat{B}-B)\rangle|+\lambda(r-\widehat{r})$
+$\leq 2\|\widehat{B}-B\|_{*}{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}+\lambda(r-\widehat{r})$
+$\leq 2\sqrt{r+\widehat{r}}~{}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+\lambda(r-\widehat{r})$ by $\text{by $\mathcal{E}_{\lambda}^{\prime}$}.$
+Using the properties on $\mathcal{E}_{X}$, we find
+$\|\widehat{B}-B\|_{F}^{2}$ $\leq{2\over c}2\sqrt{\widehat{r}+r}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+{2\over c}\lambda(r-\widehat{r})$
+$\leq{8\lambda\over c^{2}C}(\widehat{r}+r)+{1\over 2}\|\widehat{B}-B\|_{F}^{2}+{2\over c}\lambda(r-\widehat{r})$ by $2xy\leq x^{2}/2+2y^{2}$
+$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ for $\text{for $C=4/c$}.$
+In the same way, on the event $\mathcal{E}_{X}$,
+${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq\sqrt{{8\over cC}\lambda(r+\widehat{r})}~{}{1\over\sqrt{n}}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}+\lambda(r-\widehat{r})$
+$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ for $\text{for $C=4/c$}.$
+This implies the rates for both $\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ and $\|\widehat{B}-B\|_{F}^{2}$.
+They imply in turn the rates for $\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}$ and $\|\widehat{B}-B\|_{2,1}$ invoking the inequalities
+$\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}^{2}$ $\leq\|\Sigma^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq{2\over c}\|\widehat{B}-B\|_{F}^{2}$ by (2.6)
+and
+$\|\widehat{B}-B\|_{1,2}^{2}$ $\leq{p}\|\widehat{B}-B\|_{F}^{2}.$
+This completes the proof. ∎
 Lemma 13.
-Under conditions of Theorem 3, with probability at least $1-n^{-1}$,
-$\|\widehat{H}-H\|_{\infty,1}\lesssim\Delta_{\infty}\sqrt{\log(n)\over nL}+\sqrt{\Delta_{\infty}}\|{\mathbb{M}}\|_{\infty}\sqrt{s\log(n\vee p)\over nL}.$
+Under model (1.1) and 1, assume $(p+L)\log(n\vee p)\leq n$. With probability at least $1-2n^{-2}$, we have
+${1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}\lesssim\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}.$
 Proof.
-Recall that
-$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (A.17)
-We first bound the $\|\cdot\|_{\infty,1}$ norm. By triangle inequality, we have
-$\|\widehat{H}-H\|_{\infty,1}$ $\leq\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|(\widehat{B}-B)^{\top}\widehat{\Sigma}B\|_{\infty,1}$
-$\quad+\|B^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\infty,1}+\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty,1}.$
-We proceed to bound each term separately. On the event $\mathcal{E}_{\pi}$, we find that
-$\|D_{\widehat{\pi}}-D_{\pi}\|_{\infty,1}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (A.18)
-Next, pick any $k,\ell\in[L]$. We bound from above
-$\textrm{I}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}(\widehat{B}_{k}-B_{k})\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2},$ (A.19)
-$\textrm{II}=(\widehat{B}_{\ell}-B_{\ell})^{\top}\widehat{\Sigma}B_{k}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}\|\widehat{\Sigma}^{1/2}B_{k}\|_{2},$
-$\textrm{III}=B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}.$
-Note that
-$\|\widehat{\Sigma}^{1/2}B_{k}\|_{2}\leq\left(B_{k}^{\top}\Sigma B_{k}+B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}\right)^{1/2}.$
-Meanwhile Lemma 14 ensures that, on the event $\mathcal{E}_{\pi}$ in (A.1),
-$\textrm{III}\leq\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\infty}\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}$ (A.20)
-with probability at least $1-n^{-3}$. The above two bounds together with (B.5) in Lemma 9 give that
-$\textrm{II}\lesssim{\sqrt{\Delta_{\infty}}\over L}\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}.$ (A.21)
-By combining (A.18), (A.19), (A.20) and (A.21) and taking the union bounds over $k,\ell\in[L]$, we conclude that
-$\|\widehat{H}-H\|_{\infty,1}$ $\lesssim~{}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{L\log(n)\over n}$
-$\qquad+\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}\sqrt{\Delta_{\infty}}+L\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{2,\infty}^{2}$
-with probability at least $1-n^{-1}$. Invoking Theorem 2 and using (3.3) to simplify the expression complete the proof. ∎
-A.4 Proof of Theorem 4
-Proof.
-Pick any $\ell\in[L]$. Start with
-$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$
-We work on $\mathcal{E}_{\pi}$ in (A.1) throughout the rest of the proof.
-For the first term in $\widehat{G}_{\ell}(X)-G_{\ell}(X)$, invoking Lemma 16 with $k=\ell$ and $v=B^{*}_{\ell}$, using the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and taking a union bounds over $\ell\in[L]$ gives that
-$\mathbb{P}\left\{\left\{\max_{\ell\in[L]}|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|\leq\sqrt{\Delta_{\infty}}\sqrt{aL\log(n)\over n}\right\}\cap\mathcal{E}_{\pi}\right\}\geq 1-2n^{-a}.$ (A.22)
-To bound from above
-$|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|,$
-using Cauchy-Schwarz inequality gives
-$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\mu_{\ell}^{\top}\Sigma_{w}^{-1/2}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}.$ (A.23)
-On the other hand, by using the same arguments of proving (A.3), we have that on the event $\mathcal{E}_{H}$
-$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B(\widehat{\Omega}_{\ell}-\Omega_{\ell})|$
-$\leq\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}\|\widehat{\Omega}_{\ell}\|_{1}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H\widehat{\Omega}-\mathbb{I}_{L})_{\ell}|$
-$\leq\left(\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}\|\widehat{B}-B\|_{1,\infty}+\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{\infty}\delta\right)2\|\Omega_{\ell}\|_{1}.$
-Here $\delta$ is defined in (A.13). Invoking (A.15), (A.22), (A.23) and Lemma 16 with $v={\bm{e}}_{j}$ and union bounds over $j\in[p]$ yields that, on $\mathcal{E}_{H}$ defined in (A.14),
-$|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $~{}\lesssim L\Delta_{\infty}\sqrt{\|\Sigma_{w}\|_{\infty}}\sqrt{L\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}$
-$\qquad+L\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L\log(n)\over n}$
-with probability at least $1-4(n\vee p)^{-2}$. Therefore, with probability at least $1-5(n\vee p)^{-2}$,
-$\begin{split}|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|&~{}\lesssim~{}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}}+\delta\sqrt{\Delta_{\infty}^{3}}\sqrt{L^{3}\log(n)\over n}\\ &\qquad+\Delta_{\infty}\sqrt{L^{3}\log(n\vee p)\over n}\|\widehat{B}-B\|_{1,\infty}.\end{split}$ (A.24)
-Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\mathbb{X}},{\mathbb{Y}}$ gives that, for all $t\geq 0$,
-$\mathbb{P}\left\{|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\geq|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+t\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}~{}\Big{|}~{}Y={\bm{e}}_{k},{\mathbb{X}},{\mathbb{Y}}\right\}\leq 2e^{-t^{2}/2}.$
-Choosing $t=\sqrt{4a\log(n)}$, taking the union bounds over $\ell\in[L]$ and unconditioning yield
-$\max_{\ell\in[L]}|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\geq\max_{k,\ell\in[L]}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+\max_{\ell\in[L]}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{4a\log(n)}$
-$\geq 3\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{2,\infty}\sqrt{(a\vee\Delta_{\infty})\log(n)}$ (A.25)
-with probability less than $2n^{-a}$. The last step uses (A.23). Collecting (A.2), (A.22), (A.24) and (A.4) and invoking Theorem 3 complete the proof. ∎
-A.5 Proof of Theorem 5
-For any $x\in\mathbb{R}^{p}$, write
-$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
-For any $t\geq 0$, define the event
-$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t~{}\Big{|}~{}{\mathbb{X}},{\mathbb{Y}}\right\}.$ (A.26)
-Proof.
-By definition, we start with
-$\quad\mathcal{R}(\widehat{g})$
-$=\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]\Bigr{\}}$
-$=\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{k\in[L]}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)\neq k\}\mid Y={\bm{e}}_{k}\right]$
-$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=k,g^{*}(X)=\ell\}\mid Y={\bm{e}}_{k}\right]$
-$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y={\bm{e}}_{k}\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid L=\ell\right]\Bigr{\}}.$
-Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
-[…]
-[…]
-where, for all $k,\ell\in[L]$,
-$G^{(\ell|k)}(x):=\left(x-{\mu_{\ell}+\mu_{k}\over 2}\right)^{\top}\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{k})+\log{\pi_{\ell}\over\pi_{k}}={1\over 2}\left(G_{k}(x)-G_{\ell}(x)\right).$
-Observe that the event $\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\cap\mathcal{E}_{t}$ implies
-$00>G_{z}^{(\ell|k)}(z)\overset{\mathcal{E}_{t}}{\geq}{1\over 2}\left(\widehat{G}_{k}(X)-\widehat{G}_{\ell}(X)\right)-t\geq-t,\quad\forall~{}t>$ 0 .
-By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that, for any $t>0$,
-$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\left\{t\pi_{k}\mathbb{E}\left[\mathbbm{1}\{-t\leq G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y={\bm{e}}_{k}\}\right]+\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
-$\leq t\sum_{k\in[L]}\pi_{k}\left\{\mathbb{P}\left(-t\leq\max_{\ell\in[L]\setminus\{k\}}G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right)+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y={\bm{e}}_{k})\right\}$
-$\leq t(m-1)\sum_{k\in[L]}\pi_{k}\max_{\ell\in[L]\setminus\{k\}}\mathbb{P}\left\{-t\leq G^{(\ell|k)}(X)\leq 0\mid Y={\bm{e}}_{k}\right\}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c})$
-Note that
-$G^{(\ell|k)}(X)\mid Y={\bm{e}}_{k}~{}\sim~{}N\left(-{1\over 2}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}+\log\left(\pi_{\ell}\over\pi_{k}\right),~{}\|\mu_{k}-\mu_{\ell}\|_{\Sigma_{w}}^{2}\right).$
-By the mean-value theorem, we conclude
-$\mathcal{R}(\widehat{g})$ $\leq(m-1)t^{2}+(m-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$
-Choosing
-$t=\Delta_{\infty}^{3}\sqrt{sL\log(n\vee p)\log(n)\over n}$
-and invoking Theorem 4 complete the proof. ∎
-Appendix B Technical lemmas
+By (B.31), it suffices to bound from above
+$\textrm{I}=\left\|(\widehat{\Sigma}-\Sigma)B\right\|_{\rm op},\qquad\textrm{II}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}\right\|_{\rm op}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right\|_{\rm op}.$
+For I, by invoking Lemma 20, we find with probability at least $1-n^{-2}$,
+I $\leq\|\Sigma^{1/2}\|_{\rm op}\|\Sigma^{-1/2}(\widehat{\Sigma}-\Sigma)\Sigma^{-1/2}\|_{\rm op}\|\Sigma^{1/2}B\|_{\rm op}$
+$\lesssim\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$
+where in the last step we also use
+$\|\Sigma\|_{\rm op}=\|\Sigma_{w}^{1/2}(\mathbf{I}_{p}+\Sigma_{w}^{-1/2}{M}D_{\pi}{M}^{\top}\Sigma_{w}^{-1/2})\Sigma_{w}^{1/2}\|_{\rm op}\leq\|\Sigma_{w}\|_{\rm op}\left(1+{\Delta_{\rm op}\over L}\right)\leq\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})$
+as well as
+$\|\Sigma^{1/2}B\|_{\rm op}^{2}=\|B^{\top}\Sigma B\|_{\rm op}=\|D_{\pi}{M}^{\top}B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L(L+\Delta_{\rm op})}\leq{1\over L}$
+from Lemma 14.
+Regarding II, by standard discretization argument, it suffices to bound from above
+$\max_{v\in\mathcal{N}_{p}(1/3),u\in\mathcal{N}_{L}(1/3)}v^{\top}\left({1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right)u$
+with $\mathcal{N}_{p}(1/3)$ and $\mathcal{N}_{L}(1/3)$ being the $(1/3)$-epsilon net of $\mathcal{S}_{p}$ and $\mathcal{S}_{m}$, respectively.
+Fix any $v\in\mathcal{N}_{p}(1/3)$ and $u\in\mathcal{N}_{L}(1/3)$.
+Observe that
+$v^{\top}\left({1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right)u$ $=v^{\top}(\widehat{M}D_{\widehat{\pi}}-{M}D_{\pi})u$
+$=v^{\top}(\widehat{M}-{M})D_{\widehat{\pi}}u+v^{\top}{M}\left(D_{\widehat{\pi}}-D_{\pi}\right)u$
+$=\sum_{k=1}^{L}u_{k}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})u_{k}v^{\top}\mu_{k}.$
+Invoking Lemma 15 and Lemma 16 yields that, for any $t\geq 0$, with probability $1-4e^{-t/2}$,
+II $\lesssim\|u\|_{2}\sqrt{tv^{\top}\Sigma_{w}v\over n}\sqrt{\max_{k}n_{k}\over n}+\sqrt{t\sum_{k=1}^{m}\pi_{k}u_{k}^{2}(v^{\top}\mu_{k})^{2}\over n}+{t\|u\|_{\infty}\|v^{\top}{M}\|_{\infty}\over n}$
+$\lesssim\sqrt{t\|\Sigma_{w}\|_{\rm op}\over nL}+\sqrt{t\|\Sigma_{w}\|_{\rm op}\Delta_{\infty}\over nL}+\sqrt{\|\Sigma_{w}\|_{\rm op}\Delta_{\infty}}{t\over n}$
+where in the last step we use
+$\|v^{\top}{M}\|_{\infty}=\|v^{\top}\Sigma_{w}^{1/2}\|_{2}\|\Sigma_{w}^{-1/2}{M}\|_{2,\infty}\leq\|\Sigma_{w}\|_{\rm op}^{1/2}\sqrt{\Delta_{\infty}}.$
+Recall that $|\mathcal{N}_{p}(1/3)|\leq 7^{p}$ and $|\mathcal{N}_{L}(1/3)|\leq 7^{L}$.
+Choosing $t=C(p+L)\log(n)$, taking the union bounds over $\mathcal{N}_{p}(1/3)$ and $\mathcal{N}_{L}(1/3)$ and combining with the bound of I complete the proof.
+∎
+Appendix C Technical lemmas
 The following lemma provides a few useful facts on quantities related with $B$ in (1.12). Write
-$\Omega:=D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}.$ (B.1)
+$\Omega:=D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}.$ (C.1)
 Lemma 14.
 With $B$ defined in (1.12), we have
-${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1},$ (B.2)
-$H=D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\Omega^{-1},$ (B.3)
-$B^{\top}\Sigma_{w}B=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}~{}\Omega^{-1}.$ (B.4)
+${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1},$ (C.2)
+$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (C.3)
+$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1}.$ (C.4)
 Consequently, for any $k,\ell\in[L]$, we have
-$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (B.5)
-$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (B.6)
-$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (B.7)
+$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (C.5)
+$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (C.6)
+$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (C.7)
 Proof.
 From Lemma 1 and its proof, we know that
-$B=B^{*}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)=\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)$
+$B=B^{*}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)=\Sigma_{w}^{-1}{M}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)$
 so that
-${\mathbb{M}}^{\top}B={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B\right)={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}-{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}B.$
+${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}-D_{\pi}{M}^{\top}B\right)={M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}-{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}{M}^{\top}B.$
 Rearranging terms gives
-${\mathbb{M}}^{\top}B$ $=\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}$
-$=\mathbb{I}_{L}-\left(\mathbb{I}_{L}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}D_{\pi}\right)^{-1}$
-$=\mathbb{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$ (B.8)
-$={\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1}$
-proving (B.2). By using the above identity (B.8), we have
-$D_{\pi}-D_{\pi}{\mathbb{M}}^{\top}B=\left(D_{\pi}^{-1}+{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\right)^{-1},$
-proving (B.3). Since Lemma 1 and (B.3) imply
+${M}^{\top}B$ $=\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}$
+$=\mathbf{I}_{L}-\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}$
+$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (C.8)
+$={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$
+proving (C.2). By using the above identity (C.8), we have
+$D_{\pi}-D_{\pi}{M}^{\top}B=\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1},$
+proving (C.3). Since Lemma 1 and (C.3) imply
 $B=B^{*}\Omega^{-1},$
-(B.4) follows from
-$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}}\Omega^{-1}.$
-To prove (B.5) – (B.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
-$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (B.9)
-yielding the second bound in (B.5). The other bound in (B.5) follows by observing that
+(C.4) follows from
+$B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}\Omega^{-1}.$
+To prove (C.5) – (C.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
+$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (C.9)
+yielding the second bound in (C.5). The other bound in (C.5) follows by observing that
 […]
-For (B.6), on the one hand, similar arguments yield
-$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{\mathbb{M}}D_{\pi}{\mathbb{M}}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
-On the other hand, using (B.2) and (B.3) proves
-$B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({\mathbb{M}}^{\top}\Sigma_{w}^{-1}{\mathbb{M}})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
+For (C.6), on the one hand, similar arguments yield
+$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{M}D_{\pi}{M}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
+On the other hand, using (C.2) and (C.3) proves
+$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({M}^{\top}\Sigma_{w}^{-1}{M})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
 Finally, the last statement follows by noting that
-$|\mu_{k}^{\top}B_{\ell}|^{2}\leq B_{\ell}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{\ell}\leq 1$
+$|\mu_{k}^{\top}B_{\ell}|^{2}\leq B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq 1$
 and
 $|\mu_{k}^{\top}B_{\ell}|=\pi_{\ell}|\mu_{k}^{\top}\Sigma^{-1}\mu_{\ell}|\leq\pi_{\ell}\sqrt{\mu_{k}^{\top}\Sigma^{-1}\mu_{k}}\sqrt{\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}}.$
-The following lemmas contain deviation inequalities related with $\widehat{\mathbb{M}}-{\mathbb{M}}$ and $\widehat{\pi}-\pi$.
+The following lemmas contain deviation inequalities related with $\widehat{M}-{M}$ and $\widehat{\pi}-\pi$.
 Lemma 15.
@@
 $\sum_{k=1}^{L}{n_{k}\over n}v^{\top}(\widehat{\mu}_{k}-\mu_{k})u_{k}$ $=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-\mu_{k}\right)u_{k}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
-$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
-$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{\mathbb{M}}Y_{i}\right)u^{\top}Y_{i}.$
-For $i\in[n]$, by conditioning on ${\mathbb{Y}}$, we know that $W_{i}:=X_{i}-{\mathbb{M}}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
-${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbb{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbb{Y}}^{\top}{\mathbb{Y}}u\over n}\right).$
+$=v^{\top}\sum_{k=1}^{L}{1\over n}\sum_{i=1}^{n}\left(X_{i}-{M}Y_{i}\right)u^{\top}Y_{i}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}$
+$={1\over n}\sum_{i=1}^{n}v^{\top}\left(X_{i}-{M}Y_{i}\right)u^{\top}Y_{i}.$
+For $i\in[n]$, by conditioning on ${\mathbf{Y}}$, we know that $W_{i}:=X_{i}-{M}Y_{i}$ are i.i.d. from $\mathcal{N}_{p}(0,\Sigma_{w})$, hence
+${1\over n}\sum_{i=1}^{n}v^{\top}W_{i}u^{\top}Y_{i}~{}\Big{|}~{}{\mathbf{Y}}\sim\mathcal{N}\left(0,{v^{\top}\Sigma_{w}v\over n}{u^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}u\over n}\right).$
 The claim follows by invoking the standard Gaussian tail probability bounds and after unconditioning. ∎
@@
 $\widehat{\Sigma}_{w}-\Sigma_{w}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}-\Sigma_{w}$
-$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{\mathbb{M}}Y_{i})(X_{i}-{\mathbb{M}}Y_{i})^{\top}-\Sigma_{w}\right].$
-Recall that, conditioning ${\mathbb{Y}}$, $W_{i}=X_{i}-{\mathbb{M}}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$.
+$={1\over n}\sum_{i=1}^{n}\left[(X_{i}-{M}Y_{i})(X_{i}-{M}Y_{i})^{\top}-\Sigma_{w}\right].$
+Recall that, conditioning ${\mathbf{Y}}$, $W_{i}=X_{i}-{M}Y_{i}$ for $i\in[n]$ are independent, $\mathcal{N}_{p}(0,\Sigma_{w})$.
 The result follows from standard concentration inequalities for the quadratic term of Gaussian random vectors.
 ∎
-B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
-Recall that $\widehat{\Sigma}=n^{-1}{\mathbb{X}}^{\top}{\mathbb{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (A.1).
+C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
+Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (B.1).
 Lemma 18.
@@
 $\left|v_{1}^{\top}(\widehat{\Sigma}-\Sigma)v_{2}\right|$ $\lesssim~{}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}\left(\sqrt{t\over n}+{t\over n}\right)$
-$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$
-$\quad+\left({\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
+$\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$
+$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
 Proof.
 By definition, we have
-$\widehat{\Sigma}$ $={1\over n}{\mathbb{X}}^{\top}{\mathbb{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
+$\widehat{\Sigma}$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{X}}={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}X_{i}X_{i}^{\top}$
 $={1\over n}\sum_{k=1}^{L}\left[\sum_{i=1}^{n}\mathbbm{1}\{Y_{i}={\bm{e}}_{k}\}(X_{i}-\mu_{k})(X_{i}-\mu_{k})^{\top}+n_{k}(\widehat{\mu}_{k}\mu_{k}^{\top}+\mu_{k}\widehat{\mu}_{k}^{\top})-n_{k}\mu_{k}\mu_{k}^{\top}\right]$
@@
 $\widehat{\Sigma}-\Sigma$ $=\widehat{\Sigma}_{w}-\Sigma_{w}+\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})\mu_{k}\mu_{k}^{\top}+\sum_{k=1}^{L}\widehat{\pi}_{k}\left[(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}+\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}\right].$
-The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 12.
-We bound the second term by invoking 1 and Lemma 11 with $u={\mathbb{M}}^{\top}v_{1}$ and $v={\mathbb{M}}^{\top}v_{2}$.
-This yields
-$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}\over n}$
-$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{\mathbb{M}}^{\top}v_{1}\|_{\infty}\|{\mathbb{M}}^{\top}v_{2}\|_{\infty}{t\over n}$ (B.11)
+The first term $\widehat{\Sigma}_{w}-\Sigma_{w}$ is bounded in Lemma 17. We bound the second term by invoking 1 and Lemma 16 with $u={M}^{\top}v_{1}$ and $v={M}^{\top}v_{2}$. This yields
+$\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}\over n}$
+$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (C.11)
 with probability at least $1-2e^{-t}$. Furthermore, on the event $\mathcal{E}_{\pi}$, 1 and $L\log(n)\leq n$ imply
 $\widehat{\pi}_{k}\lesssim\pi_{k}+\sqrt{\log(n)\over nL}\lesssim{1\over L},\quad\forall\ k\in[L],$
-and after we invoke Lemma 10 twice with $v=v_{1},u={\mathbb{M}}^{\top}v_{2}$ and $v=v_{2},u={\mathbb{M}}^{\top}v_{1}$, respectively, we get that
-$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (B.12)
-$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{\mathbb{M}}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{\mathbb{M}}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (B.13)
+and after we invoke Lemma 15 twice with $v=v_{1},u={M}^{\top}v_{2}$ and $v=v_{2},u={M}^{\top}v_{1}$, respectively, we get that
+$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (C.12)
+$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (C.13)
 hold with probability at least $1-4e^{-t}$. Collecting all three bounds completes the proof. ∎
@@
 For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$.
-Invoking Lemma 13 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (B.5), (B.6) and (B.7) of Lemma 9 gives that, with probability at least $1-n^{-2}$,
-$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
-$\quad+\|{\mathbb{M}}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
+Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (C.5), (C.6) and (C.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,
+$|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{M}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
+$\quad+\|{M}^{\top}B_{k}\|_{\infty}^{2}{\log(n)\over n}+2\sqrt{B_{k}^{\top}\Sigma_{w}B_{k}B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
 $\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)\sqrt{\log(n)\over nL^{2}}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}+\left(1\wedge{\Delta_{\infty}\over L}\right)^{2}{\log(n)\over n}$
 $\lesssim\left(1\wedge{\Delta_{\infty}\over L}\right)^{3/2}\sqrt{\log(n)\over nL}.$
-The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim. Regarding the operator norm bound, a standard discretization argument gives
+The last step uses $\Delta_{\infty}\gtrsim 1$ to collect terms. This proves the first claim.
+Regarding the operator norm bound, a standard discretization argument gives
 $\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\leq 3\max_{u\in\mathcal{N}_{L}(1/2)}u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$
@@
 $\|B^{\top}\Sigma_{w}B\|_{\rm op}\leq{\Delta_{\rm op}\over L(L+\Delta_{\rm op})},$
-$\|B^{\top}{\mathbb{M}}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
-$\|{\mathbb{M}}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
+$\|B^{\top}{M}\|_{\rm op}\leq{\Delta_{\rm op}\over L+\Delta_{\rm op}},$
+$\|{M}^{\top}Bu\|_{\infty}\leq\sqrt{\Delta_{\infty}\Delta_{\rm op}\over L(L+\Delta_{\rm op})}$
 deduced from Lemma 14 give that, with probability at least $1-8n^{-CL+L\log(7)}\geq 1-n^{-L}$,
-$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{\mathbb{M}}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
-$\quad+\|{\mathbb{M}}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{\mathbb{M}}{\mathbb{M}}^{\top}Bu}\sqrt{\log(n)\over n}$
+$u^{\top}B^{\top}(\widehat{\Sigma}-\Sigma)Bu$ $\lesssim u^{\top}B^{\top}\Sigma_{w}Bu\sqrt{L\log(n)\over n}+\|{M}^{\top}Bu\|_{\infty}\sqrt{u^{\top}B^{\top}{M}{M}^{\top}Bu}\sqrt{\log(n)\over n}$
+$\quad+\|{M}^{\top}Bu\|_{\infty}^{2}{L\log(n)\over n}+2\sqrt{u^{\top}B^{\top}\Sigma_{w}Buu^{\top}B^{\top}{M}{M}^{\top}Bu}\sqrt{\log(n)\over n}$
 $\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}$
 holds uniformly over $u\in\mathcal{N}_{L}(1/3)$. We also use $\Delta_{\infty}\gtrsim 1$ to simplify expressions in the last step above. This completes the proof. ∎
-Appendix C Auxiliary lemmas
+The following lemma provides upper bounds of the operator norm of $(\widehat{\Sigma}-\Sigma)$.
+Lemma 20.
+Under model (1.1) and 3, assume $(p+L)\log(n)\leq n$. Then with probability at least $1-n^{-2}$, the following holds uniformly over $v\in\mathcal{S}^{p}$
+$v^{\top}(\widehat{\Sigma}-\Sigma)v~{}\lesssim~{}v^{\top}\Sigma v\left(\sqrt{p\log(n)\over n}+\sqrt{L\log(n)\over n}\right).$
+As a result, with the same probability, we have
+$\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma).$
+Proof.
+The proof of the first statement follows from the same arguments of proving Lemma 12 and is thus omitted.
+For the second statement, for any $v\in\mathcal{S}^{p}$, we have
+$v^{\top}\widehat{\Sigma}v=v^{\top}\Sigma^{1/2}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right)\Sigma^{1/2}v\geq v^{\top}\Sigma v~{}\lambda_{p}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right).$
+Since, by Weyl’s inequality and on the event that the first result holds,
+$\lambda_{p}\left(\Sigma^{-1/2}\widehat{\Sigma}\Sigma^{-1/2}\right)\geq 1-\|\Sigma^{-1/2}(\widehat{\Sigma}-\Sigma)\Sigma^{-1/2}\|_{\rm op}\geq 1-\left(\sqrt{p\log(n)\over n}+\sqrt{L\log(n)\over n}\right)$
+we conclude
+$v^{\top}\widehat{\Sigma}v\geq{1\over 2}v^{\top}\Sigma v$
+uniformly over $\mathcal{S}^{p}$, completing the proof. ∎
+Appendix D Auxiliary lemmas
 The following lemmas are proved in Bing and Wegkamp (2023).

### v3

@@
 Linear Discriminant Analysis (LDA) is an important classification approach.
-Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses.
+Its simple linear form makes it easy to interpret, and it is well-suited for handling multi-class responses.
 It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression.
-In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix.
+In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between discriminant directions and regression coefficients.
 This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods.
-Moreover, our new formulation is amenable to analysis: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques.
+In contrast with the existing regression-based approaches, our new formulation is particularly amenable to establish provable guarantees: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques.
 As applications, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization as well as for using the reduced-rank regression, neither of which has yet been fully analyzed in the LDA context.
 Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
-Keywords: Multi-class classification, discriminant analysis, high-dimensional data, regularization, multivariate response regression, dimension reduction.
+Keywords: Dimension reduction, discriminant analysis, high-dimensional data, multi-class classification, multivariate response regression, regularization.
 1 Introduction
+Linear Discriminant Analysis (LDA) is a popular tool for predicting a categorical response using a set of explanatory variables.
+Its popularity stems from several favorable properties, such as ease of interpretation, reasonable robustness to departures from normality, and the capability to handle responses with multiple classes.
+See, Hastie et al.
+(1995) and many references in Seber (2009).
 This paper introduces a new regression-based approach to multi-class, high-dimensional linear discriminant analysis (LDA).
-The main premise that leads to LDA is that we have a sample of $n$ i.i.d.
-copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, with the label $Y\in\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ taking values in $L\geq 2$ classes and the conditional distributions of the feature vector $X\in\mathbb{R}^{p}$ satisfy
-$X\mid Y={\bm{e}}_{\ell}~{}\sim~{}\mathcal{N}_{p}\left(\mu_{\ell},\Sigma_{w}\right),\qquad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$ (1.1)
-Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite.
-For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$.
-Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem.
-Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$ throughout this paper mainly for simplicity, and furthermore that the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ are strictly positive for all $\ell\in[L]$.
-The Bayes rule under model (1.1) is to classify a new feature $x\in\mathbb{R}^{p}$ as
-$g^{*}(x):=\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.2)
+Assume that $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ are i.i.d.
+copies of a random pair $(X,Y)$, where the label $Y$ takes values in $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$, the canonical basis vectors in $\mathbb{R}^{L}$, with $L\geq 2$ classes, and the feature vector $X\in\mathbb{R}^{p}$ has conditional means collected in the $p$ by $L$ matrix ${M}=(\mu_{1},\ldots,\mu_{L})$ and the within-class covariance matrix
+$\Sigma_{w}={\rm Cov}(X\mid Y={\bm{e}}_{\ell}),\quad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$
+We assume $\Sigma_{w}$ to be strictly positive definite and the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$, for all $\ell\in[L]$, are strictly positive.
+We denote by $\Sigma$ the unconditional covariance matrix of $X$.
+Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$.
+For a new feature $x\in\mathbb{R}^{p}$, LDA predicts its corresponding label as
+$\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.1)
+The LDA rule in (1.1) coincides with the Bayes rule when $X\mid Y$ is Gaussian.
 See, for instance, Izenman (2008).
-Estimation of this classifier based on $(X_{i},Y_{i})$ with $i\in[n]$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
-Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022).
+Estimation of the classifier (1.1) based on $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
+1.1 Existing approaches
+Few approaches consider (ii) but notable exceptions are Levy and Abramovich (2023); Abramovich and Pensky (2019); Nibbering and Hastie (2022).
 We recall various approaches to deal with (i).
-We will use the following notation.
-Write
-$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\quad D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L},\quad{M}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.3)
-and denote the unconditional covariance matrix of $X$ by
-$\Sigma={\rm Cov}(X)=\Sigma_{w}+{M}D_{\pi}{M}^{\top}.$ (1.4)
-We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={M}D_{\pi}{M}^{\top}$ the between-class covariance matrix.
-The latter has rank less than $L$ due to $\mathbb{E}(X)=0_{p}$.
-In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$ to estimate this direction.
+In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\beta^{*}:=\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure to estimate this direction by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$.
 Other procedures based on the same or similar equations include, for instance, Qiao et al.
@@
 (2011); Fan et al.
-(2012) and Tibshirani et al.
-(2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022) for $L>2$.
-However, it isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed below shortly.
-Witten and Tibshirani (2011) alternatively use the fact that the Bayes rule (1.2) is equivalent with Fisher’s discriminant rule
-$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.5)
-Here the columns of the matrix $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are implicitly defined via, for $1\leq k\leq K$,
-$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~{}\beta^{\top}\Sigma_{b}\beta\quad\text{subject to}&~{}\beta^{\top}\Sigma_{w}\beta=1,~{}\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.6)
+(2012) for $L=2$, and Tibshirani et al.
+(2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022); Mai et al.
+(2019); Gaynanova et al.
+(2016) for $L>2$.
+Specifically, for multi-class responses, there are L1L-1 directions $\beta_{\ell}^{*}=\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{1})$ for $2\leq\ell\leq L$, which can be solved via
+$(\beta_{2}^{*},\ldots,\beta_{L}^{*})=\operatorname*{arg\,min}_{\beta_{2},\ldots,\beta_{L}}~\sum_{\ell=2}^{L}\left({1\over 2}\beta_{\ell}^{\top}\Sigma_{w}\beta_{\ell}-\beta_{\ell}^{\top}(\mu_{\ell}-\mu_{1})\right).$ (1.2)
+For large $p>n$, sparsity is imposed on $\beta_{2}^{*},\ldots,\beta_{L}^{*}$.
+These directions can be estimated by replacing $\Sigma_{w}$ and $\mu_{\ell}$ in (1.2) by their sample counterparts, and adding a penalty that encourages sparsity (Mai et al., 2019; Gaynanova et al., 2016; Wang et al., 2021; Zeng et al., 2024).
+However, it is not always reasonable to assume that the directions $(\beta_{2}^{*},\ldots,\beta_{L}^{*})$ are sparse, especially when $\Sigma_{w}$ has many non-negligible off-diagonal entries.
+Moreover, the above procedure–based on solving a quadratic program–is less appealing than the regression-based approaches discussed below in (1.5), (1.6) and (1.10), particularly in high-dimensional settings where structural or penalized estimation is desired and issues such as tuning parameter selection and computational efficiency become more critical.
+Also, see part (a) of our main contributions in Section 1.2.
+Witten and Tibshirani (2011) alternatively use the fact that the first term in the LDA rule (1.1) is equivalent with Fisher’s discriminant rule
+$\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell}),$ (1.3)
+where the columns of $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are defined via, for $k\in[K]$,
+$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~\beta^{\top}\Sigma\beta\quad\text{subject to}&~\beta^{\top}\Sigma_{w}\beta=1,~\beta^{\top}\Sigma_{w}F_{i}=0,~~\forall~i<k.\end{split}$ (1.4)
+Criteria similar to (1.3) – (1.4) and their connection to discriminant analysis have been considered in Safo and Ahn (2016); Ahn et al.
+(2021); Jung et al.
+(2019).
 Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$.
-However, the implicit formulation (1.6) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
+However, the formulation (1.4) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
 More closely related to our approach is optimal scoring which dates back to De Leeuw et al.
@@
 (2011); Gaynanova (2020) and the references therein.
-By writing the response matrix and feature matrix as
-${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.7)
-optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
+By writing the response matrix and feature matrix as ${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L}$ and ${\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p}$, optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
 $\begin{split}\min_{\ \theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}\quad\text{s.t.}~\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~\forall~i<k.\end{split}$ (1.5)
@@
 As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$.
-In the high-dimensional setting, regularization in (1.8) is needed.
+In the high-dimensional setting, regularization is needed.
 For instance, Hastie et al.
-(1995) consider
-$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.9)
-for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution to (1.9) can be computed.
+(1995) considers adding $\beta^{\top}\Omega\beta$ to the loss function of (1.5) for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution can be computed.
 A possible choice of $\Omega=\lambda\mathbf{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989).
@@
 $\begin{split}\min_{\Theta,\ B}&~\|{\mathbf{Y}}\Theta-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B)\quad\text{s.t.}~\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\Theta=n\mathbf{I}_{K},\quad\Theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}1_{n}=0_{K},\end{split}$ (1.6)
-and proves that the global solution to (1.10) can be computed, provided pen$(B)$ satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$.
-Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ and analyzes the resulting classifier.
+and proves that the global solution to (1.6) can be computed, provided pen(B)(B) satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$.
+Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (which is orthogonal invariant) and analyzes the resulting classifier when $X\mid Y$ is Gaussian.
 However, it is not clear how the analysis can be extended to other penalized estimation.
-Moreover, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ as well as its extensions such as the elastic net penalty.
-That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.10), let alone any theoretical property of the resulting estimator.
-Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1)
+In particular, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ and the elastic net penalty.
+Worse, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.6), let alone any theoretical property of the resulting estimator.
+1.2 Our contributions
+We propose a new regression based multi-class classification approach that is suitable in both low- and high-dimensional settings and has provable theoretical guarantees.
+Key to our approach is the following reformulation of the problem.
+We show in Lemma 1 of Section 2 that the discriminant directional matrix in (1.1)
 $B^{*}=\Sigma_{w}^{-1}{M}\in\mathbb{R}^{p\times L}$ (1.7)
@@
 $B:=\Sigma^{-1}\Sigma_{XY}:=[{\rm Cov}(X)]^{-1}{\rm Cov}(X,Y)\in\mathbb{R}^{p\times L}$ (1.8)
-not only share the same column space, but have a closed-form connection, to wit,
+not only share the same column space, but have a closed-form connection, specifically,
 $B^{*}$ $=BH^{-1}$ (1.9)
-for some invertible $L\times L$ matrix $H$ with explicit expression.
+for some invertible $L\times L$ matrix $H$, which we derive explicitly.
 This connection is novel and distinguishes our approach from the current literature.
-Although the fact that columns of $B$ and $B^{*}$ share the same span is known (see, for instance, Hastie et al.
+While existing literature has leveraged the fact that $B$ and $B^{*}$ share the same column space to develop two-step classification procedures, typically by first estimating this shared subspace (see, e.g., Hastie et al.
 (1995); Ye (2007); Lee and Kim (2015); Nie et al.
-(2022) and the references therein), their closed-form connection given by (1.13) has not been discovered, and a thorough theoretical investigation on the corresponding classifier has not been conducted.
+(2022); Gaynanova et al.
+(2016); Ahn et al.
+(2021)), these approaches lack theoretical guarantees for the resulting classifier.
+In contrast, our procedure is built upon the new closed-form characterization of $B$ given in (1.9), which not only leads to a straightforward classifier but also enables a rigorous theoretical analysis of its misclassification risk.
 To the best of our knowledge, only in the binary case ($L=2$), Mai et al.
-(2012) leverages sparse linear regression between an appropriately encoded label vector and the feature matrix ${\mathbf{X}}$ for classification and proves consistency.
-Extension of such procedure to handle multi-class responses with high-dimensional features has been a longstanding open problem.
+(2012) employs sparse linear regression between an appropriately encoded label vector and the feature matrix $\mathbf{X}$} for classification and proves consistency of their classifier.
+Extension of such regression formulation to handle multi-class responses with high-dimensional features has been a longstanding open problem.
 In view of (1.9), we propose to estimate the discriminant direction matrix $B^{*}$ by first estimating the regression matrix $B$ as
-$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
-and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are three-fold:
+$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B)$ (1.10)
+and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are two-fold:
   1. (a)
-    Akin to optimal scoring, our approach in (1.14) is regression-based, which, as pointed out in Hastie et al.
-(1995, 1994), is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the aforementioned procedures based on approximately solving the equation $\Sigma_{w}B^{*}=M$ for $B^{*}$.
+    From a methodological perspective, our approach in (1.10), akin to optimal scoring, is regression-based.
+As pointed out by Hastie et al.
+(1995, 1994), regression-based methods are generally much easier to compute – especially in high-dimensional settings – than canonical correlation analysis (Gaynanova, 2020), Fisher’s discriminant rule (Witten and Tibshirani, 2011), and the aforementioned procedures based on approximately solving in (1.2) as in (Mai et al., 2019; Cai and Zhang, 2019).
+Moreover, regression-based approaches are more amenable to extensions involving regularized and non-parametric regression, as well as model selection.
 We can also easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$ in (1.10).
+However, we caution that the nonlinear relation $\mathbb{E}[{\mathbf{Y}}\mid{\mathbf{X}}]$ and $\mathbf{X}$} complicates the theoretical analysis.
+    After computing $\widehat{B}$ from (1.10), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
+Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys both computational advantage and numerical stability, compared to the existing approaches.
+In Section 4.1, we further offer an alternative way to estimate the LDA rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$.
+This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization (see, for instance, the real data analysis in Section 5).
   2. (b)
-    After computing $\widehat{B}$ from (1.14), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context.
-Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys computational advantage as well as numerical stability, compared to the existing approaches.
-In Section 2.3, we further offer an alternative way to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$.
-This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization.
-  3. (c)
-    We are able to quantify the errors of estimating both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier, thanks to (1.13).
-Concretely, in Section 3.1 and Section 3.2 we provide a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for a generic estimator $\widehat{B}$ of $B$.
+    Theoretically, unlike the intractable analysis of optimal scoring with general regularization, we are able to quantify the errors in estimating both $B$ and $B^{*}$, and provide theoretical guarantees for the resulting classifier, thanks to (1.9).
+Concretely, under the assumption that the distributions of $X\mid Y={\bm{e}}_{\ell}$ are Gaussian $\mathcal{N}_{p}(\mu_{\ell},\Sigma_{w})$, we provide in Sections 3.1 and 3.2 a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for any generic estimator of $B^{*}$, including all existing estimators as well as our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{-1}$.
 While various choices of $\widehat{B}$ are feasible in our approach, more specifically in (1.10), we apply the general result to two popular estimators: the lasso estimator in Section 3.3 and the reduced rank estimator in Section 3.4.
@@
 Finally, we provide a thorough simulation study in Section 4 to corroborate our theoretical findings.
-In Section 5, we report our findings in several real data studies.
-The Appendix contains most of the proofs and additional simulation study.
-Notation.
+Section 5 contains our findings in several real data studies.
+The Supplement (Bing et al., 2025) contains all the proofs.
+1.3 Notation
 For any numbers $a,b\in\mathbb{R}$, we write $a\vee b=\max\{a,b\}$.
 For any vector $v$ and $1\leq q\leq\infty$, we use $\|v\|_{q}$ to denote the standard $\ell_{q}$-norm.
-For any positive integer $d$, we write $[d]:=\{1,\ldots,d\}$.
+For any positive integer $d$, we write $d]:=\{1,\ldots,d\}$.
 For any matrix $Q\in\mathbb{R}^{d_{1}\times d_{2}}$, any $i\in[d_{1}]$ and $j\in[d_{2}]$, we write $Q_{i\cdot}$ for its $i$th row and $Q_{j}$ for its $j$th column.
@@
 2 Methodology
-We discuss our classification procedure in this section. Our goal is to estimate the Bayes rule in (1.2) which, for a given $x\in\mathbb{R}^{p}$, requires estimation of the discriminant functions
-$G_{\ell}(x):=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}),\qquad\text{for all }\ell\in[L].$ (2.1)
-This, in turn, requires to estimate the discriminant directional matrix $B^{*}=\Sigma_{w}^{-1}{M}$.
-The following lemma establishes the connection between $B^{*}$ and the regression matrix $B$ in (1.12).
-Recall the definitions of $\pi$, $D_{\pi}$ and ${M}$ from (1.3) as well as $\Sigma$ from (1.4).
+We discuss our classification procedure in this section. Our goal is to estimate the rule in (1.1) which entails estimating the discriminant functions:
+$\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2x^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell})=\mu_{\ell}^{\top}B^{*}_{\ell}-2x^{\top}B^{*}_{\ell}-2\log(\pi_{\ell}),\quad\text{for all }\ell\in[L].$ (2.1)
+Note that there is no need to estimate the $p\times p$ matrix $\Sigma_{w}^{-1}$ explicitly.
+The following lemma establishes the connection between the discriminant directional matrix $B^{*}=(B_{1}^{*},\cdots,B_{L}^{*})$ in (1.7) and the regression matrix $B$ in (1.8).
+Write $\pi=(\pi_{1},\ldots,\pi_{L})^{\top}$ and $D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L}$.
 Lemma 1.
-Under model (1.1), we have
+We have
 $B^{*}=BH^{-1}$ (2.2)
@@
 $B=\Sigma^{-1}{M}D_{\pi}.$ (2.5)
-Next, using
-$\Sigma=\Sigma_{w}+\Sigma_{b}=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
+Next, using the fact that
+$\Sigma=\Sigma_{w}+{M}D_{\pi}{M}^{\top},$ (2.6)
 the Woodbury matrix identity yields
@@
 Lemma 1 suggests that estimation of $B^{*}$ could be done by first estimating the matrix $B$ and then estimating $H^{-1}$ according to (2.3).
-We discuss these two steps in details in the next two sections.
+We discuss these two steps in detail in the next two sections.
 2.1 Estimation of the regression matrix
-To accommodate high-dimensional data, we estimate $B$ via regularized regression methods in (1.14).
-Choices of the penalty term $\text{pen}(B)$ depend on the concrete problem at hand.
+To accommodate high-dimensional data, we estimate $B$ via penalized regression methods in (1.10).
+Choices of the penalty term depend on the concrete problem at hand.
 For instance, when entries of $B$ exhibit certain smoothness structure such as in some imaging or audio applications (Hastie et al., 1995), one could choose $\text{pen}(B)={\rm tr}(B^{\top}\Omega B)$ for some pre-specified “roughness” penalty matrix $\Omega$.
 In this paper we mainly focus on the following two types of structural estimation of $B$.
-Sparsity-based structural estimation.
+2.1.1 Sparsity-based structural estimation.
 If we believe that only a subset of the original features are important for predicting the label, then sparsity structures of the regression matrix $B$ are reasonable assumptions.
@@
 The corresponding matrix estimator $\widehat{B}$ is defined in (1.10) with the following lasso penalty (Tibshirani, 1996)
-$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}$ (2.7)
+$\text{pen}(B)$ $=\lambda\sum_{j=1}^{p}\|B_{j\cdot}\|_{1}=\lambda\sum_{\ell=1}^{L}\|B_{\ell}\|_{1}$ (2.7)
 for some tuning parameter $\lambda>0$.
@@
 In this case, one can retain a much smaller subset of important features and achieve feature selection.
+From Lemma 1 we see that row-sparsity of $B$ is equivalent to that of $B^{*}$.
+Our focus in this paper, however, is on prediction, not feature selection.
 Furthermore, one can also consider either a row-wise or a column-wise elastic net penalty (Zou and Hastie, 2005) if the features tend to (highly) correlate with each other.
 Our numerical results in Sections 4 and 5 include both the lasso penalty and the elastic net penalty.
-Low-rank structural estimation.
+2.1.2 Low-rank structural estimation.
 Different from sparsity, if the goal is to construct a low-dimensional subset of linear combinations of the original feature for classification purpose, then a low-rank penalty in (1.10) could be more appropriate.
@@
 It is entirely different from our approach in Section 2.2 but connected to the one in Section 4.1 in the way that our approach in Section 4.1 is also a two-step procedure but reduces the feature space in the first step in a supervised way via the regression step in (1.10).
-Numerical comparison of both procedures are also included in our simulation study in Appendix A.
-2.2 Estimation of the discriminant matrix and the Bayes rule
-The discriminant functions in (2.1) that form the basis of the Bayes classification rule (1.2) necessitate that we estimate the matrix of the conditional means ${M}$, the vector of prior probabilities $\pi$ as well as the discriminant directional matrix $B^{*}$.
+Numerical comparison of both procedures are also included in our simulation study in Section 4.3.2.
+2.2 Estimation of the discriminant functions
+The discriminant functions in (2.1) necessitates that we estimate the conditional mean matrix M{M} and the discriminant directional matrix $B^{*}$.
+The latter also requires estimating the vector of prior probabilities $\pi$.
 We adopt the following standard estimators
 $\widehat{M}={\mathbf{X}}^{\top}{\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1},\qquad D_{\widehat{\pi}}={\rm diag}(\widehat{\pi})={1\over n}{\mathbf{Y}}^{\top}{\mathbf{Y}}$ (2.9)
-to estimate ${M}$ and $D_{\pi}$, respectively.
+to estimate M{M} and $D_{\pi}$, respectively.
 Recall from Lemma 1 that $B^{*}=BH^{-1}$.
 Since the estimator $\widehat{B}$ of $B$ is computed in (1.10), in view of the closed-form of $H$ in (2.3), we propose to estimate $H$ by the plug-in estimator
-$\widehat{H}=D_{\widehat{\pi}}-{1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}={1\over n}\left({\mathbf{Y}}^{\top}{\mathbf{Y}}-\widehat{B}^{\top}{\mathbf{X}}^{\top}{\mathbf{X}}\widehat{B}\right).$ (2.10)
-Note that ${\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.14).
-Computation of ${\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of ${\mathbf{X}}$ consist of basis expansions.
+[…]
+Note that $^{\mathbf{X}}\widehat{B}$ is simply the in-sample fit of the regression in (1.10).
+Computation of $^{\mathbf{X}}\widehat{B}$ thus can be done efficiently even in non-parametric regression settings, for instance, when the columns of $\mathbf{X}$} consist of basis expansions.
 Now, using $B^{*}=BH^{-1}$, we can further estimate $B^{*}$ by
 $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}.$ (2.11)
-Here for any matrix $M$, $M^{+}$ denotes its moore-penrose inverse. Finally, in view of (2.1), we estimate the discriminant functions by
-$\widehat{G}_{\ell}(x)=\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$. }$ (2.12)
-The corresponding classifier is
-$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad\text{ for any $x\in\mathbb{R}^{p}$}.$ (2.13)
-Theoretically, we show in Theorem 5 of Section 3.2 that $\widehat{H}$ is invertible with overwhelming probability under the same mild condition that is required for consistent classification.
-In Theorem 6 of Section 3.3 and Theorem 8 of Section 3.4, the condition is verified for the lasso estimator, and the reduced rank estimator, of $B$, respectively.
+Here for any matrix $M$, $M^{+}$ denotes its Moore-Penrose inverse. Finally, we estimate the discriminant functions in (2.1) by
+$\widehat{\mu}_{\ell}^{\top}\widehat{B}^{*}_{\ell}-2x^{\top}\widehat{B}^{*}_{\ell}-2\log(\widehat{\pi}_{\ell}),\qquad\text{for all $\ell\in[L]$,}$ (2.12)
+based on which classification can be done subsequently.
+Theoretically, we show in Theorem 5 of Section 3.2 that $\widehat{H}$ is invertible with overwhelming probability under a mild condition on $\widehat{B}$ that is required for classification consistency.
+In Theorem 6 of Section 3.3 and Theorem 8 of Section 3.4, this mild condition is verified for both the lasso estimator and the reduced rank estimator of $B$, respectively.
 Moreover, it is worth mentioning that the inverse of $\widehat{H}$ always exists when any $\ell_{2}$-norm related penalty is deployed in (1.10).
-2.3 Classification based on reduced dimensions
-Since in practice it is often desirable to extract fewer discriminant directions than those in $B^{*}$, we discuss in this section an alternative procedure for this purpose, which estimates the Bayes rule by also using $\widehat{B}$ from (1.14) but in a different way.
-We start with the fact (its proof is stated in Section B.1 for completeness) that the Bayes rule in (1.2) is equivalent to
+3 Theoretical guarantees
+In this section we provide a unified theory for analyzing the classifier based on the estimated discriminant functions in (2.12) via estimating the regression coefficient matrix.
+In Section 3.1 we start with full generality and bound the excess risk of the classifier by a quantity that is related with the error of estimating $B^{*}$.
+The results there hold for any estimator $\widehat{B}^{*}$ of $B^{*}$.
+In Section 3.2 we focus on the proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ and provide a unified analysis of its estimation error that is valid for any generic estimator $\widehat{B}$.
+Finally we apply the general result to two particular choices of $\widehat{B}$, the lasso estimator in Section 3.3 and the reduced-rank estimator in Section 3.4.
+Throughout our analysis, both $p$ and $L$, as well as the parameters $M$ and $\Sigma_{w}$, are allowed to depend on the sample size $n$.
+For notational simplicity, we omit this dependence in our presentation.
+3.1 A reduction scheme for bounding the excess risk
+We adopt the following distributional assumption
+$X\mid Y={\bm{e}}_{\ell}~\sim~\mathcal{N}_{p}\left(\mu_{\ell},\Sigma_{w}\right),\qquad\text{for all }\ell\in[L].$ (3.1)
+See Remark 2 for extension to sub-Gaussian distributions of $X\mid Y$.
+Let $(X,Y)$ be a new pair from (3.1) that is independent of the training data ${\bm{D}}:=\{{\mathbf{X}},{\mathbf{Y}}\}$.
+The Bayes rule under (3.1) reads as
+$g^{*}(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x),\qquad\text{with }~G_{\ell}(x)=\mu_{\ell}^{\top}B^{*}_{\ell}\,-\,2x^{\top}B^{*}_{\ell}-2\log(\pi_{\ell}).$ (3.2)
+For any classifier
+[…]
+its excess risk relative to the error rate of the Bayes rule is defined as
+$\mathcal{R}(\widehat{g}):=\mathbb{P}\{Y\neq\widehat{g}(X)\mid{\bm{D}}\}-\mathbb{P}\{Y\neq g^{*}(X)\}.$
+[…]
+Assumption 1.
+There exists some absolute constants $0<c\leq C<\infty$ such that
+${c\over L}\leq\min_{k\in[L]}\pi_{k}\leq\max_{k\in[L]}\pi_{k}\leq{C\over L}.$ (3.4)
+This is a common regularity assumption in binary classification problems (see, for instance, Cai and Zhang (2019); Mai et al.
+(2019); Abramovich and Pensky (2019)).
+We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 21 of Section B.2) at the cost of more technical proofs and a less transparent presentation.
+An important quantity in this problem is the pairwise Mahalanobis distance in the feature space between any two label classes.
+Assumption 2.
+There exist some absolute constants $0<c\leq C<\infty$ and some $\Delta>0$ such that
+$c\Delta\leq\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq C\Delta.$ (3.5)
+[…]
+Proposition 2.
+Under model (3.1) with 1 and 2, for any positive sequences $t_{1},\ldots,t_{L}$ (that possibly depend on $\bm{D}$}), with probability equal to one, one has
+$\mathcal{R}(\widehat{g})~\leq~{C\over\sqrt{\Delta}}\sum_{\ell=1}^{L}t_{\ell}^{2}+\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$ (3.6)
+Proof.
+See Section A.1. ∎
+We remark that both 1 and 2 are needed to derive the above “fast-rate” bound in (3.6). Although our proof reveals that the following “slow-rate” bound
+$\mathcal{R}(\widehat{g})\leq\max_{1\leq\ell\leq L}t_{\ell}+\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}$ (3.7)
+holds without these assumptions, we will mainly focus on explicit excess risk bounds based on (3.6).
+2 is valid for any estimator $\widehat{G}_{\ell}$ of the discriminant function, in particular, for our proposed $\widehat{G}_{\ell}$ in (3.3).
+The choice of $(t_{1},\ldots,t_{L})$ depends on the estimation error of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
+We therefore proceed to analyze
+[…]
+for each $\ell\in[L]$.
+It is worth mentioning that in this section we allow $\widehat{B}^{*}$ in the above display to be any estimator of $B^{*}$ whereas $\widehat{\mu}_{\ell}$ and $\widehat{\pi}_{\ell}$ are given in (2.9).
+1, the Gaussian tail of $X\mid Y$ and the independence between $X$ and $({\mathbf{X}},{\mathbf{Y}})$ in (3.8) entail the following upper bound of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
+Theorem 3.
+Under model (3.1) with Assumptions 1 and 2, assume $\Delta\geq 1$ and $L\log(n)\leq n$. With probability at least $1-3n^{-1}$, we have, for all $\ell\in[L]$,
+$\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|~\lesssim~\sqrt{\Delta}\sqrt{L\log(n)\over n}$ $+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
+Proof.
+See Section A.2. ∎
+Condition $L\log(n)\leq n$ together with 1 ensures that $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
+The reasonable condition $\Delta\geq 1$ requires that the pairwise separation of conditional distributions between distinct classes does not vanish.
+It is assumed to simplify the presentation, but our analysis can be easily extended to the case $\Delta=\Delta(p)\to 0$ as $p\to\infty$.
+We refer to Bing and Wegkamp (2023), in particular Theorems 3 & 7 and Corollary 11 for $L=2$ and Theorem 12 for $L>2$.
+Combining 2 with Theorem 3 immediately yields the following corollary.
+Denote by $\mathbb{E}_{{\bm{D}}}$ the expectation taken with respect to the training data $\bm{D}$} only.
+For any estimator $\widehat{B}^{*}$ of $B^{*}$, define the following (random) quantity
+$\mathcal{Q}(\widehat{B}^{*}):=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}(\Delta+\log n)+\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}.$ (3.9)
+Corollary 4.
+Under model (3.1) with Assumptions 1 and 2, assume $\Delta\geq 1$. For any estimator $\widehat{B}^{*}$ of $B^{*}$, the corresponding classifier $\widehat{g}$ in (3.3) satisfies
+$\mathbb{E}_{\bm{D}}\left[\mathcal{R}(\widehat{g})\right]~\lesssim~\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~1\right\}\right].$
+Proof.
+See Section A.3. ∎
+[…]
+3.2 A unified analysis of the estimation of $B^{*}$ based on generic regression coefficient estimation
+In this section we provide more explicit bounds of $\mathcal{Q}(\widehat{B}^{*})$ in (3.9) for the estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ with $\widehat{H}$ obtained from (2.10).
+Our results in this section are valid for any estimator $\widehat{B}$ of $B$ that satisfies the following assumption.
+Condition 1.
+There exists some deterministic sequences $\omega_{1}$ and $\omega_{2}$ such that the following holds with probability at least $1-n^{-1}$,
+$\max\left\{n^{-1/2}\|{\mathbf{X}}(\widehat{B}-B)\|_{F},~\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\right\}\leq\omega_{2},\quad\|\widehat{B}-B\|_{1,2}\leq\omega_{1}.$
+We need the following regularity conditions on $\Sigma_{w}$.
+It assumes the entries in $\Sigma_{w}$ are bounded from above and the smallest eigenvalue of $\Sigma_{w}$ is bounded away from zero.
+The lower bound condition on $\lambda_{p}(\Sigma_{w})$ can be relaxed by a restricted eigenvalue condition on $\Sigma_{w}$, see the discussion after Theorem 6.
+Assumption 3.
+There exists some positive constants such that $c<\lambda_{p}(\Sigma_{w})\leq\|\Sigma_{w}\|_{\infty}\leq C<\infty$.
+Theorem 5.
+Under model (3.1) with Assumptions 1, 2 & 3, assume $\Delta\asymp 1$ and $L\log(n)\leq cn$ for some small constant $c>0$.
+For any $\widehat{B}$ satisfying 1 with $\omega_{2}\sqrt{L}\leq c,$ we have, with probability at least $1-2n^{-1}$, that
+the matrix $\widehat{H}$ in (2.10) is invertible,
+and
+$\mathcal{Q}(\widehat{B}^{*})\lesssim~L^{2}{\log^{2}(n)\over n}+\omega_{2}^{2}L^{2}\log(n)+\omega_{1}^{2}L^{3}{\log(n\vee p)\over n}.$ (3.10)
+Proof.
+See Section A.4. ∎
+Theorem 5 relates $\mathcal{Q}(\widehat{B}^{*})$ to the estimation error of $\widehat{B}$ only in both the Frobenius norm and the $\ell_{1}/\ell_{2}$-norm.
+In the bound of (3.10), the second term on the right-hand-side is usually the leading term and its rate depends on specific choice of $\widehat{B}$.
+This allows us to exploit certain properties, such as sparsity or low-rank, of the regression matrix $B$.
+We provide concrete examples in Section 3.3 for the lasso estimator and in Section 3.4 for the reduced-rank estimator.
+Condition $\omega_{2}\sqrt{L}\leq c$ ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds so that $\widehat{H}$ is invertible, as $H$ is invertible by Lemma 1.
+Furthermore, the bound in (3.10) suggests that the condition $\omega_{2}\sqrt{L}\leq c$ is also needed for consistent classification.
+3.3 Estimation of $B$ with entry-wise $\ell_{1}$-regularization
+In this section we focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns.
+Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
+$\Theta_{s}:=\left\{B\in\mathbb{R}^{p\times L}:\max_{\ell\in[L]}\|B_{\ell}\|_{0}\leq s\right\}.$
+For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$.
+However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$.
+Our analysis also needs the following mild regularity condition on the matrix M{M} of conditional means.
+Assumption 4.
+There exists an absolute constant $C<\infty$ such that $\|{M}\|_{\infty}\leq C$.
+The following theorem establishes explicit rates of $\omega_{2}$ and $\omega_{1}$ in 1 for the regression matrix estimator $\widehat{B}$ in (1.10) using the lasso-penalty in (2.7).
+Theorem 6.
+Under model (3.1) with Assumptions 1, 2, 3 and 4, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with $(s\vee L)\log(n\vee p)\leq c~n.$ Let the estimator $\widehat{B}$ be obtained from (1.10) with any $\lambda$ satisfying
+$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.11)
+and some finite, large enough constant $C>0$. Then, with probability at least $1-3n^{-1}$, 1 holds for
+$\omega_{2}\asymp\lambda\sqrt{sL},\qquad\omega_{1}\asymp\lambda sL.$
+Proof.
+See Section A.5. ∎
+Our proof is in fact based on a weaker condition than 3 by replacing $\lambda_{p}(\Sigma_{w})$ with the Restricted Eigenvalue condition (RE) on $\Sigma_{w}$ (see Definition 1 in Section A.5 of the supplement).
+The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$.
+For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact $\mathbf{Y}$} is not a linear model in $\mathbf{X}$}.
+We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$ in Lemma 11.
+For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of $\mathbf{X}$} are generated from a mixture of Gaussians.
+In Lemma 12 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b).
+Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
+Gaynanova (2020) considers a group-lasso regularized regression similar to (1.10) and derives the bound of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}^{\prime}-{\mathbf{X}}B^{\prime})\|_{2}$ for some different response matrix $\mathbf{Y}$$^{\prime}$ and different target matrix $B^{\prime}$.
+Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
+In the rest of the paper, we assume that $\lambda$ is chosen as the order specified in (3.11).
+Although both $\|\Sigma_{w}\|_{\infty}$ and $\|M\|_{\infty}$ in (3.11) could, in principal, be consistently estimated, we recommend, on the basis of our regression formulation, to select $\lambda$ in practice simply via cross-validation, for instance, via the built-in function cv.glmnet of the R-package glmnet.
+As an immediate corollary of 4, Theorem 5 and Theorem 6, we have the following upper bound of the misclassification rate of $\widehat{g}$ based on the entry-wise lasso estimator with $\lambda$ chosen as the rate in (3.11).
+Corollary 7.
+Under the conditions in Theorem 6, assume that $\Delta\asymp 1$. Then
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~\lesssim~\min\left\{{sL^{2}\over n}\log^{2}(n\vee p),~1\right\}.$
+Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$.
+The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk under model (3.1) only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$.
+Guarantees on the excess risk are stronger since they together with the Bayes error, which is oftentimes easy to calculate, imply guarantees of the risk.
+7 focuses on the statistically most interesting case $\Delta\asymp 1$ where the pairwise separation between different classes are constant order.
+Otherwise, if $\Delta\to\infty$ as $p\to\infty$, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)).
+Indeed, using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~\lesssim~{sL^{2}\over n}\log^{2}(n\vee p)e^{-c\Delta}+n^{-C_{0}}$
+for some arbitrary large constant $C_{0}>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta$.
+We conclude this section with a comparison between our penalized linear discriminant method and the penalized multinomial logistic regression approach in Levy and Abramovich (2023); Abramovich et al.
+(2021); Lei et al.
+(2019).
+The main difference between both approaches is that logistic regression models the conditional distribution of the label $Y$ given the feature $X=x$ via
+$\log{\mathbb{P}\{Y=1\mid X=x\}\over\mathbb{P}\{Y=k\mid X=x\}}=\langle\beta_{k}^{*},x\rangle+\beta_{0,k}^{*}$
+while discriminant analysis postulates the distribution of the feature $X$ given the label $Y=k$.
+This distinction complicates any comparison between both methods.
+In particular, comparison of minimax lower bounds between the two methods is irrelevant.
+The sparse linear classifier $\widehat{g}$ in Abramovich et al.
+(2021) satisfies
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~\lesssim~\sqrt{sL+s\log(ep/s)\over n},$ (3.12)
+when the true coefficients $\beta_{k}^{*}$ are $s$-sparse ($s\leq p$). Furthermore, Abramovich et al. (2021) prove that this rate is minimax optimal in their sparse setting.
+We emphasize that Levy and Abramovich (2023); Abramovich et al.
+(2021); Lei et al.
+(2019) assume that the directions $\beta_{k}^{*}$ are sparse.
+In contrast, we assume structure on the regression matrix $B$ (not the matrix $B^{*}$ of directions).
+In particular, if we assume that our regression matrix has $s$ entire rows equal to zero, then $B^{*}$ has $s$-sparse columns.
+In this setting, we observe that the “slow-rate” bound in (3.7) implies a risk bound of order $\sqrt{sL/n}$ for our procedure (up to logarithmic factors) and that this rate matches the rate in (3.12).
+Hence, the excess risk of our classifier is no worse than the bound in (3.12) and is in fact faster, using the rate in 7, when either $sL^{3}\ll n$ or $\Delta\gg\log(L)$.
+[A reasonable assumption, given that we primarily focus on small to moderate values of $L$ in this paper.
+Determination of the optimal dependence on $L$ in the excess risk under our model remains an interesting open question, which we leave for future research.] Finally, although Abramovich et al.
+(2021) show that the rate in (3.12) can be improved under an additional margin condition that extends the one introduced by Tsybakov (2004) in binary classification, a direct comparison with our risk bound is challenging, as verifying this margin condition under our model poses a difficult problem in its own right.
+3.4 Estimation of $B$ via reduced-rank regression
+We analyze the classifier that estimates the regression coefficient matrix $B$ by reduced-rank regression.
+Specifically, let $\widehat{B}$ be obtained from (1.10) with $\text{pen}(B)=\lambda~{\rm rank}(B)$ for some tuning parameter $\lambda>0$.
+The following theorem establishes the rates of $\omega_{2}$ and $\omega_{1}$ for which 1 holds.
+Write $r={\rm rank}(B)$.
+Theorem 8.
+Under model (3.1) with Assumptions 1,2 and 3, assume $(p+L)\log(n)\leq c~n$ for some sufficiently small constant $c>0$. For the estimator $\widehat{B}$ in (1.10) with any
+$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta)}\sqrt{(p+L)\log(n)\over nL}$ (3.13)
+and some large constant $C>0$, with probability at least $1-2n^{-1}$, 1 holds for
+$\omega_{2}\asymp\sqrt{\lambda r},\qquad\omega_{1}\asymp\sqrt{\lambda rp}.$
+Proof.
+The proof is deferred to Section A.6. ∎
+For the reduced-rank estimator, establishing its in-sample prediction risk can be done without any condition on the design matrix (see, for instance, Bunea et al.
+(2011); Giraud (2011)).
+However, since 1 is also related with the out-of-sample prediction risk, we need the smallest eigenvalue of the Gram matrix $\widehat{\Sigma}$ to be bounded away from zero.
+The inequality $(p+L)\log(n)\leq cn$ is assumed for this purpose.
+In case $p$ is large, it is recommended that an additional ridge penalty be added in $\text{pen}(B)$ to alleviate the singularity issue of $\widehat{\Sigma}$ (Mukherjee and Zhu, 2011).
+Another main ingredient of our proof of Theorem 8 is to control $\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}$ in Lemma 13.
+It requires non-standard analysis for the same reason mentioned before that $\mathbf{Y}$} is not linearly related with $\mathbf{X}$}.
+Combining 4, Theorem 5 and Theorem 8 immediately yields the following guarantee on the excess risk of the classifier using the reduced-rank estimator of $B$.
+Corollary 9.
+Under model (3.1) with Assumptions 1 and 2, assume $(p+L)\log(n)\leq c~n$ for some sufficiently small constant $c>0$.
+Further assume $\Delta\asymp 1$ and $\lambda_{p}(\Sigma_{w})\asymp\lambda_{1}(\Sigma_{w})\asymp 1$.
+Then
+$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~\lesssim~\min\left\{{(p+L)rL\over n}\log^{2}(n),~1\right\}.$
+Note that for the $p\times L$ matrix $B$ with rank $r$, its effective number of parameters is exactly $(p+L)r$ (Izenman, 1975) which is much smaller than $pL$ when $r\ll(p\wedge L)$.
+9 suggests that the classifier using the reduced-rank estimator has benefit when $r$ is relatively small comparing to $p$ and $L$.
+The same remark as the one at the end of Section 3.3 applies, and our analysis can be extended to $\Delta\to\infty$, where the rate in 9 becomes exponentially fast in $\Delta$.
+Comparing to 3, a stronger condition of bounded eigenvalues of $\Sigma_{w}$ is assumed here to simplify the presentation.
+The latter holds, for instance, when the features in $X$ exhibit weak dependence or are uncorrelated within each response class.
+Remark 2 (Extension to sub-Gaussian conditional distributions).
+[…]
+$\mathcal{R}(\widehat{g})~\leq~\max_{1\leq\ell\leq L}t_{\ell}+\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$
+[…]
+$\max_{\ell\in[L]}~\mathbb{P}\left(0<\min_{k\in[L]\setminus\{\ell\}}G_{k}(X)-G_{\ell}(X)<2t\mid Y={\bm{e}}_{\ell}\right)\leq Ct^{\alpha},\quad\text{for all }t\geq 0,$
+and for some $C\geq 0$ and $\alpha\geq 1$.
+4 Simulation study
+In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR).
+We also include the procedure in Section 4.1 that is based on Fisher’s discriminant rule after estimating $B$ (called LDRR-F).
+For estimating $B$ from (1.10), in the sparse scenarios of Section 4.2, we consider both the lasso penalty (2.7) (L1) and the elastic net penalty (L1+L2), $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{1}+(1-\alpha)/2\|B_{\ell}\|_{2}^{2}]$, for some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet.
+In the low-rank scenarios of Section 4.3, we consider both the reduced-rank penalty (2.8) (RR) and the reduced-rank penalty plus the ridge penalty (RR+L2), $\text{pen}(B)=\lambda[\alpha~{\rm rank}(B)+(1-\alpha)/2\|B\|_{F}^{2}]$, with some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation.
+We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al.
+(2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al.
+(2007), the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011), the multiclass sparse discriminant analysis (MSDA) method of Mai et al.
+(2019) and the sparse multiclass discrimination with trace regularization (SLDTR) method of Ahn et al.
+(2021).
+These methods are available in R-packages pamr, rda, penalizedLDA, msda and Matlab codes for SLDTR and we use their default methods for selecting tuning parameters.
+Finally, the oracle procedure that uses the true values of M{M}, $\Sigma_{w}$, and $\pi$ in computing the discriminant functions $G_{k}(x)$ in (3.2) serves as our benchmark.
+4.1 A practical classification based on reduced dimensions
+Since in practice it is often desirable to extract fewer discriminant directions than those in $B^{*}$, we discuss in this section an alternative procedure for this purpose, which estimates the LDA rule in (1.1) by also using $\widehat{B}$ from (1.10) but in a different way.
+We start with the fact (see Section A.7 for a proof) that the rule in (1.1) is equivalent to
 $\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$ (4.1)
@@
     Perform linear discriminant analysis on the space of $z$.
-The second point follows by noticing that
-$C_{w}:=B^{\top}\Sigma_{w}B\in\mathbb{R}^{L\times L}$
-is indeed the within-class covariance matrix of $Z=B^{\top}X$ under model (1.1).
+The second point follows by noticing that $C_{w}:=B^{\top}\Sigma_{w}B\in\mathbb{R}^{L\times L}$ is indeed the within-class covariance matrix of $Z=B^{\top}X$ under model (3.1).
 Due to the fact that $C_{w}$ is rank deficient with ${\rm rank}(C_{w})<L$ and for the purpose of extracting fewer discriminant directions, we adopt the equivalent formulation of (4.2) from the Fisher’s perspective in the sequel.
-Write the between-class covariance matrix of $Z=B^{\top}X$ as
-$C_{b}:=B^{\top}{M}D_{\pi}{M}^{\top}B.$
-For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the optimization problem
+Write the between-class covariance matrix of $Z=B^{\top}X$ as $C_{b}:=B^{\top}{M}D_{\pi}{M}^{\top}B.$ For a given reduced dimension $1\leq K<L$, we can find $K$ discriminant directions in the space of $Z$ by solving, for each $k\in[K]$, the optimization problem
 $\begin{split}\alpha_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~\alpha^{\top}C_{b}~\alpha\quad\text{ s.t.}~\alpha^{\top}C_{w}~\alpha=1,~\alpha^{\top}C_{w}~\alpha_{i}=0,~~\forall~i<k.\end{split}$
@@
 In practice, after computing $\widehat{B}$ from (1.10), we propose to estimate $C_{b}$ and $C_{w}$ by
-$\widehat{C}_{b}=\widehat{B}^{\top}\widehat{M}D_{\widehat{\pi}}\widehat{M}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}P_{{\mathbf{Y}}}{\mathbf{X}}\widehat{B},\qquad\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}(\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B},$
+$\widehat{C}_{b}~=\widehat{B}^{\top}\widehat{M}D_{\widehat{\pi}}\widehat{M}^{\top}\widehat{B}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}P_{{\mathbf{Y}}}{\mathbf{X}}\widehat{B}$,
+$\widehat{C}_{w}={1\over n}\widehat{B}^{\top}{\mathbf{X}}^{\top}(\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B}$,
 with $P_{{\mathbf{Y}}}={\mathbf{Y}}({\mathbf{Y}}^{\top}{\mathbf{Y}})^{-1}{\mathbf{Y}}^{\top}$.
@@
 $\begin{split}\widehat{\alpha}_{k}=\operatorname*{arg\,max}_{\alpha\in\mathbb{R}^{L}}&~\alpha^{\top}\widehat{C}_{b}~\alpha\quad\text{ s.t.}~\alpha^{\top}\widehat{C}_{w}~\alpha=1,~\alpha^{\top}\widehat{C}_{w}~\widehat{\alpha}_{i}=0,~~\forall~i<k,\end{split}$ (4.4)
-By similar reasoning for solving the Fisher’s discriminant analysis in (1.6), we propose to solve the optimization problem (2.17) via eigen-decomposing
-$[\widehat{C}_{w}^{+}]^{1\over 2}\widehat{C}_{b}[\widehat{C}_{w}^{+}]^{1\over 2}.$
-The eigenvalues of the above matrix can also be used to determine the upper bound for $K$. For instance, $K$ should be chosen no greater than the rank of the above matrix.
+By similar reasoning for solving the Fisher’s discriminant analysis in (1.4), we propose to solve the optimization problem (4.4) via eigen-decomposing $\widehat{C}_{w}^{+}]^{1\over 2}\widehat{C}_{b}[\widehat{C}_{w}^{+}]^{1\over 2}.$ The eigenvalues of the above matrix can also be used to determine the upper bound for $K$.
+For instance, $K$ should be chosen no greater than the rank of the above matrix.
 Remark 3 (Numerical stability of computing (4.4)).
-The procedure in (2.17) computes Fisher’s discriminant directions $(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K})$ in
-$\text{span}({\mathbf{X}}\widehat{B})\subset\text{span}({\mathbf{X}})$
-of dimension less than $L$. Consequently, as long as
-${\rm rank}((\mathbf{I}_{n}-P_{{\mathbf{Y}}}){\mathbf{X}}\widehat{B})={\rm rank}({\mathbf{X}}\widehat{B}),$ (2.18)
-we do not suffer numerically in solving (2.17), in stark contrast with the algorithmic issue of solving (1.6) (or a penalized version thereof) in the original $p$-dimensional space.
+The procedure in (4.4) computes Fisher’s discriminant directions $(\widehat{\alpha}_{1},\ldots,\widehat{\alpha}_{K})$ within the subspace $\text{span}({\mathbf{X}}\widehat{B})\subset\text{span}({\mathbf{X}})$, whose dimension is strictly less than $L$.
+Consequently, as long as
+[…]
+we do not encounter numerical difficulties when solving (4.4), in stark contrast with the algorithmic issue of solving (1.4) (or a penalized version thereof) in the original $p$-dimensional space.
 As pointed out by Wu et al.
@@
 Finally, we remark that the requirement in (4.5) is easy to satisfy as ${\rm rank}({\mathbf{X}}\widehat{B})\leq L$ and ${\rm rank}(\mathbf{I}_{n}-P_{{\mathbf{Y}}})=n-L$ when we observe at least one data point per class.
-3 Theoretical guarantees
-In this section we provide a unified theory for analyzing the classifier $\widehat{g}$ in (2.13) based on estimating the regression coefficient matrix.
-In Section 3.1 we start with full generality and bound the excess risk of $\widehat{g}$ by a quantity that is related with the error of estimating $B^{*}$.
-The results there hold for any estimator $\widehat{B}^{*}$ of $B^{*}$.
-In Section 3.2 we focus on the proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ and provide a unified analysis of its estimation error that is valid for any generic estimator $\widehat{B}$.
-Finally we apply the general result to two particular choices of $\widehat{B}$, the lasso estimator in Section 3.3 and the reduced-rank estimator in Section 3.4.
-3.1 A reduction scheme for bounding the excess risk
-Let $(X,Y)$ be a new pair from (1.1) that is independent of the training data ${\bm{D}}:=\{{\mathbf{X}},{\mathbf{Y}}\}$.
-For any classifier $\widehat{g}:\mathbb{R}^{p}\to[L]$ in (2.13), its excess risk relative to the error rate of the Bayes rule in (1.2) is defined as
-$\mathcal{R}(\widehat{g}):=\mathbb{P}\{Y\neq\widehat{g}(X)\mid{\bm{D}}\}-\mathbb{P}\{Y\neq g^{*}(X)\}.$
-Note that we use the identification $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$ in the above definition.
-Our analysis of $\mathcal{R}(\widehat{g})$ requires the following condition on the priors $\pi=(\pi_{1},\ldots,\pi_{L})$.
-Assumption 1.
-There exists some absolute constants $0<c\leq C<\infty$ such that
-${c\over L}\leq\min_{k\in[L]}\pi_{k}\leq\max_{k\in[L]}\pi_{k}\leq{C\over L}.$ (3.1)
-This is a common regularity assumption even in binary classification problems.
-We may relax Assumption 1 to $\min_{k\in[L]}\pi_{k}\geq c\log(n)/n$ in our analysis (see, for instance, Lemma 21 of Appendix D) at the cost of more technical proofs and a less transparent presentation.
-The following proposition states that for the purpose of bounding $\mathcal{R}(\widehat{g})$, it suffices to study the estimation error of each discriminant function, that is, $\widehat{G}_{\ell}(X)-G_{\ell}(X)$ for $\ell\in[L]$.
-The proof essentially follows the argument of proving Theorem 12 in Bing and Wegkamp (2023) but with modifications to capture the sum of squared errors of estimating all discriminant functions.
-Proposition 2.
-Under model (1.1) with 1, for any sequences $t_{1},\ldots,t_{L}$ (that possibly depend on ${\bm{D}}$), with probability equal to one, one has
-$\mathcal{R}(\widehat{g})~{}\leq~{}C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\}.$
-Proof.
-See Section B.2. ∎
-2 is valid for any estimator $\widehat{G}_{\ell}$ of the discriminant function, in particular, for our proposed $\widehat{G}_{\ell}$ in (2.12).
-The choice of $(t_{1},\ldots,t_{L})$ depends on the estimation error of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
-We therefore proceed to analyze
-$\widehat{G}_{\ell}(X)-G_{\ell}(X)$ $=(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}+(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})+2\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)$ (3.2)
-for each $\ell\in[L]$.
-It is worth mentioning that in this section we allow $\widehat{B}^{*}$ in the above display to be any estimator of $B^{*}$ whereas $\widehat{\mu}_{\ell}$ and $\widehat{\pi}_{\ell}$ are given in (2.9).
-Since the variance of the first term in (3.2) depends on $\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$, we define
-$\Delta_{\infty}:=\max_{\ell\in[L]}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
-This quantity essentially reflects the largest Mahalanobis distance in the feature space between any two label classes as can be seen from the relation (under $\mathbb{E}(X)=0_{p}$)
-${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}).$
-1, the Gaussian tail of $X\mid Y$ and the independence between $X$ and $({\mathbf{X}},{\mathbf{Y}})$ in (3.2) entail the following upper bound of $|\widehat{G}_{\ell}(X)-G_{\ell}(X)|$.
-Theorem 3.
-Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$ and $L\log(n)\leq n$. With probability at least $1-3n^{-1}$, we have, for all $\ell\in[L]$,
-$\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|~{}\lesssim~{}\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}$ $+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}$
-$+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
-Proof.
-See Section B.3. ∎
-Condition $L\log(n)\leq n$ together with 1 ensures that $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
-The reasonable condition $\Delta_{\infty}\geq 1$ requires that the largest separation among all pairs of conditional distributions does not vanish.
-It is assumed to simplify the presentation, and our analysis can be easily extended to the case $\Delta_{\infty}\to 0$ as $n\to\infty$.
-Combining 2 with Theorem 3 immediately yields the following corollary.
-Denote by $\mathbb{E}_{{\bm{D}}}$ the expectation taken with respect to the training data ${\bm{D}}$ only.
-For any estimator $\widehat{B}^{*}$ of $B^{*}$, define the following (random) quantity
-$\mathcal{Q}(\widehat{B}^{*}):=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}(\Delta_{\infty}+\log n)+\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}.$ (3.3)
-Corollary 4.
-Under model (1.1) with 1, assume $\Delta_{\infty}\geq 1$. For any estimator $\widehat{B}^{*}$ of $B^{*}$, the corresponding classifier $\widehat{g}$ in (2.13) satisfies
-$\mathbb{E}_{\bm{D}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\mathbb{E}_{{\bm{D}}}\left[\min\left\{{\Delta_{\infty}}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right].$
-Proof.
-See Section B.4. ∎
-4 ensures that bounding $\mathcal{R}(\widehat{g})$ can be reduced to controlling $\mathcal{Q}(\widehat{B}^{*})$ for any estimator $\widehat{B}^{*}$ of $B^{*}$.
-In particular, 4 is valid for any existing approach based on estimating $B^{*}$.
-In the next section we provide a unified analysis of $\mathcal{Q}(\widehat{B}^{*})$ for our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ based on a generic $\widehat{B}$.
-As shown later, the first term in (3.3) is typically the dominating term.
-3.2 A unified analysis of the estimation of $B^{*}$ based on generic regression coefficient estimation
-In this section we provide more explicit bounds of $\mathcal{Q}(\widehat{B}^{*})$ in (3.3) for the estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{+}$ with $\widehat{H}$ obtained from (2.10).
-Our results in this section are valid for any estimator $\widehat{B}$ of $B$ that satisfies the following assumption.
-Assumption 2.
-There exists some deterministic sequences $\omega_{1}$ and $\omega_{2}$ such that the following holds with probability at least $1-n^{-1}$,
-$\max\left\{n^{-1/2}\|{\mathbf{X}}(\widehat{B}-B)\|_{F},~{}\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}\right\}\leq\omega_{2},\quad\|\widehat{B}-B\|_{1,2}\leq\omega_{1}.$
-We need the following regularity conditions on $\Sigma_{w}$. It assumes the largest entry of $\Sigma_{w}$ is bounded from above and the smallest eigenvalue of $\Sigma_{w}$ is bounded away from zero.
-Assumption 3.
-There exists some positive constants $0<c\leq C<\infty$ such that $c<\lambda_{p}(\Sigma_{w})\leq\|\Sigma_{w}\|_{\infty}\leq C$.
-Theorem 5.
-Under model (1.1) with Assumptions 1 & 3, assume $\Delta\asymp 1$ and $L\log(n)\leq cn$ for some small constant $c>0$.
-For any $\widehat{B}$ satisfying 2 with $\omega_{2}\sqrt{L}\leq c,$ with probability at least $1-2n^{-1}$,
-the matrix $\widehat{H}$ in (2.10) is invertible,
-and
-$\mathcal{Q}(\widehat{B}^{*})\lesssim~{}L^{2}{\log^{2}(n)\over n}+\omega_{2}^{2}L^{2}\log(n)+\omega_{1}^{2}L^{3}{\log(n\vee p)\over n}.$ (3.4)
-Proof.
-See Section B.5. ∎
-Theorem 5 relates $\mathcal{Q}(\widehat{B}^{*})$ to the estimation error of $\widehat{B}$ in both the Frobenius norm and the $\ell_{1}/\ell_{2}$-norm.
-In the bound of (3.4), the second term on the right-hand-side is usually the leading term and its rate depends on specific choice of $\widehat{B}$.
-We provide concrete examples in Section 3.3 for the lasso estimator and in Section 3.4 for the reduced-rank estimator.
-Condition $\omega_{2}\sqrt{L}\leq c$ ensures that, with high probability, $\lambda_{K}(\widehat{H})\geq\lambda_{K}(H)/2$ holds so that $\widehat{H}$ is invertible.
-Furthermore, the bound in (3.4) suggests that the condition $\omega_{2}\sqrt{L}\leq c$ is also needed for consistent classification.
-3.3 Application to estimation of $B$ with entry-wise $\ell_{1}$-regularization
-In this section we focus on the case where the columns of $B$ are (potentially) sparse and allow for different sparsity patterns across its columns.
-Therefore, for some positive integer $s\leq p$, we consider the following parameter space of $B$:
-$\Theta_{s}:=\left\{B\in\mathbb{R}^{p\times L}:\max_{\ell\in[L]}\|B_{\ell}\|_{0}\leq s\right\}.$
-For ease of presentation, we simply use the largest number of non-zero entries among all columns of $B$.
-However, our analysis can be easily extended to $\|B_{\ell}\|_{0}\leq s_{\ell}$ by allowing different $s_{\ell}$ for $\ell\in[L]$.
-Our analysis also needs the following mild regularity condition on the matrix ${M}$ of conditional means.
-Assumption 4.
-There exists an absolute constant $C<\infty$ such that $\|{M}\|_{\infty}\leq C$.
-The following theorem establishes explicit rates of $\omega_{2}$ and $\omega_{1}$ in 2 for the regression matrix estimator $\widehat{B}$ in (1.14) using the lasso-penalty in (2.7).
-Theorem 6.
-Under model (1.1) with Assumptions 1, 3 and 4, assume that there exists some sufficiently small constant $c>0$ such that $B\in\Theta_{s}$ for some $1\leq s\leq p$ with $(s\vee L)\log(n\vee p)\leq c~{}n.$ Let the estimator $\widehat{B}$ be obtained from (1.14) with any $\lambda$ satisfying
-$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (3.5)
-and some finite, large enough constant $C>0$. Then, with probability at least $1-3n^{-1}$, 2 holds for
-$\omega_{2}\asymp\lambda\sqrt{sL},\qquad\omega_{1}\asymp\lambda sL.$
-Proof.
-See Section B.6. ∎
-Our proof is in fact based on a weaker condition than 3 by replacing $\lambda_{p}(\Sigma_{w})$ with the Restricted Eigenvalue condition (RE) on $\Sigma_{w}$.
-The two main difficulties in our proofs are (a) to control $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$, and (b) to bound from below the restricted eigenvalue of the Gram matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$.
-For (a), as pointed out in Gaynanova (2020), the difficulty in this model is elevated relative to the existing analysis in sparse linear regression models due to the fact ${\mathbf{Y}}$ is not a linear model in ${\mathbf{X}}$.
-We establish a sharp control of $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}$ in Lemma 11.
-For (b), the existing result in Rudelson and Zhou (2012) is not readily applicable as rows of ${\mathbf{X}}$ are generated from a mixture of Gaussians.
-In Lemma 12 we establish a uniform bound of $v^{\top}(\widehat{\Sigma}-\Sigma)v$ over $s$-sparse vectors $v$ which, in conjunction with the reduction arguments in Rudelson and Zhou (2012), proves (b).
-Our analysis in this paper allows the number of classes $L$ to grow with the sample size $n$
-Gaynanova (2020) considers a group-lasso regularized regression similar to (1.14) and derives the bound of the $\ell_{2}$-norm $\max_{j\in[p]}\|{\bm{e}}_{j}^{\top}{\mathbf{X}}^{\top}({\mathbf{Y}}^{\prime}-{\mathbf{X}}B^{\prime})\|_{2}$ for some different response matrix ${\mathbf{Y}}^{\prime}$ and different target matrix $B^{\prime}$.
-Their proof is different and in particular requires that $L$ is fixed, independent of $n$.
-In the rest of the paper, we assume that $\lambda$ is chosen as the order specified in (3.5).
-In practice, one can simply choose $\lambda$ via cross-validation, for instance, via the built-in function of the R-package glmnet.
-As an immediate corollary of 4, Theorem 5 and Theorem 6, we have the following upper bound of the misclassification rate of $\widehat{g}$ based on the entry-wise lasso estimator with $\lambda$ chosen as the rate in (3.5).
-Corollary 7.
-Under the conditions in Theorem 6, assume that $\Delta_{\infty}\asymp 1$. Then
-$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{sL^{2}\over n}\log^{2}(n\vee p),~{}1\right\}.$
-Our results allow both the sparsity level $s$ and the number of classes $L$ to grow with the sample size $n$.
-The existing literature (Cai and Zhang, 2019; Gaynanova, 2020) that analyze the excess risk only consider the case $L\asymp 1$, while Abramovich and Pensky (2019) does allow $s$, $L$ and $n$ to grow, but studies the risk $\mathbb{P}\{Y\neq g(X)\}$ instead of the excess risk $\mathcal{R}(g)$.
-Guarantees on the excess risk are stronger since they together with the Bayes error, which is oftentimes easy to calculate, imply guarantees of the risk.
-7 focuses on the statistically most interesting case $\Delta_{\infty}\asymp 1$.
-Otherwise, if the pairwise separation between different classes grows to infinity, the classification problem becomes much easier and we have super-fast rates of convergence (see, for instance, Cai and Zhang (2019); Bing and Wegkamp (2023)).
-Indeed, provided $\min_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\asymp\max_{k\neq\ell}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell}),$ using the arguments in the proof of Bing and Wegkamp (2023, Theorem 12), our analysis can be modified easily to show that
-$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}{sL^{2}\over n}\log^{2}(n\vee p)e^{-c\Delta_{\infty}}+n^{-C_{0}}$
-for some arbitrary large constant $C_{0}>0$ and some constant $c>0$. The first term on the right in the above rate is exponentially fast in $\Delta_{\infty}\to\infty$.
-3.4 Application to estimation of $B$ via reduced-rank regression
-We provide analysis for the classifier that estimates the regression coefficient matrix $B$ by reduced-rank regression.
-Specifically, let $\widehat{B}$ be obtained from (1.14) with $\text{pen}(B)=\lambda~{}{\rm rank}(B)$ for some tuning parameter $\lambda>0$.
-The following theorem establishes the rates of $\omega_{2}$ and $\omega_{1}$ for which 2 holds.
-Write $r={\rm rank}(B)$.
-Theorem 8.
-Under model (1.1) with Assumptions 1 and 3, assume $(p+L)\log(n)\leq c~{}n$ for some sufficiently small constant $c>0$. For the estimator $\widehat{B}$ in (1.14) with any
-$\lambda\geq C\sqrt{\|\Sigma_{w}\|_{\rm op}(1+\Delta_{\infty})}\sqrt{(p+L)\log(n)\over nL}$ (3.6)
-and some finite, large enough constant $C>0$, with probability at least $1-2n^{-1}$, 2 holds for
-$\omega_{2}\asymp\sqrt{\lambda r},\qquad\omega_{1}\asymp\sqrt{\lambda rp}.$
-Proof.
-The proof is deferred to Section B.7. ∎
-For the reduced-rank estimator, establishing its in-sample prediction risk can be done without any condition on the design matrix (see, for instance, Bunea et al.
-(2011); Giraud (2011)).
-However, since 2 is also related with the out-of-sample prediction risk, we need the smallest eigenvalue of the Gram matrix $\widehat{\Sigma}$ to be bounded away from zero.
-The inequality $(p+L)\log(n)\leq cn$ is assumed for this purpose.
-In case $p$ is large, it is recommended that an additional ridge penalty be added in $\text{pen}(B)$ to alleviate the singularity issue of $\widehat{\Sigma}$ (Mukherjee and Zhu, 2011).
-Another main ingredient of our proof of Theorem 8 is to control $\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}$ in Lemma 13.
-It requires non-standard analysis for the same reason mentioned before that ${\mathbf{Y}}$ is not linearly related with ${\mathbf{X}}$.
-Combining 4, Theorem 5 and Theorem 8 immediately yields the following guarantee on the excess risk of the classifier using the reduced-rank estimator of $B$.
-Corollary 9.
-Under the conditions in Theorem 8, assume $\Delta_{\infty}\asymp 1$ and $\|\Sigma_{w}\|_{\rm op}\lesssim 1$. Then
-$\mathbb{E}_{{\bm{D}}}\left[\mathcal{R}(\widehat{g})\right]~{}\lesssim~{}\min\left\{{(p+L)rL\over n}\log^{2}(n),~{}1\right\}.$
-Note that for the $p\times L$ matrix $B$ with rank $r$, its effective number of parameters is exactly $(p+L)r$ (Izenman, 1975) which could be much smaller than $pL$ when $r\ll(p\wedge L)$.
-9 suggests that the classifier using the reduced-rank estimator has benefit when $r$ is relatively small comparing to $p$ and $L$.
-The same remark as the one at the end of Section 3.3 applies, and our analysis can be extended to $\Delta_{\infty}\to\infty$, where the rate in 9 becomes exponentially fast in $\Delta_{\infty}$.
-4 Simulation study
-In our simulation study, we evaluate our proposed procedure in Section 2.2 that we term as Linear Discriminant Regularized Regression (LDRR).
-We also include the procedure in Section 2.3 that is based on Fisher’s discriminant rule after estimating $B$ (called LDRR-F).
-For estimating $B$ from (1.14), in the sparse scenarios of Section 4.1, we consider both the lasso penalty (2.7) (L1) and the elastic net penalty (L1+L2), $\text{pen}(B)=\lambda\sum_{\ell=1}^{L}[\alpha\|B_{\ell}\|_{1}+(1-\alpha)/2\|B_{\ell}\|_{2}^{2}]$, for some $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation from the R-package glmnet.
-In the low-rank scenarios of Section 4.2, we consider both the reduced-rank penalty (2.8) (RR) and the reduced-rank penalty plus the ridge penalty (RR+L2), $\text{pen}(B)=\lambda[\alpha~{}{\rm rank}(B)+(1-\alpha)/2\|B\|_{F}^{2}]$, with tuning parameters $\alpha\in[0,1]$ and $\lambda>0$ chosen by cross-validation.
-We compare our proposed methods with the nearest shrunken centroids classifier (PAMR) of Tibshirani et al.
-(2002), the shrunken centroids regularized discriminant analysis (RDA) method of Guo et al.
-(2007) and the $\ell_{1}$-penalized linear discriminant (PenalizedLDA) method of Witten and Tibshirani (2011).
-These methods are available in R-packages pamr, rda, penalizedLDA, and we use their default methods for selecting tuning parameters.
-Finally, the oracle procedure that uses the true values of ${M}$, $\Sigma_{w}$, and $\pi$ in computing the discriminant functions $G_{k}(x)$ in (2.1) serves as our benchmark.
 4.2 Sparse scenarios
@@
 A similar setting is also used in Witten and Tibshirani (2011).
-Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $[\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0.
+Specifically, for any $\ell\in[L]$ and its corresponding mean vector $\mu_{\ell}\in\mathbb{R}^{p}$, we sample $\mu_{\ell}]_{j}$, for $5(\ell-1)\leq j\leq 5\ell$, independently from $N(0,2^{2})$ and set the other entries of $\mu_{\ell}$ to 0.
 Regarding the within-class covariance matrix, we set $\Sigma_{w}=\sigma^{2}W$ where $W$ is generated by independently sampling its diagonal elements from Uniform(1,3), and setting its off-diagonal elements as $W_{ij}=\sqrt{W_{ii}W_{jj}}\rho^{|i-j|}$, for all $i\neq j$.
@@
 For a given $\alpha\geq 0$, we set $\pi_{\ell}=\nu_{\ell}^{\alpha}/(\sum_{i=1}^{L}\nu_{i}^{\alpha})$ for all $\ell\in[L]$.
-Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform$(0,1)$.
+Here $\nu_{1},\ldots,\nu_{L}$ are independent draws from the Uniform(0,1)(0,1).
 The quantity $\alpha$ controls the imbalance of each class size: a larger $\alpha$ corresponds to more imbalanced class sizes.
@@
 From Figure 1, it is clear that our procedures consistently outperform other methods across all settings.
-We observe that the performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases.
+The performance of our algorithms improves as $n$ increases, $L$ decreases and $\sigma$ decreases.
 Moreover, neither the imbalance of the class probabilities nor the ambient dimension $p$ seems to affect the misclassification error of our algorithms.
@@
 This is as expected as the data is simulated according to model (3.1), although in practice the Fisher’s discriminant analysis (LDRR-F) with the elastic net penalty could have more robust performance, as revealed in our real data analysis in the next section.
-Figure 1: The averaged misclassification errors of each algorithm for varying each parameter one at a time.
+Figure 1: The averaged misclassification errors in sparse scenarios.
 4.3 Low-rank scenarios
 In this section we evaluate our proposed procedure when the rank of the conditional mean matrix $M$ is low and hence $B$ is also low-rank.
-We consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Appendix A for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank.
-Specifically, let the within-class covariance matrix $[\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$.
+4.3.1 Low-rank model (1)
+We first consider a general within-class covariance matrix $\Sigma_{w}$ here, and defer to Section 4.3.2 for the factor model mentioned in Remark 1 where $\Sigma_{w}$ has approximately low-rank.
+Specifically, let the within-class covariance matrix $\Sigma_{w}]_{ij}=\rho^{|i-j|}$ with $\rho=0.6$.
 To generate $M$, for some scalar $\eta>0$ and some randomly generated orthogonal matrix $A\in\mathbb{R}^{p\times r}$ with $A^{\top}A=\mathbf{I}_{r}$, we set $M=\eta A\alpha$ with entries of $\alpha\in\mathbb{R}^{r\times L}$ generated i.i.d.
@@
 The quantity $\eta$ controls the magnitude of separation between the conditional means.
-We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$.
+We examine how the performance of all algorithms depends on various parameters by varying one at a time: the number of classes $L\in\{5,15,25,35,45\}$; the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the separation scalar $\eta\in\{0.3,0.6,0.9,1.2,1.5\}$.
 For each setting, we fix $n=1000,L=10,r=3,p=100$ and $\eta=1$ when they are not varied.
@@
 For relative large $p$, incorporating the ridge penalty is beneficial as it improves the estimation of $B$ when the sample covariance matrix $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ becomes ill-conditioned.
-Figure 2: The averaged misclassification errors of each algorithm.
-5 Real data analysis
-In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets and one low-dimensional football dataset.
-The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001).
-The SRBCT dataset (Khan et al., 2001) has $n=83$ samples and consists of microarray analyses of childhood cancer tumors, more specifically, small round blue cell tumors (SRBCT).
-These tumors are categorized into $L=4$ classes: Burkitt lymphoma, Ewing sarcoma, neuroblastoma, and rhabdomyosarcoma.
-Each sample has $p=2,308$ gene expressions as features.
-The third dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes.
-These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia.
-Each sample has $p=4026$ gene expressions as its features.
-We used the preprocessed Lymphoma data in Dettling (2004).
-The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes.
-We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction.
-We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$.
-For all four data sets, the features are centered to have zero mean.
-Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data.
-This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
-Table 1: The averaged misclassification errors (in percentage) and standard errors
-LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA
-Ramaswamy 30.7(0.051) 14.4(0.054) 22.6(0.048) 40.0(0.058) 30.4(0.058) 16.7(0.05)
-SRBCT 0.7(0.022) 0.3(0.011) 0.1(0.007) 4.4(0.052) 1.0(0.019) 0.0(0.0)
-Lymphoma 0.3(0.012) 0.5(0.017) 0.4(0.015) 1.9(0.034) 1.3(0.025) 1.3(0.025)
-Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014)
-Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification.
-We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison.
-From Table 1, we see that our procedure LDRR-F is close to have the best performance across all datasets.
-Choice of the best penalty depends on the data structure.
-For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent.
-For the Football dataset with $p=118$ and $L=10$, the selected rank is found to be $\widehat{r}=3$.
-Furthermore, in Figure 3, we show that LDRR-F with $K=2$ in (2.16) can be used for visualization in the 2D space of the first two discriminant directions.
-The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
-Figure 3: Visualization of Ramaswamy, SRBCT, Lymphoma and Football data points in the space of the first two discriminant vectors.
-Acknowledgements
-Wegkamp is supported in part by the National Science Foundation grant NSF DMS-2210557.
-Appendix A Additional simulation study when the conditional mean matrix is low-rank and $\Sigma_{w}$ is approximately low-rank
+Figure 2: The averaged misclassification errors in low-rank model (1).
+4.3.2 Low-rank model (2).
 In this setting, we consider the setting in Remark 1 and also compare with the method (PCLDA) in Bing and Wegkamp (2023).
 As mentioned in Remark 1, PCLDA is a two-step procedure similar to LDRR-F but with the first step done in a unsupervised way by using PCA.
-We generate the data according to model (1.1) with $M=\eta A\alpha$ as in Section 4.2 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $[\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.1 with $\rho=0.2$ and its diagonal elements from Uniform(0,1).
+We generate the data according to model (3.1) with $M=\eta A\alpha$ as in Section 4.3 and $\Sigma_{w}=\eta^{2}A\Sigma_{z}A^{\top}+{\rm Cov}(W)$ where $\Sigma_{z}]_{ij}=(0.6)^{|i-j|}$ for $i,j\in[r]$ and ${\rm Cov}(W)$ is generated the same as $\Sigma_{w}$ in Section 4.2 with $\rho=0.2$ and its diagonal elements from Uniform(0,1).
 The quantity $\eta$ represents the signal-to-noise ratio (SNR) between the low-dimensional signal $\eta AZ$ and the noise $W$.
 As pointed out in Bing and Wegkamp (2023), classification becomes easier as $\eta$ increases.
-We examine how the performance of all algorithms depends on different parameters by varying one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$.
+We examine how the performance of all algorithms depends on different parameters by varying the parameters one at a time: the sample size $n\in\{300,500,700,900,1100\}$; the feature dimension $p\in\{100,200,300,400,500\}$; the number of classes $L\in\{5,15,25,35,45\}$; the scalar $\eta\in\{0.5,1.0,2.0,4.0,6.0\}$.
 For all settings, we fix $n=1000,L=10,r=3,p=100$ and $\eta=2$ when they are not varied.
@@
 But when $\eta$ is small such as $\eta<2$, LDRR-F has benefit by reducing the dimension in a supervised way.
-Figure 4: The averaged misclassification errors of each algorithm.
-Appendix B Proofs
-Throughout the proofs, we will use the following notation. Define
-$\Delta={M}^{\top}\Sigma_{w}^{-1}{M}$
-and write
-$\Delta_{\infty}=\max_{\ell\in[L]}\Delta_{\ell\ell},\qquad\Delta_{\rm op}=\|\Delta\|_{\rm op}.$
-Further write
+Figure 3: The averaged misclassification errors in low-rank model (2).
+5 Real data analysis
+In this section we evaluate the performance of our approach and compare it with other algorithms on three high-dimensional biological data sets and one low-dimensional football dataset.
+The Ramaswamy dataset has $n=198$ samples and each sample has features of $p=16,063$ gene expressions and belongs to one of the $L=14$ distinct cancer subtypes (Ramaswamy et al., 2001).
+The second dataset has $n=62$ samples and consists of $L=3$ distinct lymphoma subtypes.
+These lymphoma are categorized into diffuse large B-cell lymphoma, follicular lymphoma and chronic lymphocytic leukemia.
+Each sample has $p=4026$ gene expressions as its features.
+We used the preprocessed Lymphoma data in Dettling (2004).
+The Football dataset from kaggle [^1] https://www.kaggle.com/datasets/vivovinco/20222023-football-player-stats/data contains 2022 – 2023 football player stats per 90 minutes.
+We excluded categorical data such as the player’s nationality, name, team name, and league from our prediction.
+We used age, year of birth and performance data $(p=118)$ to predict the $(n=2689)$ football player’s position $(L=10)$.
+Brain A is another dataset used in (Dettling, 2004).
+It has sample size $n=42$ with $p=5597$ features and contains $L=5$ different tumor types.
+For all four data sets, the features are centered to have zero mean.
+Each dataset is randomly divided with 75% of the samples to serve as the training data and the remaining 25% serves as the test data.
+This randomization was repeated 50 times and the averaged misclassification errors of each algorithm are reported in Table 1.
+Table 1: The averaged misclassification errors (in percentage) and standard errors
+LDRR-F(L1+L2) LDRR-F(L12+L2) LDRR-F(RR+L2) PenalizedLDA PAMR RDA SLDTR MSDA
+Ramaswamy 22.3(0.051) 15.9(0.042) 22.9(0.044) 37.4(0.052) 28.6(0.043) 17.1(0.038) 17.9(0.045) 30.4(0.067)
+Lymphoma 1.9(0.029) 3.1(0.051) 1.9(0.029) 3.1(0.044) 2.0(0.029) 1.9(0.029) 1.8(0.037) 5.9(0.050)
+Football 31.5(0.013) 32.2(0.011) 30.5(0.008) 52.6(0.008) 43.8(0.016) 31.9(0.014) 34.11(0.011) 31.6 (0.011)
+Brain A 24.3(0.106) 16.3(0.105) 13.8(0.082) 18.8 (0.107) 18.8(0.090) 19.3(0.080) 12.8(0.062) 39.0(0.134)
+Given the similar performance in Section 4 between LDRR and LDRR-F as well as their variants of using the lasso and the elastic-net penalty, we only compare LDRR-F using the elastic net penalty (L1+L2) and the reduced rank plus ridge penalty (RR+L2) with PenalizedLDA, PAMR and RDA in our real data analysis due to its robustness against model misspecification.
+We also include LDRR-F that uses the group-lasso penalty plus the ridge penalty (L12+L2) for comparison.
+From Table 1, we observe that our procedure LDRR-F exhibits relatively better performance across all datasets.
+Choice of the best penalty depends on the data structure.
+For instance, when the data has the group-sparsity structure such as in Ramaswamy data set, the classifier using the group-lasso penalty has better performance; when the sparsity structure varies across the response levels such as in Lymphoma, the classifier using the elastic-net penalty performs better; when the data meets the low rank structure, the benefit of the classifier using the reduced-rank estimator is apparent (the selected rank is found to be $\widehat{r}=3$ in the Football dataset with $p=118$ and $L=10$).
+We remark that in practice one can also use cross-validation to find the most suitable choice of penalty.
+Furthermore, in Figure 4, we show that LDRR-F with $K=2$ in (4.3) can be used for visualization in the 2D space of the first two discriminant directions.
+The algorithm RDA has comparable performance in all settings, but it cannot be used for low-dimensional representation / visualization.
+Figure 4: Visualization in the space of the first two discriminant vectors.
+Acknowledgements
+Wegkamp is supported in part by the National Science Foundation grant NSF DMS-2210557.
+Appendix A contains the main proofs. Technical lemmas are stated in Appendix B.
+Appendix A Main proofs
+Throughout the proofs, we will use the following notation. Write
+$\Delta_{\infty}=\max_{\ell\in[L]}{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell},\qquad\Delta_{\rm op}=\|{M}^{\top}\Sigma_{w}^{-1}{M}\|_{\rm op}.$
+For future reference, we note that (under $\mathbb{E}(X)=0_{p}$)
+${1\over 4}\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\leq\Delta_{\infty}\leq\max_{k,\ell\in[L]}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})$
+so that $\Delta_{\infty}\asymp\Delta$ under 2. Further, we write
 $\pi_{\min}=\min_{\ell\in[L]}\pi_{\ell},\qquad\pi_{\max}=\max_{\ell\in[L]}\pi_{\ell}.$
 The following event will be frequently used in our proofs. Define the event
-$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (B.1)
+$\mathcal{E}_{\pi}=\left\{\max_{k\in[L]}\left|\widehat{\pi}_{k}-\pi_{k}\right|\leq C\sqrt{\log(n)\over nL}\right\}$ (A.1)
 for some absolute constant $C>0$.
-Lemma 21 in Appendix D ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1.
+Lemma 21 in Section B.2 ensures that $\mathbb{P}(\mathcal{E}_{\pi})\geq 1-2n^{-2}$ under 1.
 Also note that, on $\mathcal{E}_{\pi}$ and under the condition $L\log(n)\leq n$,
-$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (B.2)
+$\max_{\ell\in[L]}\left|\log\left({\pi_{\ell}\over\widehat{\pi}_{\ell}}\right)\right|\leq\max_{\ell\in[L]}{|\widehat{\pi}_{\ell}-\pi_{\ell}|\over\pi_{\ell}}\lesssim\sqrt{L\log(n)\over n}$ (A.2)
 as well as $\widehat{\pi}_{\ell}\asymp\pi_{\ell}$ for all $\ell\in[L]$.
-B.1 Proof of the equivalence between (1.2) and (2.14)
-Proof.
-Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
-$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}$
-so that, for any $\ell\in[L]$,
-$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})-2\log(\pi_{\ell})$
-$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$
-$\qquad+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}-2\log(\pi_{\ell}).$
-Write the singular value decomposition of $\Sigma_{w}^{-1/2}{M}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{M})$.
-It follows that the above equals to
-$\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
-$=\operatorname*{arg\,min}_{\ell\in[L]}~{}2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}-2\log(\pi_{\ell})$
-which, in view of (1.2) and (2.1), completes the proof. ∎
-B.2 Proof of 2
-For any $x\in\mathbb{R}^{p}$, write
-$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
+A.1 Proof of 2
+Recall that, for any $x\in\mathbb{R}^{p}$,
+$\widehat{g}(x)=\operatorname*{arg\,min}_{\ell\in[L]}\widehat{G}_{\ell}(x),\qquad g^{*}(x)=\operatorname*{arg\,min}_{\ell\in[L]}G_{\ell}(x).$
 For any $t_{1},\ldots,t_{L}\geq 0$, define the event
-$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (B.3)
+$\mathcal{E}_{t}=\bigcap_{\ell\in[L]}\left\{\left|\widehat{G}_{\ell}(X)-G_{\ell}(X)\right|\leq t_{\ell}\right\}.$ (A.3)
 In the proof of 2, we use the notation $\{Y=k\}:=\{Y={\bm{e}}_{k}\}$.
 Proof.
-We condition on ${\bm{D}}$ throughout the proof.
-The law of $\mathbb{E}$ and $\mathbb{P}$ is with respect to the randomness of $(X,Y)$, which is independent of ${\bm{D}}$.
-By definition, the excess risk $\mathcal{R}(\widehat{g})$ equals to
+[…]
 $\sum_{k\in[L]}\pi_{k}\Bigl{\{}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)\neq k\}\mid Y=k\right]-\mathbb{E}\left[\mathbbm{1}\{g^{*}(X)\neq k\}\mid Y=k\right]\Bigr{\}}$
@@
 $=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\Bigl{\{}\pi_{k}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=k\right]-\pi_{\ell}\mathbb{E}\left[\mathbbm{1}\{\widehat{g}(X)=\ell,g^{*}(X)=k\}\mid Y=\ell\right]\Bigr{\}}.$
-Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y={\bm{e}}_{k}$ for each $k\in[L]$. We find that
+Write $f_{k}(x)$ as the p.d.f. of $X=x\mid Y=k$ for each $k\in[L]$. We find that
 […]
-[…]
+$=\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\int_{\widehat{g}=\ell,g^{*}=k}\pi_{k}f_{k}(x)\left(1-\exp\left\{G^{(\ell|k)}(x)\right\}\right)dx$
 where, for all $k,\ell\in[L]$,
@@
 By the basic inequality $1+z\leq\exp(z)$ for all $z\in\mathbb{R}$, we obtain that,
-$\mathcal{R}(\widehat{g})$ $\leq\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$
+[…]
 $+{1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{E}\left[\mathbbm{1}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y=k\right\}\right].$
-For the first term, we have
-$\sum_{k\neq\ell}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)$ $\leq(L-1)\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=(L-1)\mathbb{P}(\mathcal{E}_{t}^{c}).$ (B.4)
+For the first term, it is bounded from above by
+$\sum_{k=1}^{L}\pi_{k}\mathbb{P}(\mathcal{E}_{t}^{c}\mid Y=k)=\mathbb{P}(\mathcal{E}_{t}^{c}).$ (A.4)
 Regarding the second term, it is bounded from above by
-${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (B.5)
+${1\over 2}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})\mathbb{P}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0\mid Y=k\right\}.$ (A.5)
 Note that $G^{(\ell|k)}(X)\mid Y=k$ is normally distributed
 $N\left(-{1\over 2}(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})+\log\left(\pi_{\ell}/\pi_{k}\right),~(\mu_{k}-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(\mu_{k}-\mu_{\ell})\right).$
-By the mean-value theorem, the quantity in (B.5) is no greater than
-${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}(t_{k}+t_{\ell})^{2}$ $\leq{1\over 2}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
-Together with (B.4) and $\pi_{k}\leq C/L$, the proof is complete. ∎
-B.3 Proof of Theorem 3
+By the mean-value theorem and 2, the quantity in (A.5) is no greater than
+${1\over 4}\sum_{k=1}^{L}\pi_{k}\sum_{\ell\in[L]\setminus\{k\}}{(t_{k}+t_{\ell})^{2}\over\sqrt{c\Delta}}$ $\leq{1\over 2\sqrt{c\Delta}}\sum_{k=1}^{L}\pi_{k}\left[(L-1)t_{k}^{2}+\sum_{\ell\in[L]\setminus\{k\}}t_{\ell}^{2}\right]={1\over 2\sqrt{c\Delta}}\sum_{k=1}^{L}(L\pi_{k}+1)t_{k}^{2}.$
+Together with (A.4) and $\pi_{k}\leq C/L$, the proof of (3.6) is complete.
+The bound in (3.7) follows by noting that
+${1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{E}\left[\mathbbm{1}\left\{-t_{k}-t_{\ell}\leq 2G^{(\ell|k)}(X)\leq 0,\widehat{g}(X)=\ell\mid Y=k\right\}\right]$
+$\leq{1\over 2}\sum_{\begin{subarray}{c}k,\ell\in[L]\\ k\neq\ell\end{subarray}}\pi_{k}(t_{k}+t_{\ell})\mathbb{P}\left\{\widehat{g}(X)=\ell\mid Y=k\right\}$
+$\leq\max_{\ell\in[L]}t_{\ell}.$
+The proof is complete. ∎
+A.2 Proof of Theorem 3
 Proof.
 Pick any $\ell\in[L]$. Recall that
-$|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\leq|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}|+|(\widehat{\mu}_{\ell}-2X)^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2|\log\left({\pi_{\ell}/\widehat{\pi}_{\ell}}\right)|.$ (B.6)
-Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (B.1) which implies (B.2), an upper bound for the last term on the right of (B.6).
+[…]
+Throughout the proof, we work on $\mathcal{E}_{\pi}$ in (A.1) which implies (A.2), an upper bound for the last term on the right of (A.6).
 Recall that $\mathbb{P}\{\mathcal{E}_{\pi}\}\geq 1-2n^{-2}$ and note that we consider both the randomness of $(X,Y)$ and that of $\bm{D}$} in this proof.
-For the first term in (B.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}=\Delta_{\ell\ell}$ and take a union bound over $\ell\in[L]$ to conclude that the event
-$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~{}\text{for all }\ell\in[L]\right\}$ (B.7)
-has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (B.6)
+For the first term in (A.6), we invoke Lemma 22 with $k=\ell$ and $v=B^{*}_{\ell}$, use the fact that $B_{\ell}^{*\top}\Sigma_{w}B_{\ell}^{*}=\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}\leq\Delta_{\infty}$ and take a union bound over $\ell\in[L]$ to conclude that the event
+$\mathcal{E}_{\mu}:=\left\{\left|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}_{\ell}\right|\leq\sqrt{3\Delta_{\ell\ell}}\sqrt{L\log(n)\over n},~\text{for all }\ell\in[L]\right\}$ (A.7)
+has probability at least $1-2Ln^{-3}\geq 1-n^{-2}$. Regarding the second term in (A.6)
 $|\widehat{\mu}_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|,$
 the Cauchy-Schwarz inequality gives
-$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\ell\ell}}$ (B.8)
+$|\mu_{\ell}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|\leq\|\Sigma_{w}^{-1/2}\mu_{\ell}\|_{2}\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}=\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}}$ (A.8)
 Finally, conditioning on $Y={\bm{e}}_{k}$, the Gaussian tail of $X\mid Y,{\bm{D}}$ gives that, for all $t\geq 0$,
@@
 $|X^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|$ $\leq\sum_{k=1}^{L}\pi_{k}|\mu_{k}^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|+2\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\log(n)}$
-$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (B.9)
-holds with probability at least $1-2Ln^{-2}$. The last step uses (B.8). Collecting (B.2), (B.7), (B.8) and (B.3) completes the proof. ∎
-B.4 Proof of 4
+$\leq\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+4\log(n)}$ (A.9)
+holds with probability at least $1-2Ln^{-2}$. The last step uses (A.8). Collecting 2, (A.2), (A.7), (A.8) and (A.2) and $\Delta_{\infty}\lesssim\Delta$ completes the proof. ∎
+A.3 Proof of 4
 Proof.
 For any $\ell\in[L]$, let
-${t_{\ell}\over C}=\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
+${t_{\ell}\over C}=\sqrt{\Delta}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
 Recall from the proof of Theorem 3 that $1-\mathbb{P}(\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi})\leq 2n^{-2}+n^{-2}=3n^{-2}$ and
@@
 By the fact $\mathcal{R}(\widehat{g})\leq 1$ almost surely and invoking 2, we find that
-$\mathbb{E}_{{\bm{D}}}[\mathcal{R}(\widehat{g})]$ $\leq\mathbb{E}_{{\bm{D}}}\left[\min\left\{C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]$
+[…]
 $\quad+1-\mathbb{P}_{{\bm{D}}}\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}$
-$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
-$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right],$
+$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
+$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~1\right\}\right],$
 completing the proof. ∎
-B.5 Proof of Theorem 5
+A.4 Proof of Theorem 5
 The proof of Theorem 5 uses the following lemma on $\|\widehat{H}-H\|_{\rm op}$. The proof of Lemma 10 appears at the end of this section.
@@
 Define
-$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (B.10)
+$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (A.10)
 for some constant $C>0$. We work on the event
@@
 Lemma 10 ensures that $\mathcal{E}_{H}$ holds with probability at least $1-n^{-1}$. Moreover, on the event $\mathcal{E}_{H}$, we find that
-$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
-$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ by $\text{by }\Omega=H^{-1}$
+[…]
+$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ $\text{by }\Omega=H^{-1}$
 […]
@@
 Hence, on the event $\mathcal{E}_{H}$, $\widehat{H}$ is invertible. We proceed to prove
-$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~{}\lesssim~{}\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (B.11)
-$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (B.12)
-which, in conjunction with (B.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
-To prove (B.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
+$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~\lesssim~\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (A.11)
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (A.12)
+which, in conjunction with (A.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
+To prove (A.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
 $\widehat{B}^{*}-B^{*}$ $=(\widehat{B}-B)\widehat{\Omega}+B(\widehat{\Omega}-\Omega)$
 $=(\widehat{B}-B)\widehat{\Omega}+B^{*}H(\widehat{\Omega}-\Omega)$ by Lemma 1
-$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ by $\text{by }H\Omega=\mathbf{I}_{L}$
-$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (B.13)
+$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ $\text{by }H\Omega=\mathbf{I}_{L}$
+$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (A.13)
 It then follows that
@@
 we conclude
-$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (B.14)
-It now follows from 2 that
+$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (A.14)
+It now follows from 1 that
 $\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}$ $\lesssim\left(\|\Sigma_{w}^{1/2}(\widehat{B}-B)\|_{F}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty}$
-$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (B.15)
-proving (B.11).
-We now prove (B.12). Using the identity (B.5), we have, on the event $\mathcal{E}_{H}$,
+$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (A.15)
+proving (A.11).
+We now prove (A.12). Using the identity (A.4), we have, on the event $\mathcal{E}_{H}$,
 $\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}$
-$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (B.16)
+$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (A.16)
 For the first term on the right, we find
@@
 On the one hand, on the event $\mathcal{E}_{\pi}$, Lemma 22 with $k=\ell$ and $v={\bm{e}}_{j}$, in conjunction with union bounds over $j\in[p]$ and $\ell\in[L]$, yields that
-$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (B.17)
+$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (A.17)
 with probability at least $1-n^{-2}$. On the other hand, we notice that
@@
 $\leq\|\widehat{\Omega}\|_{\rm op}^{2}\|\widehat{B}-B\|_{1,2}^{2}.$
-By invoking 2 and applying inequalities (B.14) and (B.17), we obtain
-$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (B.18)
-with probability at least $1-n^{-2}$. Regarding the second term on the right in (B.5), we have
+By invoking 1 and applying inequalities (A.14) and (A.17), we obtain
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (A.18)
+with probability at least $1-n^{-2}$. Regarding the second term on the right in (A.4), we have
 $\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $\leq\sum_{\ell=1}^{L}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}\|(H-\widehat{H})\widehat{\Omega}_{\ell}\|_{2}^{2}$
@@
 $\lesssim\delta^{2}\max_{\ell\in[L]}\|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}\|_{2}^{2}~L\|\widehat{\Omega}\|_{\rm op}^{2}$
-on the event $\mathcal{E}_{H}$. Further invoking (B.7) and (B.14) yields that
-$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~{}\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (B.19)
-with probability at least $1-n^{-2}$. Combining (B.18) with (B.19) proves (B.12), and hence completes the proof. ∎
+on the event $\mathcal{E}_{H}$. Further invoking (A.7) and (A.14) yields that
+$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (A.19)
+with probability at least $1-n^{-2}$. Recall that $\Delta_{\infty}\leq C\Delta$ from 2. Combining (A.18) with (A.19) proves (A.12), and hence completes the proof. ∎
 Proof of Lemma 10.
 Recall that
-$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (B.20)
+$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (A.20)
 By triangle inequality, we have
@@
 We proceed to bound each term on the right-hand-side separately. On the event $\mathcal{E}_{\pi}$, we first find that
-$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (B.21)
-Regarding the second term, 2 ensures that
-$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (B.22)
+$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (A.21)
+Regarding the second term, 1 ensures that
+$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (A.22)
 For the last two terms, invoking Lemma 19 ensures with probability at least $1-n^{-2}$,
-$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (B.23)
+$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (A.23)
 It also follows that
@@
 […]
-deduced from Lemma 14. Combining the above bounds with (B.21), (B.22) and (B.23) gives
+deduced from Lemma 14. Combining the above bounds with (A.21), (A.22) and (A.23) gives
 $\|\widehat{H}-H\|_{\rm op}$ $\lesssim\sqrt{\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}$
@@
 with probability at least $1-n^{-1}$. The result follows by using $\Delta_{\infty}\asymp 1$ and $\omega_{2}\sqrt{L}\leq c$ to collect terms. ∎
-B.6 Proof of Theorem 6: the lasso estimator of $B$
+A.5 Proof of Theorem 6: the lasso estimator of $B$
 The proof of Theorem 6 uses the following Restricted Eigenvalue condition (RE) on the within-class covariance matrix $\Sigma_{w}$.
@@
 We first prove that
-$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (B.24)
+$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~\lesssim~\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~\lesssim~\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~\lesssim~\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (A.24)
 hold with probability at least $1-2n^{-1}$. We work on the intersection of the event
-$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (B.25)
+$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (A.25)
 with the event
-$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (B.26)
+$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (A.26)
 Recall $\mathcal{C}(S,3):=\{u\in\mathbb{R}^{p}\setminus\{0\}:\|u_{S^{c}}\|_{1}\leq 3\|u_{S}\|_{1}\}$.
@@
 Under 5, we define the space
-$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (B.27)
+$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (A.27)
 for some constant $C=C(\kappa_{s},\|\Sigma_{w}\|_{\infty})>1$.
@@
 Lemma 12 and the condition $(s\vee L)\log(n\vee p)\leq c~n$ imply that
-${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (B.28)
-holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (B.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
-$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (B.29)
-holds for any $u\in\mathcal{C}(S,3)$. Eq. B.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
-It is now relatively straightforward to finish the proof of (B.24).
+${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~v\in\Gamma_{s}$ (A.28)
+holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (A.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
+$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (A.29)
+holds for any $u\in\mathcal{C}(S,3)$. Eq. A.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
+It is now relatively straightforward to finish the proof of (A.24).
 Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
@@
 from which one can deduce that
-$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (B.30)
-with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (B.29), we find on the event $\mathcal{E}_{RE}$
+$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (A.30)
+with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (A.29), we find on the event $\mathcal{E}_{RE}$
 ${\kappa_{s}\over 4}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}^{2}\leq{1\over 4}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}^{2}$ $\leq{1\over 2n}\|{\mathbf{X}}\widehat{B}_{\ell}-{\mathbf{X}}B_{\ell}\|_{2}^{2}$
 $\leq{3\lambda\over 2}\|\widehat{B}_{\ell}-B_{\ell}\|_{1}$
-$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (B.30)
+$\leq 6\lambda\|(\widehat{B}_{\ell}-B_{\ell})_{S_{\ell}}\|_{1}$ by (A.30)
 $\leq 6\lambda\sqrt{s}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}.$
-We have proved (B.24).
+We have proved (A.24).
 Finally, the claim follows by noting that
@@
 The proof is complete. ∎
-B.6.1 Two key lemmas used in the proof of Theorem 6
+A.5.1 Two key lemmas used in the proof of Theorem 6
 Lemma 11.
@@
 Note that
-${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (B.31)
+${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (A.31)
 We first analyze the term $(\widehat{\Sigma}-\Sigma)B$.
@@
 We observe from Lemma 21 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$.
-Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (C.5), (C.6) and (C.7) in Lemma 14 to derive that the inequalities
+Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (B.5), (B.6) and (B.7) in Lemma 14 to derive that the inequalities
 $\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{[\Sigma_{w}]_{jj}B_{\ell}^{\top}\Sigma_{w}B_{\ell}}\sqrt{\log(n\vee p)\over n}$
@@
 hold with probability at least $1-(n\vee p)^{-3}$. After we take the union bound over $j\in[p]$ and $\ell\in[L]$, we have that
-$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (B.32)
+$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (A.32)
 holds with probability $1-n^{-1}$.
-To bound the first term in (B.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
+To bound the first term in (A.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
 ${1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}$ $={1\over n}\sum_{k=1}^{L}\sum_{i=1}^{L}\left(X_{ij}Y_{i\ell}-{M}_{j\ell}\pi_{\ell}\right)\mathbbm{1}\{Y_{i}={\bm{e}}_{\ell}\}$
@@
 Invoking 1, Lemma 21 and Lemma 22 with $t=\sqrt{\log(p\vee n)}$ and taking a union bound over $j\in[p]$ and $\ell\in[L]$, we obtain
-$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (B.33)
-with probability $1-n^{-1}$. Combining (B.32) and (B.33) completes the proof. ∎
+$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (A.33)
+with probability $1-n^{-1}$. Combining (A.32) and (A.33) completes the proof. ∎
 Recall the set
@@
 and
-$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (B.34)
+$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (A.34)
 Here
@@
 $|\mathcal{N}|\leq\sum_{|S|=\lfloor Cs\rfloor}|\mathcal{N}_{S}|\lesssim{p}^{Cs}.$
-We will bound the right hand side (B.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$.
+We will bound the right hand side (A.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$.
 Specifically, repeating the proof of Lemma 18 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
@@
 $\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v\leq\max_{k\in[L]}{|\widehat{\pi}_{k}-\pi_{k}|\over\pi_{k}}v^{\top}{M}D_{\pi}{M}^{\top}v\lesssim v^{\top}{M}D_{\pi}{M}^{\top}v\sqrt{L\log(n)\over n}$
-holds uniformly over $\mathcal{N}$. The proof is complete in view of (B.34). ∎
-B.7 Proof of Theorem 8: the reduced-rank estimator of $B$
+holds uniformly over $\mathcal{N}$. The proof is complete in view of (A.34). ∎
+A.6 Proof of Theorem 8: the reduced-rank estimator of $B$
 Proof.
@@
 $\mathcal{E}_{X}=\left\{\lambda_{p}(\widehat{\Sigma})\geq{1\over 2}\lambda_{p}(\Sigma)\geq{c\over 2}\right\}.$
-Here $c$ is defined in 3.
-Lemma 20, Lemma 13 and 3 together with (3.6) ensure that $\mathcal{E}_{X}\cap\mathcal{E}_{\lambda}^{\prime}$ hold with probability at least $1-n^{-1}$.
-Write $\widehat{r}={\rm rank}(\widehat{B})$ and $r={\rm rank}(B)$.
-By the optimality of $\widehat{B}$, we have
+[…]
 ${1\over n}\|{\mathbf{Y}}-{\mathbf{X}}\widehat{B}\|_{F}^{2}+\lambda\widehat{r}$ $~\leq~{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\lambda r.$
 Working out the squares gives
-${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq{2\over n}|\langle{\mathbf{Y}}-{\mathbf{X}}B,{\mathbf{X}}(\widehat{B}-B)\rangle|+\lambda(r-\widehat{r})$
-$\leq 2\|\widehat{B}-B\|_{*}{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}+\lambda(r-\widehat{r})$
-$\leq 2\sqrt{r+\widehat{r}}~{}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+\lambda(r-\widehat{r})$ by $\text{by $\mathcal{E}_{\lambda}^{\prime}$}.$
+[…]
+[…]
+[…]
 Using the properties on $\mathcal{E}_{X}$, we find
-$\|\widehat{B}-B\|_{F}^{2}$ $\leq{2\over c}2\sqrt{\widehat{r}+r}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+{2\over c}\lambda(r-\widehat{r})$
-$\leq{8\lambda\over c^{2}C}(\widehat{r}+r)+{1\over 2}\|\widehat{B}-B\|_{F}^{2}+{2\over c}\lambda(r-\widehat{r})$ by $2xy\leq x^{2}/2+2y^{2}$
-$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ for $\text{for $C=4/c$}.$
+[…]
+[…]
+$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ $\text{for $C=4/c$}.$
 In the same way, on the event $\mathcal{E}_{X}$,
-${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq\sqrt{{8\over cC}\lambda(r+\widehat{r})}~{}{1\over\sqrt{n}}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}+\lambda(r-\widehat{r})$
-$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ for $\text{for $C=4/c$}.$
+[…]
+$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ $\text{for $C=4/c$}.$
 This implies the rates for both $\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ and $\|\widehat{B}-B\|_{F}^{2}$.
@@
 Proof.
-By (B.31), it suffices to bound from above
+By (A.31), it suffices to bound from above
 $\textrm{I}=\left\|(\widehat{\Sigma}-\Sigma)B\right\|_{\rm op},\qquad\textrm{II}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}\right\|_{\rm op}=\left\|{1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-{M}D_{\pi}\right\|_{\rm op}.$
@@
 ∎
-Appendix C Technical lemmas
+A.7 Proof of the equivalence between (1.1) and (4.1)
+Proof.
+Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
+$B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}$ $=\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}$
+so that
+$\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})$
+$=\operatorname*{arg\,min}_{\ell\in[L]}~2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
+Write the singular value decomposition of $\Sigma_{w}^{-1/2}{M}$ as $U\Lambda V^{\top}$ with $U\in\mathbb{R}^{p\times r}$ containing the first $r$ singular vectors and $r={\rm rank}(\Sigma_{w}^{-1/2}{M})$.
+It follows that the above equals to
+$\operatorname*{arg\,min}_{\ell\in[L]}~2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}$
+$=\operatorname*{arg\,min}_{\ell\in[L]}~2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}$
+which, in view of (1.1) and (3.2), completes the proof. ∎
+Appendix B Technical lemmas
 The following lemma provides a few useful facts on quantities related with $B$ in (1.8). Write
-$\Omega:=D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}.$ (C.1)
+$\Omega:=D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}.$ (B.1)
 Lemma 14.
 With $B$ defined in (1.8), we have
-${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1},$ (C.2)
-$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (C.3)
-$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1}.$ (C.4)
+${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~\Omega^{-1},$ (B.2)
+$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (B.3)
+$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~\Omega^{-1}.$ (B.4)
 Consequently, for any $k,\ell\in[L]$, we have
-$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (C.5)
-$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (C.6)
-$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (C.7)
+$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (B.5)
+$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (B.6)
+$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (B.7)
 Proof.
@@
 $=\mathbf{I}_{L}-\left(\mathbf{I}_{L}+{M}^{\top}\Sigma_{w}^{-1}{M}D_{\pi}\right)^{-1}$
-$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (C.8)
+$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (B.8)
 $={M}^{\top}\Sigma_{w}^{-1}{M}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$
-proving (C.2). By using the above identity (C.8), we have
+proving (B.2). By using the above identity (B.8), we have
 $D_{\pi}-D_{\pi}{M}^{\top}B=\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1},$
-proving (C.3). Since Lemma 1 and (C.3) imply
+proving (B.3). Since Lemma 1 and (B.3) imply
 $B=B^{*}\Omega^{-1},$
-(C.4) follows from
+(B.4) follows from
 $B^{\top}\Sigma_{w}B=\Omega^{-1}B^{*\top}\Sigma_{w}B^{*}\Omega^{-1}=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}\Omega^{-1}.$
-To prove (C.5) – (C.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
-$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (C.9)
-yielding the second bound in (C.5). The other bound in (C.5) follows by observing that
+To prove (B.5) – (B.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
+$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (B.9)
+yielding the second bound in (B.5). The other bound in (B.5) follows by observing that
 […]
-For (C.6), on the one hand, similar arguments yield
+For (B.6), on the one hand, similar arguments yield
 $B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}{M}D_{\pi}{M}^{\top}\Sigma^{-1}\mu_{\ell}\leq{\pi_{\ell}^{2}\over\pi_{\min}}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}.$
-On the other hand, using (C.2) and (C.3) proves
+On the other hand, using (B.2) and (B.3) proves
 $B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq{\bm{e}}_{\ell}^{\top}\Omega^{-1}({M}^{\top}\Sigma_{w}^{-1}{M})^{2}\Omega^{-1}{\bm{e}}_{\ell}\leq 1.$
@@
 ∎
-C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
-Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (B.1).
+B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
+Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (A.1).
 Lemma 18.
@@
 $\quad+\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$
-$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
+$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{L}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{L}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
 Proof.
@@
 $\left|\sum_{k=1}^{L}(\widehat{\pi}_{k}-\pi_{k})v_{1}^{\top}\mu_{k}\mu_{k}^{\top}v_{2}\right|$ $\lesssim\sqrt{t\sum_{k=1}^{L}\pi_{k}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over n}+{t\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}\over n}$
-$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (C.11)
+$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (B.11)
 with probability at least $1-2e^{-t}$. Furthermore, on the event $\mathcal{E}_{\pi}$, 1 and $L\log(n)\leq n$ imply
@@
 and after we invoke Lemma 15 twice with $v=v_{1},u={M}^{\top}v_{2}$ and $v=v_{2},u={M}^{\top}v_{1}$, respectively, we get that
-$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (C.12)
-$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (C.13)
+$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{L}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (B.12)
+$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{L}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (B.13)
 hold with probability at least $1-4e^{-t}$. Collecting all three bounds completes the proof. ∎
@@
 For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$.
-Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (C.5), (C.6) and (C.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,
+Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (B.5), (B.6) and (B.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,
 $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ $\lesssim B_{k}^{\top}\Sigma_{w}B_{k}\sqrt{\log(n)\over n}+\|{M}^{\top}B_{k}\|_{\infty}\sqrt{B_{k}^{\top}{M}{M}^{\top}B_{k}}\sqrt{\log(n)\over nL}$
@@
 uniformly over $\mathcal{S}^{p}$, completing the proof. ∎
-Appendix D Auxiliary lemmas
+B.2 Auxiliary lemmas
 The following lemmas are proved in Bing and Wegkamp (2023).